### v3

# Weyl symmetry without the traceless condition

Abstract
We show that the requirement that the trace of the stress-energy tensor of matter must vanish if invariance under Weyl transformations is a symmetry of a given gravitational theory is not universal. This requirement holds whenever the masses of timelike fields are constant parameters that are not transformed by conformal transformations, or when the energy density of perfect fluids transforms as $\rho\rightarrow\Omega^{-3}\rho$. In contrast, if the masses of timelike fields are point-dependent quantities transforming under conformal transformations as $m\rightarrow\Omega^{-1}m$, and the energy density of perfect fluids transforms as $\rho\rightarrow\Omega^{-4}\rho$, the Weyl symmetry does not require the vanishing of the trace of the matter SET. In consequence, any matter fields, regardless of whether the trace of their stress-energy tensor vanishes or not, can be coupled to gravity. The phenomenological and physical consequences of the novel result are drawn.
I Introduction
$g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu},\;\Phi_{i}\rightarrow\hat{\Phi}_{i}=\Omega^{w_{i}}\Phi_{i},$ (1)
where $\Omega=\Omega(x)$ is the conformal factor ($\Omega(x)>0$), $\Phi_{i}$ are other gravitational and matter fields present, and $w_{i}$ are the conformal weights of these fields. Here we call this “Dicke’s principle.”
The high-energy physics (HEP) argument that, due to the Planck mass $M_{\text{pl}}$, a dimensionful parameter, gravity cannot be conformally invariant, seems to oppose Dicke’s principle dicke-1962 . In general, the presence of any mass term is believed to break the conformal invariance deser-1970 ; callan_prd_1970 . However, strictly speaking, this is true only if we consider constant masses that do not transform under (1): $m\rightarrow m$. As stated in bekenstein_1980 , we must differentiate between LSTs:
$g_{\mu\nu}\rightarrow\Omega^{2}g_{\mu\nu},\;\Phi_{i}\rightarrow\Omega^{w_{i}}\Phi_{i},\;m\rightarrow m,$ (2)
and Weyl transformations (WTs) (same as transformations of units dicke-1962 ):
$g_{\mu\nu}\rightarrow\Omega^{2}g_{\mu\nu},\;\Phi_{i}\rightarrow\Omega^{w_{i}}\Phi_{i},\;m\rightarrow\Omega^{-1}m,$ (3)
where, in addition to CTs (1), the following transformation property of the masses is assumed:
$m\rightarrow\hat{m}=\Omega^{-1}m.$ (4)
That is, the mass parameters are fields with conformal weight $w_{m}=-1$ hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 .
According to bekenstein_1980 , scale transformations (2) are not units’ transformations; they are active enlargements of a system whose usefulness depends on the absence of a length scale. Meanwhile, WTs are units’ transformations (3) whose physical and geometrical meanings follow from the existence of some scale of length.
In the present document, as is customary when working within the framework of scalar-tensor theories (STTs) dicke-1962 ; bekenstein_1980 ; faraoni_rev ; fujii_book ; faraoni_book ; faraoni_prd_2007 ; quiros_ijmpd_2019 ; hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 , we consider WTs (3), instead of LSTs (2). In consequence, it is assumed that the masses are point-dependent quantities that, under (1), transform like (4). Whether the mass parameter transforms or not under the Weyl transformations plays a central role in the demonstration of conformal form-invariance of the action of fundamental matter fields quiros_prd_2025 :
$S_{m}=\int d^{4}x\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu}),$ (5)
where $L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$ is the Lagrangian of the fundamental matter fields $\chi_{i}$, with i=1,2,3,,Ni=1,2,3,...,N, and $N$ – the number of matter fields that are minimally coupled to the metric $g_{\mu\nu}$. Here, the stress-energy tensor (SET) of the matter fields is defined as usual;
$T^{(m)}_{\mu\nu}=-\frac{2}{\sqrt{-g}}\frac{\delta{\cal L}_{m}}{\delta g^{\mu\nu}},$ (6)
where ${\cal L}_{m}=\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$ is the Lagrangian density of matter.
For fermions and gauge bosons, it has been shown in quiros_prd_2025 that if their masses are point-dependent quantities transforming as (4), the matter action (5) is form-invariant under WTs
$g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu},\;\chi_{i}\rightarrow\hat{\chi}_{i}=\Omega^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow\hat{m}=\Omega^{-1}m,$ (7)
where $w_{\chi_{i}}$ is the conformal weight of the matter field $\chi_{i}$. But if the masses are constant parameters, then (5) is not form-invariant under (2). Take, for illustration, the Lagrangian density of the Dirac fermion: [^1] Here we use the following signature of the metric: (,+,+,+)(-,+,+,+). We follow the notation of quiros_prd_2025 (see also the Appendix A).
[…]
A large body of applications of gravitational theories is associated with background matter in the form of a perfect fluid, rather than with fundamental fields. [^3] For a perfect fluid with energy density $\rho$ and barotropic pressure $p$, the SET is given by $T^{(m)}_{\mu\nu}=(\rho+p)u_{\mu}u_{\nu}+pg_{\mu\nu}$, where $u^{\mu}=dx^{\mu}/d\tau$ is the four-velocity of the fluid ($d\tau=ids$ $\Rightarrow d\tau^{2}=-ds^{2}$). Therefore, the trace of the SET of the perfect fluid reads: $T^{(m)}=-\rho+3p$. It vanishes only for radiation with the equation of state (EOS) $p=\rho/3$. Therefore, it is of interest to look for the transformation properties of the action of perfect fluids under Weyl transformations (3). In quiros_prd_2025 , it has also been shown that, in addition to the action of fundamental matter fields, that of perfect fluids is form-invariant under WTs as well. Actually, the action of a perfect fluid with energy density $\rho$ and barotropic pressure $p$ can be written in either form: $S_{\text{fluid}}=-\int d^{4}x\sqrt{-g}\,\rho$ or $S_{\text{fluid}}=\int d^{4}x\sqrt{-g}\,p$. Unless the perfect fluid couples explicitly to the curvature (which is not the case in this paper), both Lagrangian densities are equivalent faraoni_2009 . Here, for definiteness, we choose the former action for the perfect fluid. According to dimensional arguments, if we assume point-dependent masses that transform as $m\rightarrow\Omega^{-1}m$ under CTs (1), the energy density of the fluid transforms as $\rho\rightarrow\Omega^{-4}\rho$. In fact, the energy density has units: $M]/[L]^{3}$, where $M]\rightarrow\Omega^{-1}[M]$ and $L]\rightarrow\Omega[L]$ are the transformation laws of the units of mass and length under CTs, respectively. Hence, $M]/[L]^{3}\rightarrow\Omega^{-4}[M]/[L]^{3}$. The barotropic pressure of the fluid $p$ transforms in the same way: $p\rightarrow\Omega^{-4}p$. Since under (1), $\sqrt{-g}\rightarrow\sqrt{-\hat{g}}=\Omega^{4}\sqrt{-g}$ and $\rho\rightarrow\hat{\rho}=\Omega^{-4}\rho$, then $\int d^{4}x\sqrt{-g}\,\rho=\int d^{4}x\sqrt{-\hat{g}}\,\hat{\rho}$. That is, the action of the perfect fluid is form-invariant under WTs. Note, however, that if we assume that the mass unit is not transformed, then, under CTs $\rho\rightarrow\hat{\rho}=\Omega^{-3}\rho$, which means that the action of the perfect fluid is not form-invariant: $\int d^{4}x\sqrt{-g}\,\rho=\int d^{4}x\sqrt{-\hat{g}}\,\Omega\hat{\rho}$.
The requirement of vanishing of the matter SET trace, in connection with LSS in curved spacetime, was derived in deser-1970 , in the case of a scalar field non-minimally coupled to gravity. [^4] The vanishing of the matter SET trace as a requirement for conformal symmetry to hold ccj_ann_phys_1970 was primarily investigated within the HEP framework, where gravity is ignored, so the background spacetime is flat space with the Minkowski metric. Note, in this regard, that, within the HEP context, conformal symmetry refers to the 15-parameter conformal group that consists, in addition to the Poincaré group, of the 1-parameter group of dilatations and the 4-parameter special conformal group kastrup_1966 ; jackiw_1972 . These are spacetime transformations. In contrast, the conformal transformations that we consider in this paper are transformations of fields, including the spacetime metric and any matter fields that could be coupled to gravity, that do not affect either the spacetime points or the coordinates that label these points. That is, transformations (1) are not spacetime diffeomorphisms. For that reason, in this document, to avoid confusion between conformal transformations (1) and the HEP 15-parameter group of transformations, we will refer to the former mostly as Weyl transformations. In consequence, we say, for example, “Weyl symmetry” instead of “conformal symmetry.” The gravitational action of the theory investigated in this bibliographic reference, called here “conformal general relativity” (CGR), reads:
$S_{\text{grav}}=\frac{1}{2}\int d^{4}x\sqrt{-g}\left[\frac{1}{6}\,\phi^{2}R+(\partial\phi)^{2}\right],$ (9)
where $R$ is the curvature scalar, $\phi$ is the non-minimally coupled scalar field with mass dimensions, and we use the following shorthand notation: $(\partial\phi)^{2}\equiv g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi$. In deser-1970 , it is demonstrated that the form-invariance of (9) under LSTs (2), is possible thanks to the vanishing trace of the matter SET: $T^{(m)}\equiv g^{\mu\nu}T^{(m)}_{\mu\nu}=0$.
In this document, we shall demonstrate that the traceless condition $T^{(m)}=0$ is not a universal requirement for Weyl symmetry to materialize. The result of reference deser-1970 , that only matter fields with traceless SET can be coupled to gravity in CGR theory (9), is right only if we assume that the masses of timelike fields are constant parameters that are not transformed by CTs (1), so the energy density of a perfect fluid transforms as: $\rho\rightarrow\hat{\rho}=\Omega^{-3}\rho$; that is, if we consider local scale transformations (2) instead of Weyl transformations (3). Under this assumption, the Lagrangian density of the matter fields breaks LSS. In contrast, if we assume that the masses of timelike fields are point-dependent quantities that, under the CTs (1), transform as $m\rightarrow\hat{m}=\Omega^{-1}m$, and consequently the energy density of the perfect fluid transforms as $\rho\rightarrow\hat{\rho}=\Omega^{-4}\rho$, the Lagrangian density of the fundamental matter fields, as well as the Lagrangian density of perfect fluids, are form-invariant under the WTs (3):
$\int d^{4}x\sqrt{-g}\,L_{m}(\chi_{i},g_{\mu\nu})=\int d^{4}x\sqrt{-\hat{g}}\,\Omega^{-4}\,L_{m}(\Omega^{-w_{\chi_{i}}}\hat{\chi}_{i},\Omega^{-2}\hat{g}_{\mu\nu})=\int d^{4}x\sqrt{-\hat{g}}\,L_{m}(\hat{\chi}_{i},\hat{g}_{\mu\nu}),$
that is,
$\int d^{4}x\sqrt{-g}\,L_{m}(\chi_{i},g_{\mu\nu})=\int d^{4}x\sqrt{-\hat{g}}\,L_{m}(\hat{\chi}_{i},\hat{g}_{\mu\nu}).$ (10)
Similarly, the action of a perfect fluid is form-invariant under (7) as well: $\int d^{4}x\sqrt{-g}\,\rho=\int d^{4}x\sqrt{-\hat{g}}\,\hat{\rho}$. In these cases, since the Lagrangian density of matter is form-invariant under WTs, the Ward identity associated with Weyl invariance leads to the following nonvanishing variational derivative: $\delta{\cal L}_{m}/\delta\phi=\sqrt{-g}\,T^{(m)}/\phi$, which provides the term with the trace of the stress-energy tensor of matter in the EOM of the scalar field $\phi$. In contrast, if the Lagrangian density of the matter fields is not form-invariant under CTs (1), the Ward identity does not hold and $\delta{\cal L}_{m}/\delta\phi=0$.
Our main hypothesis in this work is that Dicke’s principle holds in nature. That is, the laws of gravity are form-invariant under Weyl transformations (3). In consequence, we consider that the masses of fundamental timelike matter fields $m_{i}$ are point-dependent quantities bekenstein_1980 ; hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 : $m_{i}(x)=\kappa_{i}\,\phi(x)$, where $\kappa_{i}$ are dimensionless constants. [^5] This choice is dictated by the fact that, since the scalar field $\phi$ has mass dimensions, under WTs the mass parameter transforms as $\phi$ does; that is, $m\propto\phi$. These transform as (4) under Weyl transformations, while the energy density of a perfect fluid transforms as $\rho\rightarrow\Omega^{-4}\rho$.
We have organized the present document as follows. In Section II, we set up the notation and briefly review the most important aspects of the work of reference deser-1970 , where the invariance of CGR under LSTs (2) is considered. This is required for further comparison with the results of Section III, where the equations of motion (EOM) are derived from the overall CGR action under the assumptions of the present work, above all form-invariance of the matter action under WTs (3). In Section IV, we derive some physical consequences of the CGR theory, in particular, the fifth force required by form-invariance under WTs, which does not interact with radiation, thus acting as a ”dark force.” In this section, we also discuss the two possible approaches to CTs in the configuration space: i) passive and ii) active approaches. We shall show that only the active approach is suitable for the investigation of the physical and phenomenological consequences of the Weyl symmetry. In Section V, we discuss gauge freedom within CGR, in connection with Weyl symmetry, under the active approach to WTs. We draw the similitude of the picture arising from gauge freedom within the CGR theory and the “many-worlds” interpretation of quantum mechanics everett_rmp_1957 ; wheeler ; dewitt_phys_rev_1967 ; dewitt ; barvinsky_cqg_1990 ; omnes_rmp_1992 ; tegmark_1998 ; garriga_prd_2001 ; zurek_rmp_2003 ; tegmark_nature_2007 ; quiros_prd_2023 . In Sections VI and VII, we expose some important cosmological consequences of the present setup, including the possible explanation of the galactic rotation curves in some gauge of the CGR theory, as well as some encouraging aspects of Weyl invariant cosmology. Possible observational consequences of CGR theory, mainly related to measurements of the redshift of frequencies, are analyzed in Section VIII. The novel results of the present work are discussed in Section IX, accompanied by brief concluding remarks.
II Local scale symmetry and the trace of the stress-energy tensor of matter
The gravitational action of CGR theory (9) is not only invariant but also form-invariant under the following conformal transformations of the gravitational fields:
$g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu},\;\phi\rightarrow\hat{\phi}=\Omega^{-1}\phi.$ (11)
Independent variations of the CGR gravitational Lagrangian density
${\cal L}_{\text{grav}}=\frac{1}{2}\sqrt{-g}\left[\frac{1}{6}\phi^{2}R+(\partial\phi)^{2}\right],$ (12)
with respect to the metric and to the scalar field, neglecting divergence terms, yield deser-1970 :
$\frac{\delta{\cal L}_{\text{grav}}}{\delta g^{\mu\nu}}=\sqrt{-g}\,\frac{\phi^{2}}{12}\,{\cal E}_{\mu\nu},\;\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}=\sqrt{-g}\left(\frac{1}{6}\,\phi R-\nabla^{2}\phi\right),$ (13)
where $G_{\mu\nu}\equiv R_{\mu\nu}-g_{\mu\nu}R/2$ is the Einstein tensor and
${\cal E}_{\mu\nu}\equiv G_{\mu\nu}-\frac{6}{\phi^{2}}\,\Theta^{(\phi)}_{\mu\nu},\;\Theta^{(\phi)}_{\mu\nu}\equiv\frac{1}{6}\left(\nabla_{\mu}\nabla_{\nu}-g_{\mu\nu}\nabla^{2}\right)\phi^{2}-T^{(\phi)}_{\mu\nu}$. (14)
The modified SET of the scalar field $\Theta^{(\phi)}_{\mu\nu}$ ccj_ann_phys_1970 , has finite elements to all orders in the perturbation parameter and defines the same four-momentum and Lorentz generators as the conventional tensor: [^6] Note that in the expression for $\Theta^{(\phi)}_{\mu\nu}$ the standard stress-energy tensor of the scalar field $T^{(\phi)}_{\mu\nu}$, has the wrong sign. This is because, in the gravitational Lagrangian density (12), the kinetic energy density of the scalar field appears with the wrong sign, which is a requirement for the CGR gravitational Lagrangian density (12) to be invariant under WTs (11). This has no physical consequences, since $\phi$ is a gauge field that does not obey a specific EOM.
$T^{(\phi)}_{\mu\nu}\equiv\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}g_{\mu\nu}(\partial\phi)^{2}.$ (15)
Here we use the shorthand notation $\nabla^{2}\equiv g^{\mu\nu}\nabla_{\mu}\nabla_{\nu}$. If we take into account the trace of the tensor ${\cal E}_{\mu\nu}$:
${\cal E}\equiv g^{\mu\nu}{\cal E}_{\mu\nu}=-R+\frac{6}{\phi}\,\nabla^{2}\phi,$ (16)
the equations in (13) can be written in the following form:
$g^{\mu\nu}\frac{\delta{\cal L}_{\text{grav}}}{\delta g^{\mu\nu}}=\sqrt{-g}\frac{\phi^{2}}{12}\,{\cal E}=-\sqrt{-g}\frac{\phi}{2}\left(\frac{1}{6}\,\phi R-\nabla^{2}\phi\right),\;\frac{\phi}{2}\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}=\sqrt{-g}\frac{\phi}{2}\left(\frac{1}{6}\,\phi R-\nabla^{2}\phi\right),$
respectively. If we compare these equations, we obtain the following Ward identity:
$g^{\mu\nu}\frac{\delta{\cal L}_{\text{grav}}}{\delta g^{\mu\nu}}=-\frac{\phi}{2}\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}.$ (17)
This identity is a consequence of the invariance of the CGR gravitational action (9) under CT (11). Actually, if this action is form-invariant under infinitesimal CTs:
$\delta g_{\mu\nu}=2\theta\,g_{\mu\nu}\,\left(\delta g^{\mu\nu}=-2\theta\,g^{\mu\nu},\;\delta\sqrt{-g}=4\theta\sqrt{-g}\right),\;\delta\phi=-\theta\,\phi,$ (18)
where for convenience (and only temporarily) we rescaled the conformal factor: $\Omega=e^{\theta}$, then $S_{\text{grav}}\rightarrow S_{\text{grav}}+\delta S_{\text{grav}}$, where $\delta S_{\text{grav}}=0$; so that, under (18), $S_{\text{grav}}\rightarrow S_{\text{grav}}$. Hence;
$\delta S_{\text{grav}}=\frac{1}{2}\int d^{4}x\left(\frac{\delta{\cal L}_{\text{grav}}}{\delta g^{\mu\nu}}\delta g^{\mu\nu}+\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}\delta\phi\right)=-\theta\int d^{4}x\left(g^{\mu\nu}\frac{\delta{\cal L}_{\text{grav}}}{\delta g^{\mu\nu}}+\frac{\phi}{2}\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}\right)=0,$
where in the last equality, we substituted $\delta g^{\mu\nu}$ and $\delta\phi$ from infinitesimal CTs (18). Specifically, from the last equation, the Ward identity (17) follows.
After applying the variational principle of least action, from (13) we obtain the vacuum Einstein EOM and the Klein-Gordon-type EOM:
$\frac{\delta{\cal L}_{\text{grav}}}{\delta g^{\mu\nu}}$ $=0\;\Rightarrow\;{\cal E}_{\mu\nu}=0\;\Rightarrow\;G_{\mu\nu}=\frac{6}{\phi^{2}}\,\Theta^{(\phi)}_{\mu\nu}$, (19)
$\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}$ $=0\;\Rightarrow\;\left(\nabla^{2}-\frac{1}{6}\,R\right)\phi=0,$ (20)
respectively. The trace of the Einstein EOM (19) reads $R+6\nabla^{2}\phi/\phi=0$, which, after some simple algebra, can be written as the KG-EOM (20). Hence, the latter is not an independent equation. Consequently, the scalar field $\phi$ does not obey an independent EOM. This is consistent with the new degree of freedom (DOF) associated with the Weyl symmetry of (9) since, in addition to the four DOFs for making spacetime diffeomorphisms, there is a DOF for making CTs (11), which is exhausted by the gauge field $\phi$.
Let us next consider matter fields coupled to CGR;
$S_{\text{tot}}=\frac{1}{2}\int d^{4}x\sqrt{-g}\left[\frac{1}{6}\,\phi^{2}R+(\partial\phi)^{2}\right]+\int d^{4}x\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu}),$ (21)
where, since it is assumed in deser-1970 that the CTs do not transform the masses, the matter Lagrangian of time-like matter fields breaks LSS. In this case, the Ward identity for matter fields is not satisfied, so $\delta{\cal L}_{m}/\delta\phi=0$, because the matter Lagrangian density does not depend explicitly on $\phi$.
Variations of the total CGR action (21) with respect to the metric $g_{\mu\nu}$ and with respect to $\phi$, yield
$\frac{\delta{\cal L}_{\text{tot}}}{\delta g^{\mu\nu}}=\sqrt{-g}\,\frac{\phi^{2}}{12}\,\left[{\cal E}_{\mu\nu}-\frac{6}{\phi^{2}}\,T^{(m)}_{\mu\nu}\right],\;\frac{\delta{\cal L}_{\text{tot}}}{\delta\phi}=\sqrt{-g}\left(\frac{1}{6}\,\phi R-\nabla^{2}\phi\right),$ (22)
respectively, where we have omitted divergence terms and assumed that,
$\frac{\delta{\cal L}_{m}}{\delta\phi}=0\;\Rightarrow\;\frac{\delta{\cal L}_{\text{tot}}}{\delta\phi}=\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}.$ (23)
If we apply the variational principle, from (22) we get the following EOMs:
$\frac{\delta{\cal L}_{\text{tot}}}{\delta g^{\mu\nu}}$ $=0\;\Rightarrow\;{\cal E}_{\mu\nu}=\frac{6}{\phi^{2}}\,T^{(m)}_{\mu\nu}\;\Rightarrow\;G_{\mu\nu}=\frac{6}{\phi^{2}}\left[\Theta^{(\phi)}_{\mu\nu}+T^{(m)}_{\mu\nu}\right$], (24)
$\frac{\delta{\cal L}_{\text{tot}}}{\delta\phi}$ $=0\;\Rightarrow\;\left(\nabla^{2}-\frac{1}{6}\,R\right)\phi=0.$ (25)
Note that the Einstein EOM (19) has been modified by adding the SET of matter, whereas the KG-type EOM (20) is not modified since the Lagrangian density of the matter fields ${\cal L}_{m}=\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$ does not depend explicitly on $\phi$, that is, equation (23) takes place. This results in the following requirement. If we compare the trace of (24):
$-R+\frac{6}{\phi}\nabla^{2}\phi=\frac{6}{\phi^{2}}T^{(m)}\;\Rightarrow\;\left(\nabla^{2}-\frac{1}{6}\,R\right)\phi=\frac{1}{\phi}\,T^{(m)},$ (26)
with the EOM (25), we obtain the “traceless condition:” $T^{(m)}=0$. Therefore, only matter fields or fluids with vanishing SET trace: photons and radiation, in general, can be coupled to gravity in this local scale-invariant theory. [^7] Non-relativistic matter fields that behave like radiation can also be coupled to CGR faria-a .
This result has been the focus of studies on Weyl symmetry for decades, but not on the Weyl symmetry we are investigating in the present document. Recall that here, following bekenstein_1980 , we make a distinction between local scale transformations (2) and Weyl transformations (3), called by Dicke as transformations of units. Historically, LSTs have been the most studied Weyl-type transformations. In consequence, Weyl symmetry is usually identified with invariance under LSTs (2). Here, we associate Weyl symmetry with units’ transformations (3), which are popular in the study of scalar-tensor gravitational theories fujii_book ; faraoni_book ; flanagan_cqg_2004 ; faraoni_prd_2007 ; catena_prd_2007 ; chiba_2013 ; capozziello_prd_2013 ; jarv_2015 ; quiros_ijmpd_2019 .
III Weyl symmetry does not require vanishing trace of the stress-energy tensor of matter
In this section, we will demonstrate that, if we assume that the masses of timelike fields are themselves point-dependent fields that transform as $m\rightarrow\hat{m}=\Omega^{-1}m$, while the energy density of a perfect fluid transforms as $\rho\rightarrow\hat{\rho}=\Omega^{-4}\rho$, it is not necessary to make any changes in (21) to get a consistent coupling of any matter fields. That is, under these assumptions, the coupling of the matter fields in (21) preserves the Weyl symmetry regardless of whether $T^{(m)}$ vanishes or not. Our demonstration is based on the Weyl invariance of matter action in this case, as demonstrated in quiros_prd_2025 for fundamental fields and perfect fluids.
Under the above assumptions, for fermions and gauge bosons as well as for perfect fluids, the matter action (5) is not only invariant but also form-invariant under WTs. That is, under (7), the following equality takes place:
$\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})=\sqrt{-\hat{g}}\,L_{m}(\hat{\chi}_{i},\hat{\partial}\hat{\chi}_{i},\hat{g}_{\mu\nu}).$ (27)
From this equation it follows that if $g_{\mu\nu}$ is the physical metric with respect to the matter fields $\chi_{i}$; that is, the matter fields $\chi_{i}$ are minimally coupled to $g_{\mu\nu}$, the conformal metric $\hat{g}_{\mu\nu}$ is the physical metric with respect to the conformal matter fields $\hat{\chi}_{i}$ which are minimally coupled to $\hat{g}_{\mu\nu}$. In other words, if the metric $g_{\mu\nu}$ is the physical metric on some gauge, its conformal metric $\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu}$ is the physical metric on the conformal gauge. Equation (27) plays an important role because it shows that the total action (21) is Weyl invariant regardless of the type of matter coupled to gravity: fermions, gauge bosons, or perfect fluids.
The transformation of masses $m\rightarrow\Omega^{-1}m$ and the consequent transformation of the energy density (and pressure) of a perfect fluid: $\rho\rightarrow\Omega^{-4}\rho$, is fundamental for the demonstration of form-invariance of the Lagrangian density of timelike fundamental matter fields quiros_prd_2025 . In this case, the Lagrangian density of matter fields is form-invariant under WTs, so the variational derivative: $\delta{\cal L}_{m}/\delta\phi$ obeys the Ward identity; that is, it is nonvanishing. This is the primary assumption underlying our approach to the Weyl symmetry of the CGR theory, which distinguishes our result from the well-known result that the matter SET trace must vanish if local scale invariance is a symmetry of the CGR theory.
Let us apply the variational procedure under the assumption that the action of the matter degrees of freedom coupled to gravity: fermions, gauge bosons, and perfect fluids, is form-invariant under the following WTs:
$g_{\mu\nu}\rightarrow\Omega^{2}g_{\mu\nu},\;\phi\rightarrow\Omega^{-1}\phi,\;\chi_{i}\rightarrow\Omega^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow\Omega^{-1}m.$ (28)
Let us consider the Lagrangian density of matter fields: ${\cal L}_{m}={\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$. Since it is invariant under WTs (28), then, under infinitesimal Weyl transformations (18) plus $\delta\chi_{i}=w_{\chi_{i}}\theta\chi_{i}$ and $\delta m=-\theta m$, we have $\delta{\cal L}_{m}=\nabla_{\mu}V^{\mu}$, where $V^{\mu}$ is some vector or $\delta{\cal L}_{m}=0$. In any case, $\delta S_{m}=\int d^{4}x\,\delta{\cal L}_{m}=0$, so that under (18), $S_{m}\rightarrow S_{m}+\delta S_{m}=S_{m}$; that is, the matter action is invariant under infinitesimal Weyl transformations
$\delta g_{\mu\nu}=2\theta\,g_{\mu\nu}\,\left(\delta g^{\mu\nu}=-2\theta\,g^{\mu\nu},\;\delta\sqrt{-g}=4\theta\sqrt{-g}\right),\;\delta\phi=-\theta\,\phi,\;\delta\chi_{i}=w_{\chi_{i}}\theta\chi_{i},\;\delta m=-\theta m.$ (29)
For the variation of the Lagrangian density of matter, we have
$\delta{\cal L}_{m}=\frac{\delta{\cal L}_{m}}{\delta g^{\mu\nu}}\delta g^{\mu\nu}+\frac{\delta{\cal L}_{m}}{\delta\phi}\delta\phi=0\;\Rightarrow\;\frac{\delta{\cal L}_{m}}{\delta g^{\mu\nu}}\delta g^{\mu\nu}+\frac{\delta{\cal L}_{m}}{\delta\phi}\delta\phi=-2\theta\left(g^{\mu\nu}\frac{\delta{\cal L}_{m}}{\delta g^{\mu\nu}}+\frac{\phi}{2}\frac{\delta{\cal L}_{m}}{\delta\phi}\right)=0,$ (30)
where, in the last step, we took into account infinitesimal WTs (29) and, for simplicity, we have assumed that the EOMs of the matter fields
$\frac{\delta{\cal L}_{m}}{\delta\chi_{i}}=0\;\Rightarrow\;\frac{\partial{\cal L}_{m}}{\partial\chi_{i}}-\nabla_{\mu}\left[\frac{\partial{\cal L}_{m}}{\partial(\partial_{\mu}\chi_{i})}\right]=0,$ (31)
are satisfied. [^8] For classical gravitational problems, it is assumed that the matter fields are a priori given. From the last equality in (30), it follows the Ward identity:
$g^{\mu\nu}\frac{\delta{\cal L}_{m}}{\delta g^{\mu\nu}}=-\frac{\phi}{2}\frac{\delta{\cal L}_{m}}{\delta\phi}.$ (32)
If we substitute in this identity the definition of the matter SET (6), the Ward identity can be written in the following outstanding form:
$\frac{\delta{\cal L}_{m}}{\delta\phi}=\frac{\sqrt{-g}}{\phi}\,T^{(m)}.$ (33)
We have
$\delta S_{\text{tot}}=\delta S_{\text{grav}}+\delta S_{m}=\int d^{4}x$ $\left[\frac{\delta{\cal L}_{\text{grav}}}{\delta g^{\mu\nu}}\delta g^{\mu\nu}+\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}\delta\phi+\frac{\delta{\cal L}_{m}}{\delta g^{\mu\nu}}\delta g^{\mu\nu}+\frac{\delta{\cal L}_{m}}{\delta\phi}\delta\phi\right]=0,$
that is,
$\delta S_{\text{tot}}$ $=\int d^{4}x\left[\left(\frac{\delta{\cal L}_{\text{grav}}}{\delta g^{\mu\nu}}+\frac{\delta{\cal L}_{m}}{\delta g^{\mu\nu}}\right)\delta g^{\mu\nu}+\left(\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}+\frac{\delta{\cal L}_{m}}{\delta\phi}\right)\delta\phi\right]$
$=\int d^{4}x\sqrt{-g}\left\{\frac{\phi^{2}}{12}\left[{\cal E}_{\mu\nu}-\frac{6}{\phi^{2}}T^{(m)}_{\mu\nu}\right]\delta g^{\mu\nu}+\left[-\nabla^{2}\phi+\frac{\phi}{6}R+\frac{1}{\phi}T^{(m)}\right]\delta\phi\right\}=0,$ (34)
where in the last line of (34), in the factor multiplying $\delta\phi$ (expression within square brackets), we have taken into account the modified Ward identity (33). The variational principle leads to the following EOM:
${\cal E}_{\mu\nu}=\frac{6}{\phi^{2}}T^{(m)}_{\mu\nu}$, (35)
$\nabla^{2}\phi-\frac{\phi}{6}R=\frac{1}{\phi}T^{(m)}.$ (36)
This is an outstanding result, since the trace of the matter SET is not required to vanish. Note that the KG-EOM (36) is not an independent equation, as it coincides with the trace of Einstein-type EOM (35). That is, the scalar field $\phi$ is not a dynamical field since it does not obey an independent EOM.
IV CGR: Physics beyond GR
The CGR theory deser-1970 whose Lagrangian density is given by the following expression (see action (21)):
${\cal L}_{\text{tot}}=\frac{\sqrt{-g}}{2}\left[\frac{1}{6}\,\phi^{2}R+(\partial\phi)^{2}\right]+{\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu}),$ (37)
is form-invariant under the WTs (28). The independent EOM derived from (37), is the following Einstein-type equation
${\cal E}_{\mu\nu}=\frac{6}{\phi^{2}}\,T^{(m)}_{\mu\nu}\;\Rightarrow\;G_{\mu\nu}$ $=\frac{6}{\phi^{2}}\left[\Theta^{(\phi)}_{\mu\nu}+T^{(m)}_{\mu\nu}\right$], (38)
where the tensor $\Theta^{(\phi)}_{\mu\nu}$ is defined in (14). There is no independent EOM that the gauge field $\phi$ must obey, since the KG-type equation (36), which is derived from (37) by varying with respect to $\phi$, coincides with the trace of (38).
Another independent equation is provided by the continuity equation. If we take the divergence of (38), and use the second Bianchi identity $\nabla^{\lambda}G_{\lambda\mu}=0$, and the identity $(\nabla_{\lambda}\nabla_{\mu}-\nabla_{\mu}\nabla_{\lambda})\nabla^{\lambda}\phi^{2}=R_{\mu\lambda}\nabla^{\lambda}\phi^{2}$, we obtain
$\nabla^{\lambda}(\phi^{2}{\cal E}_{\lambda\mu})=6\partial_{\mu}\phi\left(\nabla^{2}-\frac{1}{6}\,R\right)\phi=6\nabla^{\lambda}T^{(m)}_{\lambda\mu}$.
Then, if we substitute the trace of the Einstein EOM (36) in the latter equation, we get the following inhomogeneous continuity equation:
$\nabla^{\lambda}T^{(m)}_{\lambda\mu}=\frac{\partial_{\mu}\phi}{\phi}\,T^{(m)}.$ (39)
The inhomogeneous term on the right-hand side (RHS) of (39) is required for the above equation to be form-invariant under WTs (11). This term represents a fifth force, which can be related to the point-dependent property of the mass parameter in the present theory $m(x)=\kappa\,\phi(x)$ bekenstein_1980 ; hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 . Since timelike fields acquire masses through Yukawa interactions with the Higgs field $H$ after $SU(2)\times U(1)$ symmetry breaking, it is required to modify the standard model of particles (SMP) to get point-dependent masses. That is, the SMP must be Weyl invariant, and the $SU(2)\times U(1)$ symmetry breaking must not spoil Weyl symmetry. Therefore, the Higgs Lagrangian density must be appropriately modified bars-2014 ; mohamedi-2024 . In particular, the mass parameter $v$ in the Higgs potential, $V\propto(H^{\dagger}H-v^{2})^{2}$, must be replaced by a field: $v\rightarrow v_{0}\phi$, where $v_{0}$ is some dimensionless parameter bars-2014 . After the mentioned modification, the timelike fields acquire point-dependent masses through the Yukawa interaction. For example, for a fermion field, the Yukawa coupling is given by $\kappa_{f}\bar{\psi}\phi\psi=\bar{\psi}\,m_{f}\,\psi$ $\Rightarrow$ $m_{f}=\kappa_{f}\,\phi$.
Since $\phi$ is the gravitational coupling in CGR theory, that is, $\phi\sim M_{\text{pl}}$, a mass scale issue could arise. However, we will never be able to measure the point-dependent mass, $m(x)=\kappa\,\phi(x)$, in local experiments, because measuring sticks and clocks are made of atoms (matter) that suffer the same universal point-dependence, so the mass unit also varies: $u_{m}(x)=u_{0}\,\phi(x)$. In experiments, we measure the ratio faraoni_prd_2007
$\frac{m}{u_{m}}=\frac{\kappa}{u_{0}},$
that does not depend on the spacetime point. Hence, even if we assume $\phi_{0}\sim M_{\text{pl}}$, the value $\phi_{0}$ is not what is measured in experiments. In other words, the gigantic mass $m=\kappa\,\phi_{0}$ is compensated for by an equally gigantic mass unit $u_{m}=u_{0}\,\phi_{0}$, so there is no mass scale problem associated with point-dependent masses. The point-dependence of masses (and the related fifth force) can be observed only in “nonlocal” experiments; that is, in experiments that imply emission and absorption of photons coming from distant spacetime points; for example, a photon emitted by some atom located at some point $(t,\vec{x})$ is absorbed by an identical atom at the origin $(0,\vec{0})$. Since the mass of electrons is point-dependent, so is the energy of the emitted/absorbed photon. On the other hand, photon paths are not sensitive to WTs (the Riemannian null-geodesics are already Weyl invariant). This mismatch makes it possible to reveal the point-dependence of the masses, but adequate experiments must be designed (see Section VIII for further discussion of this subject).
IV.1 The fifth force
An interesting property of the CGR theory is related to the inhomogeneous continuity equation (39): $\nabla^{\lambda}T^{(m)}_{\lambda\mu}=T^{(m)}\partial_{\mu}\phi/\phi$. While this equation is satisfied by matter fluxes of stress and energy, the equivalent EOM for point-like particles is the following inhomogeneous EOM for time-like fields:
$\frac{d^{2}x^{\alpha}}{ds^{2}}+\left\{{}^{\alpha}_{\mu\nu}\right\}\frac{dx^{\mu}}{ds}\frac{dx^{\nu}}{ds}=\frac{\partial_{\mu}\phi}{\phi}\,h^{\mu\alpha},$ (40)
where the orthogonal projection tensor is given by
$h^{\mu\nu}\equiv g^{\mu\nu}-\frac{dx^{\mu}}{ds}\frac{dx^{\nu}}{ds}=g^{\mu\nu}+u^{\mu}u^{\nu},$ (41)
with the fourth velocity vector $u^{\mu}=dx^{\mu}/d\tau$ ($d\tau^{2}=-ds^{2}$) satisfying the following conditions: $g_{\mu\nu}u^{\mu}u^{\nu}=-1$ and $h_{\mu\lambda}u^{\lambda}=0$. Fields with vanishing mass, like photons and radiation fields in general, obey an actual conservation equation, $\nabla^{\lambda}T^{\text{rad}}_{\lambda\mu}=0,$ since their SET trace vanishes, $T^{\text{rad}}=g^{\mu\nu}T^{\text{rad}}_{\mu\nu}=0$. This means that these fields follow null geodesics of Riemann space $V_{4}$:
$\frac{dk^{\mu}}{d\xi}+\left\{{}^{\mu}_{\nu\sigma}\right\}k^{\nu}k^{\sigma}=0,$ (42)
where $k^{\mu}\equiv dx^{\mu}/d\xi$ is the wave vector ($k_{\mu}k^{\mu}=0$) and $\xi$ is an affine parameter along null-geodesic. [^9] Note that the null-geodesic equation (42) is form-invariant under CTs. Actually, under (1) the wave vector transforms like $k^{\mu}\rightarrow\Omega^{-2}k^{\mu}$ $\Rightarrow$ $d\xi\rightarrow\Omega^{2}d\xi$, while: $\frac{dk^{\mu}}{d\xi}\rightarrow\Omega^{-4}\left[\frac{dk^{\mu}}{d\xi}-2k^{\mu}\frac{d\ln\Omega}{d\lambda}\right],\;\{^{\alpha}_{\lambda\nu}\}k^{\lambda}k^{\nu}\rightarrow\Omega^{-4}\left[\{^{\mu}_{\lambda\nu}\}k^{\lambda}k^{\nu}+2k^{\mu}\frac{d\ln\Omega}{d\lambda}\right],$ so that (42) is not transformed by (1). Therefore, the fifth force $f^{\alpha}=h^{\alpha\lambda}\partial_{\lambda}\phi/\phi$ is orthogonal to the four-velocity: $u^{\lambda}f_{\lambda}=0$, and it acts on time-like fields exclusively. The fifth force does not interact with null-fields nor with radiation. For that reason, we call it “dark force.” This opens up the possibility that this extra force could explain, in principle, the dark energy and dark matter components of the cosmic fluid.
IV.2 Passive conformal transformations
It has been demonstrated in woodard-1986 (see also hobson-epjc-2022 ) that there is no physics beyond GR in CGR theory. The argument goes like this. To assess the physical consequences of CGR (in woodard-1986 called “conformal scalar-metric theory” in contrast to GR, called “purely metric theory”), we must rely on Weyl-invariant objects and operators. In GR theory, every physical question can be reduced to the VEV of some coordinate-invariant functional of the metric ${\cal F}[g]$. Similarly, every physical question in CGR can be reduced to the VEV of a coordinate and conformally invariant functional of $\phi$ and $g_{\mu\nu}$: $\hat{{\cal F}}[\phi,g]$. In woodard-1986 , it is demonstrated that the VEV of ${\cal F}[g]$ for the theory whose action is $S[g]$, equals the VEV of $\hat{{\cal F}}[\phi,g]$ for the theory whose action is $\hat{S}[\phi,g]$. Hence, any conformal scalar-metric theory is gauge equivalent to the corresponding pure metric theory.
Let us demonstrate the above result in an alternative way and show that the related point of view represents only one of two possible approaches to CTs (28): i) passive Weyl transformations (PWTs), and ii) active Weyl transformations (AWT’s) quiros_prd_2025 . Consider some abstract space of fields ${\cal M}_{f}$ where the metric $g_{\mu\nu}$, the gauge field $\phi$, the masses $m$, and any matter fields $\chi_{i}$ coupled to gravity, are “generalized coordinates.” Each point $P:(\phi,g_{\mu\nu},\chi_{i},m)$ in ${\cal M}_{f}$ represents a gravitational state. Passive CTs
$g_{\mu\nu}\rightarrow g^{\prime}_{\mu\nu}=\Omega^{2}g_{\mu\nu},\;\phi\rightarrow\phi^{\prime}=\Omega^{-1}\phi,\;\chi_{i}\rightarrow\chi^{\prime}_{i}=\Omega^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow m^{\prime}=\Omega^{-1}m,$ (43)
may be viewed as a “rotation” of the coordinate system $R(g_{\mu\nu},\phi,\chi_{i},m)$ $\rightarrow$ $R^{\prime}(g^{\prime}_{\mu\nu},\phi^{\prime},\chi^{\prime}_{i},m^{\prime})$ (see Figure 1 of quiros_prd_2025 where, for simplicity, the vacuum case is considered). The physically meaningful quantities are the invariants:
$\mathfrak{g}_{\mu\nu}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{2}\,g_{\mu\nu},\;\Psi_{i}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{w_{\chi_{i}}}\chi_{i},\;\mathfrak{m}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-1}m.$ (44)
Given that these are the physical (thus measurable) quantities according to PWTs, the fields $g_{\mu\nu}$, $\phi$, $\chi_{i}$, and $m$, which are transformed according to (43) we shall call “auxiliary fields.” The gravitational state is given by ${\cal S}:(\mathfrak{g}_{\mu\nu},\Psi_{i},\mathfrak{m})$. Hence, $R$ and $R^{\prime}$ are two different “coordinate representations” of the same point $P:(\mathfrak{g}_{\mu\nu},\Psi_{i},\mathfrak{m})\in{\cal M}_{f}$; that is, of the same gravitational state $\cal S$}.
It has been demonstrated in quiros_prd_2025 that the Lagrangian density of matter fields (this includes both fundamental matter fields and perfect fluids) is form-invariant under WTs (43). That is, the relationship (27) takes place. Similarly, it can be shown that the following equation:
${\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})={\cal L}_{m}(\Psi_{i},\mathfrak{D}\Psi_{i},\mathfrak{g}_{\mu\nu}),$ (45)
is also satisfied, where $\mathfrak{D}$ is the derivative operator defined in terms of the physical metric $\mathfrak{g}_{\mu\nu}$. This equality relates the Lagrangian density of the auxiliary matter fields $\chi_{i}$ to that of the physical matter fields $\Psi_{i}$. In the Appendix A, for illustrative purposes, we show equality (45) for the Lagrangian density of the Dirac fermion and for that of perfect fluids. For other fundamental fields, e.g., gauge bosons, the demonstration is similar.
In view of equality (45), the equation (37) can be written as
${\cal L}_{\text{tot}}=\frac{\sqrt{-g}}{2}\left[\frac{1}{6}\,\phi^{2}R+(\partial\phi)^{2}\right]+{\cal L}_{m}(\Psi_{i},\mathfrak{D}\Psi_{i},\mathfrak{g}_{\mu\nu}).$ (46)
The next step is to write the gravitational part of this Lagrangian density in terms of the physical metric, just as we have done for the matter Lagrangian density. We get
${\cal L}_{\text{tot}}=\sqrt{-\mathfrak{g}}\,\frac{M^{2}_{\text{pl}}}{2}\mathfrak{R}+{\cal L}_{m}(\Psi_{i},\mathfrak{D}\Psi_{i},\mathfrak{g}_{\mu\nu}),$ (47)
where $\mathfrak{R}$ is the curvature scalar defined in terms of the Weyl invariant, thus physical, metric. Written in this (covariant) way, the total Lagrangian density coincides with that of GR. The difference between (37) and (47) according to the PWTs is that the latter Lagrangian density is written in terms of physical quantities, while the former one is written in terms of auxiliary quantities without direct physical meaning.
We notice that the auxiliary fields have been removed from the covariant formulation (47). Hence, since the WTs act precisely on the auxiliary fields, we can say that the passive approach to CTs is not suitable to search for the physical and phenomenological consequences of Weyl symmetry. In this case, as demonstrated in woodard-1986 and confirmed in hobson-epjc-2022 , there is no physics beyond GR in CGR theory. However, as we have already commented, this is only one of two possible understandings of WTs (28).
IV.3 Active transformations
[…]
$g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu},\;\phi\rightarrow\hat{\phi}=\Omega^{-1}\phi,\;\chi_{i}\rightarrow\hat{\chi}_{i}=\Omega^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow\hat{m}=\Omega^{-1}m,$ (48)
[…]
${\cal L}_{\text{tot}}=\frac{\sqrt{-g}}{2}\left[\frac{1}{6}\,\phi^{2}R+(\partial\phi)^{2}\right]+{\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu}),$
is already written covariantly in terms of physical fields $g_{\mu\nu}$, $\phi$, $\chi_{i}$, and $m$. In this case, Weyl invariants, such as those defined in (44), represent useful relationships between the physical fields in the two different gravitational states:
$\mathfrak{g}_{\mu\nu}$ $=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{2}\,g_{\mu\nu}=\left(\frac{\hat{\phi}}{\sqrt{6}M_{\text{pl}}}\right)^{2}\,\hat{g}_{\mu\nu}\;\Rightarrow\;g_{\mu\nu}=\left(\frac{\hat{\phi}}{\phi}\right)^{2}\,\hat{g}_{\mu\nu},$
$\Psi_{i}$ $=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{w_{\chi_{i}}}\chi_{i}=\left(\frac{\hat{\phi}}{\sqrt{6}M_{\text{pl}}}\right)^{w_{\chi_{i}}}\hat{\chi}_{i}\;\Rightarrow\;\chi_{i}=\left(\frac{\hat{\phi}}{\phi}\right)^{w_{\chi_{i}}}\hat{\chi}_{i},$
$\mathfrak{m}$ $=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-1}m=\left(\frac{\hat{\phi}}{\sqrt{6}M_{\text{pl}}}\right)^{-1}\hat{m}\;\Rightarrow\;m=\left(\frac{\hat{\phi}}{\phi}\right)^{-1}\hat{m}$. (49)
From now on, when we talk about Weyl transformations, we mean the active approach.
V Gauge freedom: The “many-worlds” interpretation
The freedom to choose any arbitrary function $\phi=\phi(x)$ is called “gauge freedom”. Different gauges are fixed by different choices of the scalar $\phi=\phi(x)$. For example, the GR gauge is achieved by choosing $\hat{\phi}=\sqrt{6}\,M_{\text{pl}}=$ const. Under this choice, the total Lagrangian density (37) is written as;
${\cal L}_{\text{tot}}=\sqrt{-\hat{g}}\,\frac{M^{2}_{\text{pl}}}{2}\,\hat{R}+{\cal L}_{m}(\hat{\chi}_{i},\hat{\partial}\hat{\chi}_{i},\hat{g}_{\mu\nu}),$ (50)
while the Einstein EOM (38) reads
$\hat{G}_{\mu\nu}=\frac{1}{M^{2}_{\text{pl}}}\,\hat{T}^{(m)}_{\mu\nu}$, (51)
which is the standard GR Einstein EOM.
The GR gauge is related to any other (arbitrary) gauge of (37) through the following equations (see (49)):
$g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{2}g_{\mu\nu},\;\chi_{i}\rightarrow\hat{\chi}_{i}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow\hat{m}=\left(\frac{\sqrt{6}M_{\text{pl}}}{\phi}\right)\,m,$
where $\hat{m}$ is a constant mass parameter and we took into account $\phi\rightarrow\hat{\phi}=\Omega^{-1}\phi=\sqrt{6}M_{\text{pl}}$, from which we get $\Omega=\phi/\sqrt{6}M_{\text{pl}}$. Here, we also consider the transformation $T^{(m)}_{\mu\nu}\rightarrow\hat{T}^{(m)}_{\mu\nu}=\Omega^{-2}\,T^{(m)}_{\mu\nu}$ of the matter SET; that is, $\hat{T}^{(m)}_{\mu\nu}=(\phi/\sqrt{6}M_{\text{pl}})^{-2}\,T^{(m)}_{\mu\nu}$. Since GR is a gauge of Weyl invariant theory (37), we refer to the latter theory as CGR.
Note that in an arbitrary gauge, where ${\cal L}_{m}={\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$, the matter fields $\chi_{i}$ are minimally coupled to the metric $g_{\mu\nu}$. Therefore, it is the metric that defines the proper time intervals; that is, it is the physical metric in an arbitrary gauge. In the same way, the matter fields $\hat{\chi}_{i}$ are minimally coupled to the metric $\hat{g}_{\mu\nu}$, so it is the physical metric in the GR gauge. The GR gauge $\mathfrak{G}_{\text{gr}}=\left\{\sqrt{6}M_{\text{pl}},\hat{g}_{\mu\nu}(x)\right\}$, where $\hat{g}_{\mu\nu}=\hat{g}_{\mu\nu}(x)$ is any solution of the GR-EOM (51), and any other arbitrary gauge $\mathfrak{G}=\{\phi(x),g_{\mu\nu}(x)\}$, where $g_{\mu\nu}=g_{\mu\nu}(x)$ is a solution of the Einstein-type EOM (38) once we have fixed the function $\phi=\phi(x)$, represent potential descriptions of the conformal invariant gravitational interactions of matter according to CGR theory. Both solutions are related by
$g_{\mu\nu}=\left(\frac{\sqrt{6}M_{\text{pl}}}{\phi}\right)^{2}\hat{g}_{\mu\nu}.$ (52)
Hence, we do not need to solve the Einstein-type EOM (38) after making a specific choice of $\phi=\phi(x)$. We only need to know some GR solution $\hat{g}_{\mu\nu}(x)$, then, if we fix the function $\phi=\phi(x)$, the corresponding conformal metric, according to (52), is necessarily a solution of (38).
The usefulness of this result is illustrated, for example, in a series of papers bambi-2017-a ; bambi-2017-b ; bambi-2017-c ; bambi-2018-a ; bambi-2018-b , where conformal invariance is used in several contexts, for example, to look for the properties of some astrophysical black holes, to investigate the black hole evaporation, as well as the singularity issues. Note that in these bibliographic references, each conformal solution is assumed to be physically meaningful. However, this is true only if we follow the active approach to Weyl transformations. [^10] In contrast, if we follow the passive approach, the physical metric is the conformal invariant quantity $\mathfrak{g}_{\mu\nu}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{2}\,g_{\mu\nu}=\hat{g}_{\mu\nu}.$ That is, the GR metric $\hat{g}_{\mu\nu}$ coincides with the physical metric. Therefore, only the GR solution $\hat{g}_{\mu\nu}(x)$ matters. In this case, both the GR gauge $\mathfrak{G}_{\text{gr}}$ and any arbitrary gauge $\mathfrak{G}_{k}$;
$\mathfrak{G}_{k}=\{\phi_{k}(x),g^{(k)}_{\mu\nu}(x)\},\;g^{(k)}_{\mu\nu}=\left(\frac{\sqrt{6}M_{\text{pl}}}{\phi_{k}}\right)^{2}\hat{g}_{\mu\nu},\;k=1,2,...,N,$
in the conformal equivalence class $\mathfrak{C}_{\text{conf}}=\{\mathfrak{G}_{\text{gr}},\mathfrak{G}_{1},\mathfrak{G}_{2},...,\mathfrak{G}_{N}\}$ of (38), represent different gravitational states that account for different phenomenological consequences. The choice of one gauge or another can only be an experimental matter. This may be understood as a classical version of the many-worlds interpretation of quantum mechanics everett_rmp_1957 ; wheeler ; dewitt_phys_rev_1967 ; dewitt ; barvinsky_cqg_1990 ; omnes_rmp_1992 ; tegmark_1998 ; garriga_prd_2001 ; zurek_rmp_2003 ; tegmark_nature_2007 ; quiros_prd_2023 applied to the gravitational interactions of matter.
V.1 Vacuum solution: Conformal Schwarzschild space
In a series of papers bambi-2017-a ; bambi-2017-b ; bambi-2017-c ; bambi-2018-a ; bambi-2018-b , conformal invariance is used to investigate the properties of astrophysical black holes and to examine black hole evaporation and singularity issues. In these bibliographic references, each conformal solution is assumed to be physically meaningful; that is, the authors implicitly assume the active approach to WTs. Imagine, for example, that in the GR gauge we have the Schwarzschild GR vacuum solution:
$d\hat{s}^{2}=\hat{g}_{\mu\nu}dx^{\mu}dx^{\nu}=-\left(1-\frac{2m}{r}\right)dt^{2}+\frac{dr^{2}}{\left(1-\frac{2m}{r}\right)}+r^{2}d\Omega^{2},$ (53)
where $d\Omega^{2}\equiv d\theta^{2}+sin^{2}\theta\,d\varphi^{2}$. Then, according to (52), the conformal metric:
$ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}=\left[\frac{\sqrt{6}M_{\text{pl}}}{\phi(r)}\right]^{2}\left[-\left(1-\frac{2m}{r}\right)dt^{2}+\frac{dr^{2}}{\left(1-\frac{2m}{r}\right)}+r^{2}d\Omega^{2}\right],$ (54)
is also a solution of the vacuum EOM [^11] The solutions (53) and their conformal (54), are valid in the absence of a cosmological term. Otherwise, we must replace (53) with the Schwarzschild-de Sitter solution instead.
${\cal E}_{\mu\nu}\equiv G_{\mu\nu}+\frac{6}{\phi^{2}}\left[\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}g_{\mu\nu}(\partial\phi)^{2}\right]-\frac{1}{\phi^{2}}\left(\nabla_{\mu}\nabla_{\nu}-g_{\mu\nu}\nabla^{2}\right)\phi^{2}=0.$ (55)
As an illustration of the meaning of gauge fixing in CGR within the AWTs framework, in the bibliographic reference bambi-2017-a , the authors made the following gauge choice: [^12] In this and other bibliographic references, the BD scalar field $\Phi=\phi^{2}/6$ is used instead of $\phi$, so certain minimal differences may appear in the resulting formulae with respect to the present document.
$\phi(r)=\frac{\sqrt{6}M_{\text{pl}}}{\sqrt{1+(L/r)^{4}}}=\frac{\sqrt{6}M_{\text{pl}}\,r^{2}}{\sqrt{r^{4}+L^{4}}},$ (56)
where $L$ is some length scale. It is shown in bambi-2017-a that the spacetime given by the metric (54), with $\phi(r)$ given by (56), is complete, so the resulting geometry represents a nonsingular black hole. The authors conclude that the singularity issue is an artifact of the GR gauge. According to bambi-2017-a , another possible gauge choice, which is suitable for removing the spacetime singularity, could be
$\phi(r)=\frac{\sqrt{6}M_{\text{pl}}}{1+(L/r)^{2}}=\frac{\sqrt{6}M_{\text{pl}}\,r^{2}}{r^{2}+L^{2}}.$ (57)
In these examples, the Schwarzschild vacuum solution $d\hat{s}^{2}$ given by (54) and its conformal singularity-free solutions; $\sqrt{6}M_{\text{pl}}/\phi(r)]^{2}\,d\hat{s}^{2}$, where $\phi(r)$ can be given by (56) or (57) (or possibly by many other $r$-functions), give suitable descriptions of vacuum CGR (55). The fact that gravitational states with a spacetime singularity are in the same equivalence class as singularity-free spacetimes is very interesting because this might shed some light on the singularity issue, as done, for example, in bambi-2017-a ; bambi-2017-b ; bambi-2017-c ; bambi-2018-a ; bambi-2018-b .
If instead of the AWTs we adopted the passive approach to WTs, the results in these bibliographic references would not have any physical significance. In fact, according to passive WTs, the physically meaningful metric $\mathfrak{g}_{\mu\nu}=[\phi(r)/\sqrt{6}M_{\text{pl}}]^{2}\,g_{\mu\nu}=\hat{g}_{\mu\nu}$, coincides with the GR vacuum metric. That is, only the Schwarzschild black hole solution (53) matters.
V.2 Cautionary note on gauge freedom
The interpretation of gauge freedom (also called gauge fixing) we have applied within the framework of CGR theory, as well as its physical meaning, has nothing in common with the usual interpretation of gauge freedom in standard gauge theories of fields. Consider, for example, the Lagrangian of a Maxwell field $A_{\mu}$ coupled to a Dirac fermion $\psi$:
${\cal L}_{\text{em}}=-\frac{1}{4}\,F_{\mu\nu}F^{\mu\nu}+\bar{\psi}(i\cancel{D}-m)\psi,$ (58)
where $F_{\mu\nu}\equiv\partial_{[\mu}A_{\nu]}/2$ is the Maxwell field strength, $\psi$ is the Dirac spinor of the fermion field ($\bar{\psi}$ is its adjoint), $\gamma^{\mu}$ are the Dirac gamma matrices, and $\cancel{D}\psi\equiv\gamma^{\mu}D_{\mu}\psi=\gamma^{\mu}(\partial_{\mu}\psi+iA_{\mu}\psi)$ is the gauge covariant derivative. It is assumed that the background space is Minkowski space. The Lagrangian (58) is form-invariant under gauge transformations:
$A_{\mu}\rightarrow\hat{A}_{\mu}=A_{\mu}+\partial_{\mu}\lambda(x),\;\psi\rightarrow\hat{\psi}=e^{-i\lambda(x)}\psi,\;\bar{\psi}\rightarrow\hat{\bar{\psi}}=\bar{\psi}\,e^{i\lambda(x)},$
where the free function $\lambda(x)$ is the gauge parameter.
In this case, gauge invariance of (58) means, in turn, that the measured quantities: the Lorentz scalar type $\bar{\psi}\psi$, vector type $\bar{\psi}\gamma^{\mu}\psi$, and tensor type $\bar{\psi}\gamma^{\mu}\gamma^{\nu}\psi$ quantities, as well as the time and spatial components of the gauge invariant field strength $F_{\mu\nu}$; the electric $E_{i}=F_{0i}$ and magnetic $B_{i}=-\epsilon_{ijk}F^{jk}/2$ fields ($\epsilon_{ijk}$ is the Levi-Civita tensor), respectively, are not affected by the gauge transformations, so $\bar{\psi}\psi=\hat{\bar{\psi}}\hat{\psi}$, $\bar{\psi}\gamma^{\mu}\psi=\hat{\bar{\psi}}\gamma^{\mu}\hat{\psi}$, …, $E_{i}=\hat{E}_{i}$ and $B_{i}=\hat{B}_{i}$. This means that these transformations do not modify the state of the Maxwell field in interaction with the fermion. This is in contrast with the interpretation of conformal transformations of the physically meaningful quantities; that is, the metric, the scalar field, and the other fields coupled to gravity, according to the active approach: $g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu}$, $\phi\rightarrow\hat{\phi}=\Omega^{-2}\phi$, …, where $\{\phi(x),g_{\mu\nu}(x),$...\} and $\{\hat{\phi}(x),\hat{g}_{\mu\nu}(x),$...\} represent different gravitational states, which are characterized by different curvature properties, proper time lapse, etc.
In this document, we call gauge fixing and gauge freedom the specific choice of the free scalar $\phi(x)$, as explained above, because, historically, Weyl called conformal transformations (1) as “gauge transformations”.
VI Galactic rotation curves
In harko-prd-2023 , the possibility that the observed behavior of test particles outside galaxies, which is usually explained with the help of dark matter, could be explained, alternatively, within the framework of so-called conformal invariant Weyl geometric quadratic gravity ghilen-2019 , was investigated. The static spherically symmetric solution found in harko-epjc-2022 to the equations of motion of the latter theory was one of three cases studied in harko-prd-2023 , showing a good fit of the galactic rotation curves.
For the specific form of the Weyl vector with only a radial nonvanishing component, chosen in harko-prd-2023 ; harko-epjc-2022 , the resulting vacuum EOMs coincide exactly with (55) with the addition of a symmetry-preserving potential $V_{0}\phi^{4}/24$, for the specific choice $V_{0}=1$, once the static spherically symmetric metric (in harko-prd-2023 ; harko-epjc-2022 the BD field $\Phi=\phi^{2}$ was used):
$ds^{2}=-e^{\nu(r)}dt^{2}+e^{\lambda(r)}dr^{2}+r^{2}d\Omega^{2},$ (59)
is substituted. The solution studied in harko-prd-2023 reads:
$e^{\nu}=e^{-\lambda}=\alpha-2+\frac{(\alpha-3)C_{2}}{3r}+(\alpha-1)\frac{r}{C_{2}}+C_{3}r^{2},$ (60)
which corresponds to the gauge choice:
$\phi(r)=\frac{\sqrt{C_{1}}M_{\text{pl}}}{r+C_{2}}.$ (61)
In these equations $C_{1}$, $C_{2}$ and $C_{3}$ are integration constants, while $\alpha\equiv 3C_{3}C^{2}_{2}+C_{1}/4$. It is not difficult to show that, under the coordinate transformation,
$\tilde{t}=\sqrt{\frac{C_{1}}{6C^{2}_{2}}}\,t,\;\tilde{r}=\sqrt{\frac{C_{1}}{6}}\frac{r}{r+C_{2}},$
the above solution is written as conformal Schwarzschild-de Sitter (SdS) solution: [^13] The study of conformal SdS geometry to explain the galactic rotation curves has been undertaken also in modesto-2020 ; modesto-2024 .
$ds^{2}=\left[\frac{\sqrt{6}M_{\text{pl}}}{\phi(\tilde{r})}\right]^{2}d\hat{s}^{2},$ (62)
where
$d\hat{s}^{2}=-\left(1-\frac{2m}{\tilde{r}}-\frac{\Lambda_{0}}{3}\,\tilde{r}^{2}\right)\,d\tilde{t}^{2}+\left(1-\frac{2m}{\tilde{r}}-\frac{\Lambda_{0}}{3}\,\tilde{r}^{2}\right)^{-1}d\tilde{r}^{2}+\tilde{r}^{2}d\Omega^{2},$ (63)
is the GR-SdS solution and the choice of the scalar field (61) in the new coordinates reads as
[…]
In these equations, we have defined the following constants: $2m\equiv\sqrt{C_{1}/6}(1-C_{3}C^{2}_{2}-C_{1}/12)$ and $\Lambda_{0}\equiv C_{1}/4$.
Therefore, according to active CTs, the GR Schwarzschild-de Sitter metric (63), which is not suitable for explaining the galactic rotation curves, and the conformal solution (62), which can provide a good background for explaining the rotation curves, belong to the same conformal equivalence class. Both are equally plausible solutions of vacuum EOMs (55) that represent different gravitational states according to AWTs. Only experiment (in the present case, the observational data on galactic rotation curves) can choose a specific gauge among the infinity of conformally related gauges in the conformal class.
If we look at the above problem from the point of view of the passive approach to CTs, as done, for instance, in hobson-epjc-2022 , there are no physical consequences of Weyl symmetry. Actually, in this bibliographic reference, it is stated that if one works in terms of quantities that can be physically measured, within the framework of conformally invariant gravity theories defined on Riemannian spacetime and having the SdS metric as a solution in the Einstein gauge, in any conformal frame, the trajectories followed by matter particles are merely the timelike geodesics of the SdS metric. Hence, there is no frame dependence of physical predictions, and it is clear that the rotation curves have no flat region. It is also stated in hobson-epjc-2022 that, since CGR is (by construction) conformally invariant, conformal transformations should not change the “observable predictions”. This makes clear that the authors consider that only conformal invariant quantities, in particular the conformal invariant metric $\mathfrak{g}_{\mu\nu}\equiv(\phi/\sqrt{6}M_{\text{pl}})^{2}g_{\mu\nu}$, have observational meaning. However, since the Weyl invariant metric $\mathfrak{g}_{\mu\nu}=(\phi/\sqrt{6}M_{\text{pl}})^{2}g_{\mu\nu}=\hat{g}_{\mu\nu}$, only the GR-SdS geometry (63) has an observational meaning. Yet, we must recall that this analysis is correct only if one assumes from the start that the passive Weyl transformations represent the only correct approach to conformal symmetry. As we have shown, if we follow this approach, the action and derived EOM of CGR theory, when written in terms of the physical quantities, coincide with the GR action and derived EOM, so Weyl symmetry is lost or is, at least, a fictitious symmetry. Recall that in this case the “auxiliary” fields that suffer the Weyl transformations: $g_{\mu\nu}$, $\phi$, $\chi_{i}$, and $m$ are removed from the covariant formulation (47).
VII Weyl invariant cosmology and the many-worlds perspective
The CGR theory (37) is a much more versatile alternative to solving current gravitational puzzles than GR alone, since the latter belongs to the class of conformal equivalence of CGR. Suppose that we have an GR solution $\hat{g}_{\mu\nu}(x)$ and that a gravitational state is labeled “kk”;
${\cal S}_{k}:\left(\phi_{k}(x),g^{(k)}_{\mu\nu}(x),...\right),\;\;k=1,2,...\,,N,$ (64)
where $N\rightarrow\infty$ and the ellipsis stands for other fields coupled to gravity, is such that the fields $\phi_{k}(x)$ and $g^{(k)}_{\mu\nu}(x)$ satisfy the CGR-EOM (38) for any $k$ and any gauge choice $\phi_{k}=\phi_{k}(x)$. The different metric tensors $g^{(k)}_{\mu\nu}(x)$ are related to the GR solution through (52);
$g^{(k)}_{\mu\nu}(x)=\left(\frac{\sqrt{6}M_{\text{pl}}}{\phi_{k}(x)}\right)^{2}\hat{g}_{\mu\nu}(x).$
Hence, instead of a single GR gravitational state ${\cal S}_{\text{gr}}$, in CGR there is an infinite set of gravitational states: $\mathfrak{S}_{\text{cgr}}=\{{\cal S}_{\text{gr}},{\cal S}_{1},{\cal S}_{2},...,{\cal S}_{N}\}$, including the GR gauge itself, which are available to describe a given set of gravitational phenomena.
Let us illustrate the versatility of CGR in the cosmological framework. Here, we assume a Friedmann-Robertson-Walker (FRW) line-element with flat spatial sections:
$ds^{2}=-dt^{2}+a^{2}(t)\delta_{ij}dx^{i}dx^{j},$ (65)
where $t$ is cosmic time, $a(t)$ is the scale factor, and the latin indices run through space: $i,j=1,2,3$. We further assume that the background matter content is a perfect fluid with energy density $\rho$ and pressure $p$ that satisfy the following EOS: $p=(\gamma-1)\rho$, where $\gamma$ is the barotropic index of the fluid. The independent equations are the $(0,0)$-component of the Einstein-type EOM (38) and the nonhomogeneous continuity equation (39);
$\left(\frac{a^{\prime}}{a}+\frac{\phi^{\prime}}{\phi}\right)^{2}$ $=\frac{2a^{2}}{\phi^{2}}\,\rho,$ (66)
$\rho^{\prime}+3\gamma\frac{a^{\prime}}{a}\,\rho$ $=(4-3\gamma)\frac{\phi^{\prime}}{\phi}\,\rho,$ (67)
respectively, where the tilde accounts for the derivative with respect to the conformal time $\tau=\int dt/a(t)$. [^14] We use the conformal time instead of the cosmic time $t$ because the former is a Weyl-invariant quantity, while $dt\rightarrow d\hat{t}=\Omega\,dt$ under WTs. Straightforward integration of (67) yields $\rho=(C^{2}/2)\,\phi^{4}(a\phi)^{-3\gamma}$ where $C^{2}$ is an integration constant. Taking into account this equation, if we introduce the Weyl-invariant variable $v\equiv a\phi$, the Friedmann EOM (66) can be written as
$v^{\prime}=C\,v^{\frac{4-3\gamma}{2}}.$ (68)
This is a Weyl-invariant equation whose integration yields $v(\tau)=\bar{C}\,(\tau-\tau_{0})^{\frac{2}{3\gamma-2}}$, where $\tau_{0}$ is another integration constant and $\bar{C}\equiv[(3\gamma-2)\,C/2]^{2/(3\gamma-2)}$. We have
$a(\tau)=\bar{C}\,(\tau-\tau_{0})^{\frac{2}{3\gamma-2}}\phi^{-1}(\tau).$ (69)
Different choices of the gauge field: $\phi(\tau)=\phi_{1}(\tau)$, $\phi(\tau)=\phi_{2}(\tau)$$\phi(\tau)=\phi_{N}(\tau)$ ($N\rightarrow\infty$), lead to different behaviors of the scale factor: $a_{1}(\tau)$, $a_{2}(\tau)$,…, $a_{N}(\tau)$. In the GR gauge, we have: $\phi(\tau)=\phi_{\text{gr}}=\sqrt{6}M_{\text{pl}}$, so
$a_{\text{gr}}(\tau)=\frac{\bar{C}}{\sqrt{6}M_{\text{pl}}}\,(\tau-\tau_{0})^{\frac{2}{3\gamma-2}}.$
In CGR theory, instead of a single gravitational state for each value of the barotropic index $\gamma$, as in GR, we have an infinite class of gravitational states:
$\mathfrak{S}_{\text{cgr}}=\left\{\left(\sqrt{6}M_{\text{pl}},a_{\text{gr}}(\tau)\right),\left(\phi_{1}(\tau),a_{1}(\tau)\right),...,(\phi_{N}(\tau),a_{N}(\tau))\right\}.$
These are related by the following Weyl transformation:
$a_{k}(\tau)=\left[\frac{\sqrt{6}M_{\text{pl}}}{\phi_{k}(\tau)}\right]a_{\text{gr}}(\tau),\;\;k=1,2,...,N.$
The large (in principle infinite) variety of cosmic behaviors in $\mathfrak{S}_{\text{cgr}}$ is what makes the CGR theory and, therefore, Weyl symmetry attractive within the cosmological context and, at the same time, yields the most unappealing feature of this theory when it is analyzed under the active approach: the loss of predictability. The Weyl invariant CGR theory is not capable of defining a unique gravitational state, but a whole class of them, so that the definition of the state that correctly describes a given gravitational phenomenon is left to experiment. Nevertheless, this theory can be assigned some predictive power. For example, the fifth force $f^{\mu}=h^{\mu\lambda}\partial_{\lambda}\phi/\phi$ in (40), which is associated with point-dependent masses so it can pass unnoticed in local experiments, points to a modification of the redshift measurements that could explain, in principle, the present accelerated pace of the cosmic expansion and so-called dark matter, because it does not interact with radiation ($T^{\text{rad}}=0$ $\Rightarrow$ $f_{\mu}=0$). In the last instance, the active approach to Weyl symmetry within CGR theory may be viewed as a classical version of the many-worlds interpretation of gravitational interactions of matter everett_rmp_1957 ; wheeler ; dewitt_phys_rev_1967 ; dewitt ; barvinsky_cqg_1990 ; omnes_rmp_1992 ; tegmark_1998 ; garriga_prd_2001 ; zurek_rmp_2003 ; tegmark_nature_2007 ; quiros_prd_2023 .
There is another issue that warrants discussion. To introduce the problem, we shall follow the line of reasoning of faraoni_prd_2007 . Consider the line-element in the GR gauge and in any other $k$-gauge: $ds^{2}_{\text{gr}}=a^{2}_{\text{gr}}(\tau)(-d\tau^{2}+\delta_{ij}dx^{i}dx^{j})$ and $ds^{2}_{k}=a^{2}_{k}(\tau)(-d\tau^{2}+\delta_{ij}dx^{i}dx^{j})$, respectively. According to the bibliographic reference mentioned above, the physically meaningful quantity to analyze, for instance, the singularity issue, is not the scale factor itself but a typical physical (proper) length $a(\tau)\sqrt{\delta_{ij}dx^{i}dx^{j}}$. The corresponding measured quantity would be the ratio $a(\tau)\sqrt{\delta_{ij}dx^{i}dx^{j}}/u(\tau)$, where $u$ is the unit of length. Since $u$ transforms in the same way as the proper length, then:
$\frac{a_{k}(\tau)\sqrt{\delta_{ij}dx^{i}dx^{j}}}{u_{k}(\tau)}=\frac{a_{\text{gr}}(\tau)\sqrt{\delta_{ij}dx^{i}dx^{j}}}{u_{\text{gr}}}.$ (70)
This means that there is no measurable effect that the gauge choice could have. However, there are some flaws in this line of reasoning. For example, since here we follow the active approach to Weyl transformations (the only possible approach that leads to Weyl symmetry having phenomenological consequences quiros_prd_2025 ), the GR metric $g^{\text{gr}}_{\mu\nu}(\tau)$ and any of its conformal metrics $g^{(k)}_{\mu\nu}(\tau)$ have independent physical meaning. In particular, the curvature tensors and any of their contractions are made up of the derivatives of these metric tensors. In the present cosmological example, the curvature scalar in the $k$-gauges reads: $R_{k}=6(\dot{H}_{k}+2H^{2}_{k})$, while in the GR gauge: $R_{\text{gr}}=6(\dot{H}_{\text{gr}}+2H^{2}_{\text{gr}})$, where the Hubble parameter is defined as $H_{k}=\dot{a}_{k}/a_{k}$, $H_{\text{gr}}=\dot{a}_{\text{gr}}/a_{\text{gr}}$, etc. Here, the dot accounts for the derivative with respect to cosmic time $t$, so $dt=ad\tau$. Under WTs, the Hubble parameter transforms as $H\rightarrow H_{\text{gr}}=\Omega^{-1}(H+\dot{\phi}/\phi)$. The transformation of $\dot{H}$ is even more complex. So is the transformation of the curvature scalar. Hence, the arguments of faraoni_prd_2007 do not apply when the derivatives of the scale factor are implied.
Additionally, many measurements in cosmology rest on the emission and detection of photons. These follow null-geodesics of Riemann geometry, which are themselves form-invariant under WTs. Any measuring clock or stick, which is made of a combination of emitted and/or absorbed photons, is not transformed by the Weyl transformations and so is the related measuring unit. Hence, the above argument faraoni_prd_2007 is not universal and can be circumvented. In addition, even if a measuring unit, for example, the mass unit, is point-dependent, there are ways in which this effect can be experimentally revealed. For example, the point-dependent property of the masses of timelike fields, and so the fifth force appearing in (39), cannot be detected in local experiments because any measuring stick is made up of fields, so it suffers the same universal point-dependence as the mass being measured, and the effect cancels out. However, in experiments that involve the detection of photons emitted at a distant point in spacetime, namely redshift measurements, the effect can be measured (at least in principle).
VIII Observational consequences of CGR: The redshift measurements
Let us explain the peculiar contribution to the redshift of frequencies carried by the Weyl symmetry. In CGR theory, two phenomena contribute to the redshift of frequencies: 1) the gravitational redshift (also curvature redshift) and 2) the redshift due to variation of masses in spacetime (mass redshift). The former effect is due to the propagation of photons in a curved background, while the latter is due to the point-dependent property of masses: $m=m(x)=\mu\,\phi(x)$. This mass dependence affects the redshift because the frequency of a photon emitted/absorbed by some atom (atomic number $Z$) is given by the Rydberg formula:
$\gamma_{e/a}=Z^{2}\frac{m_{e}c^{2}\alpha^{2}}{2h}\left(\frac{1}{n^{2}_{l}}-\frac{1}{n^{2}_{u}}\right),$ (71)
where $m_{e}$ is the mass of the electron, $c$ is the speed of light, $h$ is the Planck constant (here, temporarily we assume that $c\neq h\neq 1$), $\alpha$ is the fine-structure constant ($\alpha\approx 1/137$), while $n_{l}$ and $n_{u}$ are the principal quantum numbers of the lower and upper energy levels, respectively. Suppose that two identical atoms are located at spacetime points $x$ and $x_{0}$, respectively. The relative shift of the frequency of the emitted/absorbed photons is given by
$z_{m}=\frac{\gamma_{e/a}(x)-\gamma_{e/a}(x_{0})}{\gamma_{e/a}(x_{0})}=\frac{m_{e}(x)-m_{e}(x_{0})}{m_{e}(x_{0})}=\frac{\phi(x)}{\phi(x_{0})}-1.$ (72)
Note that this emission/absorption frequency does not depend on the specific atom that emits/absorbs the photon, nor on the specific atomic transition $n_{u}\rightleftharpoons n_{l}$. Note also that in the GR gauge, since $\phi(x)=\phi(x_{0})=M_{\text{pl}}$, this type of redshift does not arise. Meanwhile, the curvature redshift is due to the dynamics of the photon frequency during its propagation in a curved background. The frequency of a photon emitted/absorbed at some point $x$ in spacetime can be found as a solution of the null-geodesics (42): $\gamma_{e/a}(x)=\gamma^{0}_{e/a}/\sqrt{-g_{00}(x)}$, where $\gamma^{0}_{e/a}$ is some constant frequency. The curvature redshift resulting from the propagation of a photon from point $x$ to $x_{0}$ and the overall redshift are given by:
$z_{c}=\frac{\gamma_{e/a}(x)-\gamma_{e/a}(x_{0})}{\gamma_{e/a}(x_{0})}=\sqrt{\frac{g_{00}(x_{0})}{g_{00}(x)}}-1,$ (73)
and
$z_{\text{tot}}=z_{c}+z_{m}=\sqrt{\frac{g_{00}(x_{0})}{g_{00}(x)}}+\frac{\phi(x)}{\phi(x_{0})}-2,$ (74)
respectively. If we take into account the relationship (52) between the metric $g_{\mu\nu}(x)$ of an arbitrary gauge and the metric $\hat{g}_{\mu\nu}(x)$ in the GR gauge, we get,
$z_{\text{tot}}=\frac{\phi(x)}{\phi(x_{0})}\sqrt{\frac{\hat{g}_{00}(x_{0})}{\hat{g}_{00}(x)}}+\frac{\phi(x)}{\phi(x_{0})}-2,$
or if we realize that the GR redshift $z=\sqrt{\hat{g}_{00}(x_{0})/\hat{g}_{00}(x)}-1$, then
$z_{\text{tot}}=\frac{\phi(x)}{\phi(x_{0})}\,z+2\left(\frac{\phi(x)}{\phi(x_{0})}-1\right).$ (75)
In a cosmological context, if we replace $t\rightarrow x$, where $t$ is cosmic time, and set $x_{0}=0$ (present time), under the assumption of very small $\phi$-variations, we can write
$\frac{\phi(t)}{\phi(0)}=1+\frac{\dot{\phi}(0)}{\phi(0)}\Delta t=1+\epsilon,$ (76)
where a dot means derivative with respect to cosmic time, we assume that $\phi(0)=\sqrt{6}M_{\text{pl}}$, and that $\epsilon>0$ is a small positive number, so that
$z_{\text{tot}}=(1+\epsilon)z+2\epsilon.$ (77)
This simplified expression shows that, in CGR cosmology, the measured redshift $z_{\text{tot}}$ is greater than the redshift according to GR. Although (77) is an approximate equation, for redshifts $\sim 1$ and a bit larger, this result can have obvious implications for the explanation of the late-time dynamics of our Universe. Since the fifth force that arises in (39) and in (40) is due to the point-dependent property of the mass: $m(x)=\kappa\,\phi(x)$, accurate measurements of the second term in the RHS of (77) would entail indirect measurements of the extra force $f^{\alpha}=h^{\alpha\lambda}\partial_{\lambda}\phi/\phi$.
To illustrate the modification of the frequency redshift achieved by the assumption of CGR theory under the active approach to Weyl transformations, let us consider the Pound-Rebka and Pound-Snider experiments pound-1960 ; pound-1964 ; pound-1965 . A target $T_{h}$ made of some atom is placed at some height $h$ on top of a tower on the Earth’s surface (radial coordinate $r=R_{\oplus}+h$, where $R_{\oplus}$ is the mean radius of the Earth) and an identical target $T_{0}$ (made of the same atom) is placed at the bottom of the tower (radial coordinate $r_{0}=R_{\oplus}$.) Atoms in $T_{h}$ are in some excited state, so these can emit photons to return to the ground state. The emitted photons are then absorbed by atoms in the target $T_{0}$, which is located at the bottom of the tower. [^15] We omit experimental details such as, for instance, that the target at the bottom of the tower must be mounted in a centrifuge with controlled rotation speed, or that the temperature of the emitter had to be controlled to high precision, or that the role of the emitter and of the absorber was continuously exchanged, etc. If we expand (75) $\phi(r)=\phi(r_{0})+\phi^{\prime}(r_{0})\Delta r$, where $\Delta r=r-r_{0}=h$, then we get:
$z_{tot}=(1+\alpha)z+2\alpha,$ (78)
where we have introduced the following notation:
$\alpha\equiv\frac{\phi^{\prime}(r_{0})}{2\phi(r_{0})}\,h.$ (79)
Let us evaluate (79) for different choices of the gauge scalar $\phi=\phi(r)$. Take, for example, the case (56) investigated in bambi-2017-a : $\phi(r)=\sqrt{6}M_{\text{pl}}/\sqrt{1+m^{4}/r^{4}}$, where we assumed the constant $L=m\equiv GM_{\oplus}/c^{2}$. In this case,
$\alpha=\frac{2m^{4}}{R^{4}_{\oplus}}\left(\frac{h}{R_{\oplus}}\right)=1.66\times 10^{-42},$ (80)
where we used the following values of the parameters: the gravitational constant $G=6.67\times 10^{-11}\,\text{m}^{3}/\text{s}^{2}\text{kg}$, the mass of Earth $M_{\oplus}=5.97\times 10^{24}\,\text{kg}$, mean radius of Earth $R_{\oplus}=6.37\times 10^{6}\,\text{m}$, the speed of light in vacuum $c=2.99\times 10^{8}\,\text{m}/\text{s}$, and the distance between the source and the absorber established in the original experiment by Pound and Rebka pound-1960 : $h=22.55\,\text{m}$. The value of $\alpha\sim 10^{-42}$ in (80) is far beyond the experimental capabilities on Earth. For comparison, the value of the redshift measured in pound-1960 was $z=2.5\times 10^{-15}$.
Another illustrative example can be given by the arbitrary gauge choice $\phi(r)=\sqrt{6}M_{\text{pl}}(1-2m/r)^{-1/4}$. In this case,
$\alpha=\frac{m\,h}{2R^{2}_{\oplus}}\approx 1.23\times 10^{-15},$
which is of the order of the redshift measured in the Pound-Rebka experiment. Since no such additional contribution has been obtained in pound-1960 , this means that this gauge choice is experimentally ruled out.
IX Discussion and concluding remarks
Here we have demonstrated that the very well-known result within the HEP community that local-scale invariance in flat spacetime requires the vanishing of the trace of the matter stress-energy tensor is not universal in curved spacetime. The requirement $T^{(m)}=0$ holds wherein the masses of timelike fields are constant parameters that are not affected by the Weyl recalings. In the case of perfect fluids, this result means that, under WTs, the energy density of the fluid (the same as that of its pressure) transforms as $\rho\rightarrow\Omega^{-3}\rho$. Therefore, the Lagrangian density of matter fields and/or of perfect fluids is not form-invariant under the conformal transformations. In contrast, if the masses of timelike fields are point-dependent quantities transforming as $m\rightarrow\Omega^{-1}m$, under the WTs, and the energy density of perfect fluids transforms as $\rho\rightarrow\Omega^{-4}\rho$, the Lagrangian density of matter is indeed form-invariant under Weyl transformations. In this case, Weyl symmetry does not require vanishing trace of the matter SET. Actually, when the matter field has a nonvanishing mass, which is transformed under (11) as $m\rightarrow\Omega^{-1}m$, the Lagrangian density of the matter fields is form-invariant under WTs (48), so that the following Ward identity takes place:
$g^{\mu\nu}\frac{\delta{\cal L}_{\text{tot}}}{\delta g^{\mu\nu}}=-\frac{\phi}{2}\frac{\delta{\cal L}_{\text{tot}}}{\delta\phi}\;\Rightarrow\;\frac{\delta{\cal L}_{\text{tot}}}{\delta\phi}=\frac{\sqrt{-g}}{\phi}\,T^{(m)}.$
This variational derivative provides the term with the nonvanishing trace $T^{(m)}$ in the $\phi$-EOM (36):
$\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}+\frac{\delta{\cal L}_{m}}{\delta\phi}=-\sqrt{-g}\left[\nabla^{2}\phi-\frac{\phi}{6}R-\frac{1}{\phi}T^{(m)}\right]=0.$
Therefore, $\phi$-EOM (36), coincides with the trace of the Einstein-type EOM. That is, KG-type EOM (36) is not an independent equation, so it does not contribute to the dynamics of the CGR theory. Our result that under the WTs (48), vanishing of the trace of the stress-energy tensor of matter is not required for Weyl invariance to be a symmetry of theory (37), may be of importance for a revised perspective of the Weyl anomaly capper-duff-1974 ; duff-1977 ; rusos-1984 ; affleck-1986 ; duff-1994 ; odintsov-2000 ; cai-2010 . Unfortunately, the investigation of this issue goes clearly beyond the aims of the present document.
The results discussed in this document are valid only for the Weyl-invariant CGR theory (37). Consider, for example, conformal gravity cgrav-1989 ; cgrav-2012 . The EOMs of this theoretical framework are derived from the following gravitational action:
$S_{\text{grav}}^{W}=\alpha\int d^{4}x\sqrt{-g}\,C^{2},$ (81)
where $\alpha$ is a dimensionless constant and $C^{2}\equiv C_{\mu\lambda\nu\sigma}C^{\mu\lambda\nu\sigma}$ ($C_{\mu\lambda\nu\sigma}$ is the Weyl conformal tensor). The EOMs read (we include the coupling of matter fields);
$W^{(2)}_{\mu\nu}-\frac{1}{3}\,W^{(1)}_{\mu\nu}=\frac{1}{4\alpha}\,T^{(m)}_{\mu\nu}$, (82)
where the following tensors have been defined:
$W^{(1)}_{\mu\nu}$= $2g_{\mu\nu}\nabla^{2}R-2\nabla_{\mu}\nabla_{\nu}R-2RR_{\mu\nu}+\frac{1}{2}\,g_{\mu\nu}R^{2},$ (83)
$W^{(2)}_{\mu\nu}$= $\frac{1}{2}\,g_{\mu\nu}\nabla^{2}R+\nabla^{2}R_{\mu\nu}+\frac{1}{2}\,g_{\mu\nu}R_{\lambda\sigma}R^{\lambda\sigma}-2R_{\mu\lambda}R^{\lambda}_{\;\;\nu}-\nabla^{\lambda}\nabla_{\mu}R_{\lambda\nu}-\nabla^{\lambda}\nabla_{\nu}R_{\lambda\mu}.$ (84)
In this case, there is no scalar field $\phi$ involved, so our demonstration in Section III is not applicable. Besides, note the distinguishing property of this theory: If we take the trace of (82), we get $T^{(m)}=0$. Hence, only the matter with traceless SET can be coupled to conformal gravity (81). This time, the traceless condition is derived directly from the EOM independent of the variational procedure. The demonstration of the traceless condition is straightforward: the trace $W^{(1)}=6\nabla^{2}R$, while $W^{(2)}=3\nabla^{2}R-2\nabla^{\sigma}\nabla^{\lambda}R_{\sigma\lambda}=2\nabla^{2}R$, where we used the Bianchi identity $\nabla^{\lambda}G_{\lambda\mu}=0$. The trace of the EOM (82) is given by
$T^{(m)}=4\alpha[W^{(2)}-\frac{1}{3}\,W^{(1)}]=0.$
Nevertheless, a Weyl-invariant gravitational action leading to a theory where GR is recovered at a certain limit or in a certain gauge may be of the form:
$S_{\text{grav}}=\frac{1}{2}\int d^{4}x\sqrt{-g}\left[\frac{1}{6}\,\phi^{2}R+(\partial\phi)^{2}+2C^{2}\right],$ (85)
which is called massive conformal gravity faria-b ; faria-c ; faria-d ; faria-e ; faria-f . In this case, the demonstration in Section III is applicable again, so the traceless condition does not arise.
Interestingly, the above Weyl invariant framework may provide a suitable classical description of the gravitational interactions of matter, which meets Dicke’s principle that the (classical) laws of physics must be invariant under the transformations of units; the name used by Dicke for the Weyl transformations (28). Frequent arguments against the Weyl invariance of physical laws must be adequately identified as arguments against local scale invariance. The latter must not be confused with form-invariance under Weyl transformations:
$g_{\mu\nu}\rightarrow\Omega^{2}g_{\mu\nu},\;\phi\rightarrow\Omega^{-1}\phi,\;\chi_{i}\rightarrow\Omega^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow\Omega^{-1}m.$
Acknowledgments The author thanks Felipe F Faria for pointing out several bibliographic references to us and for the useful exchange of ideas and critical comments.
Funding: This work has been supported by FORDECYT-PRONACES-CONACYT under grant CF-MG-2558591.
Declaration of generative AI and AI-assisted technologies in the manuscript preparation process.
During the preparation of this work, the author used the WRITEFULL tool on overleaf in order to check spelling and grammar. After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the published article.
Appendix A Lagrangian density of matter in terms of Weyl invariant quantities
Here, for illustration, we demonstrate equality (45) for perfect fluids and also for the Dirac fermion. Let us start with the Lagrangian density of a perfect fluid. In this case, the demonstration of equality (45) is a very simple exercise. Let us define the Weyl invariant energy density of the fluid:
$\mathfrak{r}:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-4}\rho.$ (86)
Hence, trivially,
${\cal L}_{\text{fluid}}:=-\sqrt{-g}\,\rho\equiv-\sqrt{-\mathfrak{g}}\,\mathfrak{r},$ (87)
where the left-hand Lagrangian density is written in terms of the fields that suffer the WTs, while the right-hand Lagrangian is written in terms of Weyl invariant quantities.
A.1 The Dirac Lagrangian density
The Lagrangian density of a Dirac fermion $\psi$ ($\bar{\psi}$ is the adjoint spinor of the fermion) in curved space is given by:
[…]
where $\cancel{\cal D}:=\gamma^{\mu}{\cal D}_{\mu}$ and the following covariant derivative has been defined:
${\cal D}_{\mu}\psi=\left(D_{\mu}-\frac{1}{2}\sigma_{ab}e^{b\nu}\nabla_{\mu}e^{a}_{\nu}\right)\psi.$ (89)
In these equations $a,b,c=0,1,2,3$ are the flat spacetime indices, while $\gamma^{a}$ are the Dirac gamma matrices, $e^{a}_{\mu}$ are the tetrad fields such that $g_{\mu\nu}=\eta_{ab}e^{a}_{\mu}e^{b}_{\nu}$ ($\eta_{ab}$ is the Minkowski metric). Besides $\gamma^{\mu}=e^{\mu}_{c}\gamma^{c}$, and
$\sigma_{ab}=\frac{1}{2}\left[\gamma_{a},\gamma_{b}\right]=\frac{1}{4}\left(\gamma_{a}\gamma_{b}-\gamma_{b}\gamma_{a}\right),$
are the generators of the Lorentz group in the spin representation. Above, we have used the standard definition of the $SU(2)\times U(1)$ gauge derivative, $D_{\mu}\psi$ (see, for example, Equation (32) of quiros_prd_2025 ).
Let us define the following Weyl invariant quantities:
$\mathfrak{e}^{a}_{\mu}:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)e^{a}_{\mu},\;\mathfrak{g}_{\mu\nu}:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{2}g_{\mu\nu},\;\Psi:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-\frac{3}{2}}\psi,\;\mathfrak{m}:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-1}m.$ (90)
We can also compute the Levi-Cività (LC) connection (Christoffel symbols) of the Weyl invariant metric $\mathfrak{g}_{\mu\nu}$:
$\mathfrak{C}^{\alpha}_{\mu\nu}=\frac{1}{2}\mathfrak{g}^{\alpha\lambda}\left(\mathfrak{g}_{\mu\lambda,\nu}+\mathfrak{g}_{\nu\lambda,\mu}-\mathfrak{g}_{\mu\nu,\lambda}\right),$ (91)
and the related covariant derivative operator $\mathfrak{V}_{\mu}$. We have
$e^{b\nu}\nabla_{\mu}e^{a}_{\nu}=\mathfrak{e}^{b\nu}\mathfrak{V}_{\mu}\mathfrak{e}^{a}_{\nu}+\left(\mathfrak{e}^{a}_{\mu}\mathfrak{e}^{b\lambda}-\mathfrak{e}^{b}_{\mu}\mathfrak{e}^{a\lambda}\right)\frac{\partial_{\lambda}\phi}{\phi},$
so
$\sigma_{ab}e^{b\nu}\nabla_{\mu}e^{a}_{\nu}=\sigma_{ab}\left(\mathfrak{e}^{b\nu}\mathfrak{V}_{\mu}\mathfrak{e}^{a}_{\nu}+2\mathfrak{e}^{a}_{\mu}\mathfrak{e}^{b\lambda}\frac{\partial_{\lambda}\phi}{\phi}\right).$
Let us next define the Weyl invariant gamma matrix in a curved background;
$\Gamma^{\mu}:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-1}\gamma^{\mu}=\mathfrak{e}^{\mu}_{a}\gamma^{a}=\mathfrak{e}^{\mu a}\gamma_{a}.$ (92)
It is straightforward to find
$\cancel{D}\psi=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{\frac{5}{2}}\left[\Gamma^{\mu}\left(D_{\mu}-\frac{1}{2}\sigma_{ab}\mathfrak{e}^{b\nu}\mathfrak{V}_{\mu}\mathfrak{e}^{a}_{\nu}\right)\Psi+\left(\frac{3}{2}\mathfrak{e}^{b\lambda}\gamma_{b}-\sigma_{ab}\gamma^{a}\mathfrak{e}^{b\lambda}\right)\frac{\partial_{\lambda}\phi}{\phi}\Psi\right],$
that is,
$\cancel{D}\psi=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{\frac{5}{2}}\cancel{\mathfrak{D}}\Psi,$ (93)
where $\cancel{\mathfrak{D}}\Psi=\Gamma^{\mu}\left(D_{\mu}-\frac{1}{2}\sigma_{ab}\mathfrak{e}^{b\nu}\mathfrak{V}_{\mu}\mathfrak{e}^{a}_{\nu}\right)\Psi$ is the Weyl invariant gauge covariant slash derivative and we have used the identity $\sigma_{ab}\gamma^{a}=3\gamma_{b}/2$. Hence,
$\sqrt{-g}\,\bar{\psi}\cancel{D}\psi=\sqrt{-\mathfrak{g}}\,\bar{\Psi}\cancel{\mathfrak{D}}\Psi.$
The right-hand side (RHS) of this equality is written in terms of Weyl invariant quantities. This serves as an additional demonstration that the left-hand side (LHS) is form-invariant under WTs. It is not difficult to demonstrate that,
$\sqrt{-g}\,m\bar{\psi}\psi=\sqrt{-\mathfrak{g}}\,\mathfrak{m}\bar{\Psi}\Psi.$
Hence, we have demonstrated that the Lagrangian density of the Dirac spinor fulfills the following equality:
[…]
where the LHS is written in terms of the fields that suffer the Weyl transformations, while the RHS is written in terms of Weyl invariant quantities.

### v4

@@
-Weyl symmetry without the traceless condition
+Could an overlooked symmetry of the classical laws of gravity explain the cosmological puzzles?
 Abstract
-We show that the requirement that the trace of the stress-energy tensor of matter must vanish if invariance under Weyl transformations is a symmetry of a given gravitational theory is not universal.
-This requirement holds whenever the masses of timelike fields are constant parameters that are not transformed by conformal transformations, or when the energy density of perfect fluids transforms as $\rho\rightarrow\Omega^{-3}\rho$.
-In contrast, if the masses of timelike fields are point-dependent quantities transforming under conformal transformations as $m\rightarrow\Omega^{-1}m$, and the energy density of perfect fluids transforms as $\rho\rightarrow\Omega^{-4}\rho$, the Weyl symmetry does not require the vanishing of the trace of the matter SET.
-In consequence, any matter fields, regardless of whether the trace of their stress-energy tensor vanishes or not, can be coupled to gravity.
-The phenomenological and physical consequences of the novel result are drawn.
+We show that if the masses of timelike fields are point-dependent quantities transforming under conformal transformations as $m\rightarrow\Omega^{-1}m$, and the energy density of perfect fluids transforms as $\rho\rightarrow\Omega^{-4}\rho$, form-invariance under Weyl transformations is an actual symmetry of the gravitational interactions of matter.
+That is, under the mentioned circumstances, Weyl symmetry allows any matter field to be coupled to gravity.
+The phenomenological and physical consequences of the novel result, including the “many worlds” interpretation of gauge freedom, are drawn.
+We explore, in particular, a possible explanation of two major cosmological puzzles: dark matter and dark energy, as a consequence of Weyl symmetry.
+Quantum-mechanical removal of the spacetime singularities in this framework is briefly discussed as well.
 I Introduction
@@
 where $\Omega=\Omega(x)$ is the conformal factor ($\Omega(x)>0$), $\Phi_{i}$ are other gravitational and matter fields present, and $w_{i}$ are the conformal weights of these fields.
-Here we call this “Dicke’s principle.”
-The high-energy physics (HEP) argument that, due to the Planck mass $M_{\text{pl}}$, a dimensionful parameter, gravity cannot be conformally invariant, seems to oppose Dicke’s principle dicke-1962 .
+Here we call this invariance “Dicke’s principle.”
+The high-energy physics (HEP) argument that, due to the Planck mass $M_{\text{pl}}$, being a dimensionful parameter, gravity cannot be conformally invariant, seems to oppose Dicke’s principle dicke-1962 .
 In general, the presence of any mass term is believed to break the conformal invariance deser-1970 ; callan_prd_1970 .
-However, strictly speaking, this is true only if we consider constant masses that do not transform under (1): $m\rightarrow m$.
+However, strictly speaking, this is true only if we consider that constant masses do not transform under (1): $m\rightarrow m$.
 As stated in bekenstein_1980 , we must differentiate between LSTs:
@@
 $g_{\mu\nu}\rightarrow\Omega^{2}g_{\mu\nu},\;\Phi_{i}\rightarrow\Omega^{w_{i}}\Phi_{i},\;m\rightarrow\Omega^{-1}m,$ (3)
-where, in addition to CTs (1), the following transformation property of the masses is assumed:
-$m\rightarrow\hat{m}=\Omega^{-1}m.$ (4)
+where, in addition to CTs (1), the transformation of the masses $m\rightarrow\Omega^{-1}m$, is assumed.
 That is, the mass parameters are fields with conformal weight $w_{m}=-1$ hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 .
-According to bekenstein_1980 , scale transformations (2) are not units’ transformations; they are active enlargements of a system whose usefulness depends on the absence of a length scale.
-Meanwhile, WTs are units’ transformations (3) whose physical and geometrical meanings follow from the existence of some scale of length.
-In the present document, as is customary when working within the framework of scalar-tensor theories (STTs) dicke-1962 ; bekenstein_1980 ; faraoni_rev ; fujii_book ; faraoni_book ; faraoni_prd_2007 ; quiros_ijmpd_2019 ; hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 , we consider WTs (3), instead of LSTs (2).
-In consequence, it is assumed that the masses are point-dependent quantities that, under (1), transform like (4).
-Whether the mass parameter transforms or not under the Weyl transformations plays a central role in the demonstration of conformal form-invariance of the action of fundamental matter fields quiros_prd_2025 :
-$S_{m}=\int d^{4}x\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu}),$ (5)
+According to bekenstein_1980 , scale transformations (2) are not unit transformations; they are active enlargements of a system whose usefulness depends on the absence of a length scale.
+Meanwhile, WTs (3) are units’ transformations whose physical and geometrical meanings follow from the existence of some scale of length.
+In the present document, as is customary within the framework of scalar-tensor theories (STTs) dicke-1962 ; bekenstein_1980 ; faraoni_rev ; fujii_book ; faraoni_book ; faraoni_prd_2007 ; quiros_ijmpd_2019 ; hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 , we consider WTs (3), instead of LSTs (2).
+In consequence, it is assumed that the masses are point-dependent fields that, under (1), transform like in (3).
+Whether the mass parameter transforms or not under CTs plays a central role in the demonstration of conformal form-invariance of the action of fundamental matter fields quiros_prd_2025 :
+$S_{m}=\int d^{4}x\,{\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})=\int d^{4}x\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu}),$ (4)
 where $L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$ is the Lagrangian of the fundamental matter fields $\chi_{i}$, with i=1,2,3,,Ni=1,2,3,...,N, and $N$ – the number of matter fields that are minimally coupled to the metric $g_{\mu\nu}$.
@@
 where ${\cal L}_{m}=\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$ is the Lagrangian density of matter.
-For fermions and gauge bosons, it has been shown in quiros_prd_2025 that if their masses are point-dependent quantities transforming as (4), the matter action (5) is form-invariant under WTs
-$g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu},\;\chi_{i}\rightarrow\hat{\chi}_{i}=\Omega^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow\hat{m}=\Omega^{-1}m,$ (7)
-where $w_{\chi_{i}}$ is the conformal weight of the matter field $\chi_{i}$.
+For fermions and gauge bosons, it has been shown in quiros_prd_2025 that if their masses are point-dependent quantities transforming as in (3), the matter action (4) is form-invariant under WTs
+$g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu},\;\chi_{i}\rightarrow\hat{\chi}_{i}=\Omega^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow\hat{m}=\Omega^{-1}m.$ (6)
 But if the masses are constant parameters, then (4) is not form-invariant under (2).
 Take, for illustration, the Lagrangian density of the Dirac fermion: [^1] Here we use the following signature of the metric: (,+,+,+)(-,+,+,+).
-We follow the notation of quiros_prd_2025 (see also the Appendix A).
+We follow the notation of quiros_prd_2025 .
 […]
@@
 A large body of applications of gravitational theories is associated with background matter in the form of a perfect fluid, rather than with fundamental fields.
-[^3] For a perfect fluid with energy density $\rho$ and barotropic pressure $p$, the SET is given by $T^{(m)}_{\mu\nu}=(\rho+p)u_{\mu}u_{\nu}+pg_{\mu\nu}$, where $u^{\mu}=dx^{\mu}/d\tau$ is the four-velocity of the fluid ($d\tau=ids$ $\Rightarrow d\tau^{2}=-ds^{2}$).
-Therefore, the trace of the SET of the perfect fluid reads: $T^{(m)}=-\rho+3p$.
-It vanishes only for radiation with the equation of state (EOS) $p=\rho/3$.
-Therefore, it is of interest to look for the transformation properties of the action of perfect fluids under Weyl transformations (3).
+Therefore, it is of interest to investigate the transformation properties of the action of perfect fluids under Weyl transformations (3).
 In quiros_prd_2025 , it has also been shown that, in addition to the action of fundamental matter fields, that of perfect fluids is form-invariant under WTs as well.
@@
 That is, the action of the perfect fluid is form-invariant under WTs.
-Note, however, that if we assume that the mass unit is not transformed, then, under CTs $\rho\rightarrow\hat{\rho}=\Omega^{-3}\rho$, which means that the action of the perfect fluid is not form-invariant: $\int d^{4}x\sqrt{-g}\,\rho=\int d^{4}x\sqrt{-\hat{g}}\,\Omega\hat{\rho}$.
-The requirement of vanishing of the matter SET trace, in connection with LSS in curved spacetime, was derived in deser-1970 , in the case of a scalar field non-minimally coupled to gravity.
+Note, however, that if we assume that the mass unit is not transformed under CTs, then $\rho\rightarrow\hat{\rho}=\Omega^{-3}\rho$, which means that the action of the perfect fluid is not form-invariant: $\int d^{4}x\sqrt{-g}\,\rho=\int d^{4}x\,\Omega\sqrt{-\hat{g}}\hat{\rho}$.
+The requirement of vanishing of the matter SET trace, in connection with local scale symmetry (LSS) in curved spacetime, was derived in deser-1970 , in the case of a scalar field non-minimally coupled to gravity.
 [^3] The vanishing of the matter SET trace as a requirement for conformal symmetry to hold ccj_ann_phys_1970 was primarily investigated within the HEP framework, where gravity is ignored, so the background spacetime is flat space with the Minkowski metric.
@@
 where $R$ is the curvature scalar, $\phi$ is the non-minimally coupled scalar field with mass dimensions, and we use the following shorthand notation: $(\partial\phi)^{2}\equiv g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi$.
-In deser-1970 , it is demonstrated that the form-invariance of (9) under LSTs (2), is possible thanks to the vanishing trace of the matter SET: $T^{(m)}\equiv g^{\mu\nu}T^{(m)}_{\mu\nu}=0$.
-In this document, we shall demonstrate that the traceless condition $T^{(m)}=0$ is not a universal requirement for Weyl symmetry to materialize.
-The result of reference deser-1970 , that only matter fields with traceless SET can be coupled to gravity in CGR theory (9), is right only if we assume that the masses of timelike fields are constant parameters that are not transformed by CTs (1), so the energy density of a perfect fluid transforms as: $\rho\rightarrow\hat{\rho}=\Omega^{-3}\rho$; that is, if we consider local scale transformations (2) instead of Weyl transformations (3).
+In deser-1970 , it is demonstrated that the form-invariance of (8) plus a matter action piece $S_{m}=\int d^{4}x\sqrt{-g}\,L_{m}$ under LSTs (2), is possible only if the trace of the matter SET vanishes: $T^{(m)}\equiv g^{\mu\nu}T^{(m)}_{\mu\nu}=0$.
+Here, we shall demonstrate that the traceless condition $T^{(m)}=0$ is not a universal requirement for Weyl symmetry to materialize.
+The result of reference deser-1970 , that only matter fields with traceless SET can be coupled to gravity in CGR theory (8), is correct only if we assume that the masses of timelike fields are constant parameters that are not transformed by CTs (1), so the energy density of a perfect fluid transforms as: $\rho\rightarrow\hat{\rho}=\Omega^{-3}\rho$; that is, if we consider local scale transformations (2) instead of Weyl transformations (3).
 Under this assumption, the Lagrangian density of the matter fields breaks LSS.
-In contrast, if we assume that the masses of timelike fields are point-dependent quantities that, under the CTs (1), transform as $m\rightarrow\hat{m}=\Omega^{-1}m$, and consequently the energy density of the perfect fluid transforms as $\rho\rightarrow\hat{\rho}=\Omega^{-4}\rho$, the Lagrangian density of the fundamental matter fields, as well as the Lagrangian density of perfect fluids, are form-invariant under the WTs (3):
-$\int d^{4}x\sqrt{-g}\,L_{m}(\chi_{i},g_{\mu\nu})=\int d^{4}x\sqrt{-\hat{g}}\,\Omega^{-4}\,L_{m}(\Omega^{-w_{\chi_{i}}}\hat{\chi}_{i},\Omega^{-2}\hat{g}_{\mu\nu})=\int d^{4}x\sqrt{-\hat{g}}\,L_{m}(\hat{\chi}_{i},\hat{g}_{\mu\nu}),$
-that is,
-$\int d^{4}x\sqrt{-g}\,L_{m}(\chi_{i},g_{\mu\nu})=\int d^{4}x\sqrt{-\hat{g}}\,L_{m}(\hat{\chi}_{i},\hat{g}_{\mu\nu}).$ (10)
-Similarly, the action of a perfect fluid is form-invariant under (7) as well: $\int d^{4}x\sqrt{-g}\,\rho=\int d^{4}x\sqrt{-\hat{g}}\,\hat{\rho}$.
-In these cases, since the Lagrangian density of matter is form-invariant under WTs, the Ward identity associated with Weyl invariance leads to the following nonvanishing variational derivative: $\delta{\cal L}_{m}/\delta\phi=\sqrt{-g}\,T^{(m)}/\phi$, which provides the term with the trace of the stress-energy tensor of matter in the EOM of the scalar field $\phi$.
-In contrast, if the Lagrangian density of the matter fields is not form-invariant under CTs (1), the Ward identity does not hold and $\delta{\cal L}_{m}/\delta\phi=0$.
-Our main hypothesis in this work is that Dicke’s principle holds in nature.
+In contrast, if we assume that the masses of timelike fields are point-dependent fields that, under the CTs (1), transform as $m\rightarrow\hat{m}=\Omega^{-1}m$, and consequently the energy density of the perfect fluid transforms as $\rho\rightarrow\hat{\rho}=\Omega^{-4}\rho$, the Lagrangian density of the fundamental matter fields, as well as the Lagrangian density of perfect fluids, are form-invariant under the WTs (3):
+$\int d^{4}x\sqrt{-g}\,L_{m}(\chi_{i},g_{\mu\nu})=\int d^{4}x\sqrt{-\hat{g}}\,L_{m}(\hat{\chi}_{i},\hat{g}_{\mu\nu}),$ (9)
+and $\int d^{4}x\sqrt{-g}\,\rho=\int d^{4}x\sqrt{-\hat{g}}\,\hat{\rho}$, respectively.
+Our primary hypothesis in this work is that Dicke’s principle holds in nature.
 That is, the laws of gravity are form-invariant under Weyl transformations (3).
-In consequence, we consider that the masses of fundamental timelike matter fields $m_{i}$ are point-dependent quantities bekenstein_1980 ; hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 : $m_{i}(x)=\kappa_{i}\,\phi(x)$, where $\kappa_{i}$ are dimensionless constants.
-[^5] This choice is dictated by the fact that, since the scalar field $\phi$ has mass dimensions, under WTs the mass parameter transforms as $\phi$ does; that is, $m\propto\phi$.
-These transform as (4) under Weyl transformations, while the energy density of a perfect fluid transforms as $\rho\rightarrow\Omega^{-4}\rho$.
+In consequence, we consider that the masses of the fundamental timelike matter fields $m_{i}$ are point-dependent fields bekenstein_1980 ; hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 : $m_{i}(x)=\kappa_{i}\,\phi(x)$, where $\kappa_{i}$ are dimensionless constants.
+[^4] This choice is dictated by the fact that, since the scalar field $\phi$ has mass dimensions, under CTs the mass parameter transforms as $\phi$ does; that is, $m\propto\phi$.
+These transform as in (3) under Weyl transformations, while the energy density of a perfect fluid transforms as $\rho\rightarrow\Omega^{-4}\rho$.
+In these cases, since the Lagrangian density of matter is form-invariant under WTs, the Ward identity associated with Weyl invariance leads to the following nonvanishing variational derivative: [^5] In contrast, if the mass parameter $m$ is an untransformed constant, the Lagrangian density of the matter fields is not form-invariant under CTs (1), the Ward identity does not hold and $\delta{\cal L}_{m}/\delta\phi=0$.
+$\delta{\cal L}_{m}/\delta\phi=\sqrt{-g}\,T^{(m)}/\phi$, which provides the term with the trace of the stress-energy tensor of matter in the EOM of the scalar field $\phi$.
+We will aim to investigate the phenomenological and physical consequences of the resulting Weyl-invariant theory of gravity, that is, of CGR.
+In particular, the possibility of finding an alternative explanation for two major cosmological puzzles: dark matter (DM) and dark energy (DE), will be explored.
 We have organized the present document as follows.
@@
 This is required for further comparison with the results of Section III, where the equations of motion (EOM) are derived from the overall CGR action under the assumptions of the present work, above all form-invariance of the matter action under WTs (3).
-In Section IV, we derive some physical consequences of the CGR theory, in particular, the fifth force required by form-invariance under WTs, which does not interact with radiation, thus acting as a ”dark force.” In this section, we also discuss the two possible approaches to CTs in the configuration space: i) passive and ii) active approaches.
+In this Section, we also derive some physical consequences of the CGR theory, in particular, the fifth force required by form-invariance under WTs, which does not interact with radiation, thus acting as a ”dark force” with the potential to explain the “dark sector” of present cosmology.
+In Section IV, we discuss the two possible approaches to CTs in the configuration space quiros_prd_2025 : i) passive and ii) active approaches.
 We shall show that only the active approach is suitable for the investigation of the physical and phenomenological consequences of the Weyl symmetry.
@@
 We draw the similitude of the picture arising from gauge freedom within the CGR theory and the “many-worlds” interpretation of quantum mechanics everett_rmp_1957 ; wheeler ; dewitt_phys_rev_1967 ; dewitt ; barvinsky_cqg_1990 ; omnes_rmp_1992 ; tegmark_1998 ; garriga_prd_2001 ; zurek_rmp_2003 ; tegmark_nature_2007 ; quiros_prd_2023 .
-In Sections VI and VII, we expose some important cosmological consequences of the present setup, including the possible explanation of the galactic rotation curves in some gauge of the CGR theory, as well as some encouraging aspects of Weyl invariant cosmology.
-Possible observational consequences of CGR theory, mainly related to measurements of the redshift of frequencies, are analyzed in Section VIII.
-The novel results of the present work are discussed in Section IX, accompanied by brief concluding remarks.
+In Sections VI and VII, we expose some important cosmological consequences of the present setup, including the possible explanation of the galactic rotation curves in some gauge of the CGR theory, as well as some encouraging aspects of Weyl-invariant cosmology.
+Possible observational consequences of CGR theory related to frequency redshift measurements are analyzed in Section VIII, where a natural explanation of the measured redshifts of supernovae of type Ia (SNIa) is provided without requiring dark energy.
+A discussion of quantum-mechanical removal of spacetime singularities in the present many-worlds framework is given in Section IX.
+This is followed by brief concluding remarks of the present work in Section X.
+An appendix A is included where the curvature redshift is derived as a solution of the null-geodesic equation.
 II Local scale symmetry and the trace of the stress-energy tensor of matter
@@
 where $G_{\mu\nu}\equiv R_{\mu\nu}-g_{\mu\nu}R/2$ is the Einstein tensor and
-${\cal E}_{\mu\nu}\equiv G_{\mu\nu}-\frac{6}{\phi^{2}}\,\Theta^{(\phi)}_{\mu\nu},\;\Theta^{(\phi)}_{\mu\nu}\equiv\frac{1}{6}\left(\nabla_{\mu}\nabla_{\nu}-g_{\mu\nu}\nabla^{2}\right)\phi^{2}-T^{(\phi)}_{\mu\nu}$.
-(14)
+${\cal E}_{\mu\nu}\equiv G_{\mu\nu}-\frac{6}{\phi^{2}}\,\Theta^{(\phi)}_{\mu\nu},\;\Theta^{(\phi)}_{\mu\nu}\equiv\frac{1}{6}\left(\nabla_{\mu}\nabla_{\nu}-g_{\mu\nu}\nabla^{2}\right)\phi^{2}-T^{(\phi)}_{\mu\nu}$, (13)
+where we use the shorthand notation $\nabla^{2}\equiv g^{\mu\nu}\nabla_{\mu}\nabla_{\nu}$.
 The modified SET of the scalar field $\Theta^{(\phi)}_{\mu\nu}$ ccj_ann_phys_1970 , has finite elements to all orders in the perturbation parameter and defines the same four-momentum and Lorentz generators as the conventional tensor: [^6] Note that in the expression for $\Theta^{(\phi)}_{\mu\nu}$ the standard stress-energy tensor of the scalar field $T^{(\phi)}_{\mu\nu}$, has the wrong sign.
-This is because, in the gravitational Lagrangian density (12), the kinetic energy density of the scalar field appears with the wrong sign, which is a requirement for the CGR gravitational Lagrangian density (12) to be invariant under WTs (11).
+This is because, in the gravitational Lagrangian density (11), the kinetic energy density of the scalar field appears with the wrong sign, which is a requirement for the CGR gravitational Lagrangian density (11) to be invariant under CTs (10).
 This has no physical consequences, since $\phi$ is a gauge field that does not obey a specific EOM.
 $T^{(\phi)}_{\mu\nu}\equiv\partial_{\mu}\phi\partial_{\nu}\phi-\frac{1}{2}g_{\mu\nu}(\partial\phi)^{2}.$ (14)
-Here we use the shorthand notation $\nabla^{2}\equiv g^{\mu\nu}\nabla_{\mu}\nabla_{\nu}$. If we take into account the trace of the tensor ${\cal E}_{\mu\nu}$:
+If we take into account the trace of the tensor ${\cal E}_{\mu\nu}$:
 ${\cal E}\equiv g^{\mu\nu}{\cal E}_{\mu\nu}=-R+\frac{6}{\phi}\,\nabla^{2}\phi,$ (15)
@@
 $g^{\mu\nu}\frac{\delta{\cal L}_{\text{grav}}}{\delta g^{\mu\nu}}=\sqrt{-g}\frac{\phi^{2}}{12}\,{\cal E}=-\sqrt{-g}\frac{\phi}{2}\left(\frac{1}{6}\,\phi R-\nabla^{2}\phi\right),\;\frac{\phi}{2}\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}=\sqrt{-g}\frac{\phi}{2}\left(\frac{1}{6}\,\phi R-\nabla^{2}\phi\right),$
-respectively. If we compare these equations, we obtain the following Ward identity:
+respectively. Therefore, the following Ward identity is obtained:
 $g^{\mu\nu}\frac{\delta{\cal L}_{\text{grav}}}{\delta g^{\mu\nu}}=-\frac{\phi}{2}\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}.$ (16)
-This identity is a consequence of the invariance of the CGR gravitational action (9) under CT (11). Actually, if this action is form-invariant under infinitesimal CTs:
+This identity is a consequence of the invariance of the CGR gravitational action (8) under CTs (10). Actually, if this action is form-invariant under infinitesimal CTs:
 $\delta g_{\mu\nu}=2\theta\,g_{\mu\nu}\,\left(\delta g^{\mu\nu}=-2\theta\,g^{\mu\nu},\;\delta\sqrt{-g}=4\theta\sqrt{-g}\right),\;\delta\phi=-\theta\,\phi,$ (17)
 where for convenience (and only temporarily) we rescaled the conformal factor: $\Omega=e^{\theta}$, then $S_{\text{grav}}\rightarrow S_{\text{grav}}+\delta S_{\text{grav}}$, where $\delta S_{\text{grav}}=0$; so that, under (17), $S_{\text{grav}}\rightarrow S_{\text{grav}}$.
-Hence;
+We have;
 $\delta S_{\text{grav}}=\frac{1}{2}\int d^{4}x\left(\frac{\delta{\cal L}_{\text{grav}}}{\delta g^{\mu\nu}}\delta g^{\mu\nu}+\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}\delta\phi\right)=-\theta\int d^{4}x\left(g^{\mu\nu}\frac{\delta{\cal L}_{\text{grav}}}{\delta g^{\mu\nu}}+\frac{\phi}{2}\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}\right)=0,$
-where in the last equality, we substituted $\delta g^{\mu\nu}$ and $\delta\phi$ from infinitesimal CTs (18). Specifically, from the last equation, the Ward identity (17) follows.
-After applying the variational principle of least action, from (13) we obtain the vacuum Einstein EOM and the Klein-Gordon-type EOM:
-$\frac{\delta{\cal L}_{\text{grav}}}{\delta g^{\mu\nu}}$ $=0\;\Rightarrow\;{\cal E}_{\mu\nu}=0\;\Rightarrow\;G_{\mu\nu}=\frac{6}{\phi^{2}}\,\Theta^{(\phi)}_{\mu\nu}$, (19)
-$\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}$ $=0\;\Rightarrow\;\left(\nabla^{2}-\frac{1}{6}\,R\right)\phi=0,$ (20)
-respectively.
-The trace of the Einstein EOM (19) reads $R+6\nabla^{2}\phi/\phi=0$, which, after some simple algebra, can be written as the KG-EOM (20).
-Hence, the latter is not an independent equation.
-Consequently, the scalar field $\phi$ does not obey an independent EOM.
-This is consistent with the new degree of freedom (DOF) associated with the Weyl symmetry of (9) since, in addition to the four DOFs for making spacetime diffeomorphisms, there is a DOF for making CTs (11), which is exhausted by the gauge field $\phi$.
-Let us next consider matter fields coupled to CGR;
+where in the last equality, we substituted $\delta g^{\mu\nu}$ and $\delta\phi$ from infinitesimal CTs (17). The Ward identity (16) follows from the last equation.
+Let us consider matter fields coupled to CGR;
 $S_{\text{tot}}=\frac{1}{2}\int d^{4}x\sqrt{-g}\left[\frac{1}{6}\,\phi^{2}R+(\partial\phi)^{2}\right]+\int d^{4}x\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu}),$ (18)
-where, since it is assumed in deser-1970 that the CTs do not transform the masses, the matter Lagrangian of time-like matter fields breaks LSS.
+where, in deser-1970 , local scale transformations (2) are considered rather than Weyl transformations (3).
+Hence, it is assumed in the mentioned bibliographic reference that the CTs do not transform the masses; that is, the matter Lagrangian of time-like matter fields breaks LSS.
 In this case, the Ward identity for matter fields is not satisfied, so $\delta{\cal L}_{m}/\delta\phi=0$, because the matter Lagrangian density does not depend explicitly on $\phi$.
@@
 $\frac{\delta{\cal L}_{\text{tot}}}{\delta g^{\mu\nu}}=\sqrt{-g}\,\frac{\phi^{2}}{12}\,\left[{\cal E}_{\mu\nu}-\frac{6}{\phi^{2}}\,T^{(m)}_{\mu\nu}\right],\;\frac{\delta{\cal L}_{\text{tot}}}{\delta\phi}=\sqrt{-g}\left(\frac{1}{6}\,\phi R-\nabla^{2}\phi\right),$ (19)
-respectively, where we have omitted divergence terms and assumed that,
+respectively, where we have omitted divergence terms and took into account that,
 $\frac{\delta{\cal L}_{m}}{\delta\phi}=0\;\Rightarrow\;\frac{\delta{\cal L}_{\text{tot}}}{\delta\phi}=\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}.$ (20)
-If we apply the variational principle, from (22) we get the following EOMs:
+From (19) we get the following EOMs:
 $\frac{\delta{\cal L}_{\text{tot}}}{\delta g^{\mu\nu}}$ $=0\;\Rightarrow\;{\cal E}_{\mu\nu}=\frac{6}{\phi^{2}}\,T^{(m)}_{\mu\nu}\;\Rightarrow\;G_{\mu\nu}=\frac{6}{\phi^{2}}\left[\Theta^{(\phi)}_{\mu\nu}+T^{(m)}_{\mu\nu}\right$], (21)
 $\frac{\delta{\cal L}_{\text{tot}}}{\delta\phi}$ $=0\;\Rightarrow\;\left(\nabla^{2}-\frac{1}{6}\,R\right)\phi=0.$ (22)
-Note that the Einstein EOM (19) has been modified by adding the SET of matter, whereas the KG-type EOM (20) is not modified since the Lagrangian density of the matter fields ${\cal L}_{m}=\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$ does not depend explicitly on $\phi$, that is, equation (23) takes place.
-This results in the following requirement.
-If we compare the trace of (24):
+If we compare the trace of the Einstein EOM (21):
 $-R+\frac{6}{\phi}\nabla^{2}\phi=\frac{6}{\phi^{2}}T^{(m)}\;\Rightarrow\;\left(\nabla^{2}-\frac{1}{6}\,R\right)\phi=\frac{1}{\phi}\,T^{(m)},$ (23)
-with the EOM (25), we obtain the “traceless condition:” $T^{(m)}=0$.
-Therefore, only matter fields or fluids with vanishing SET trace: photons and radiation, in general, can be coupled to gravity in this local scale-invariant theory.
+with the Klein-Gordon (KG) type EOM (22), we obtain the “traceless condition:” $T^{(m)}=0$.
+Therefore, only matter fields and fluids with vanishing SET trace: photons and radiation, in general, can be coupled to gravity in this local scale-invariant theory.
 [^7] Non-relativistic matter fields that behave like radiation can also be coupled to CGR faria-a .
@@
 Here, we associate Weyl symmetry with units’ transformations (3), which are popular in the study of scalar-tensor gravitational theories fujii_book ; faraoni_book ; flanagan_cqg_2004 ; faraoni_prd_2007 ; catena_prd_2007 ; chiba_2013 ; capozziello_prd_2013 ; jarv_2015 ; quiros_ijmpd_2019 .
-III Weyl symmetry does not require vanishing trace of the stress-energy tensor of matter
-In this section, we will demonstrate that, if we assume that the masses of timelike fields are themselves point-dependent fields that transform as $m\rightarrow\hat{m}=\Omega^{-1}m$, while the energy density of a perfect fluid transforms as $\rho\rightarrow\hat{\rho}=\Omega^{-4}\rho$, it is not necessary to make any changes in (21) to get a consistent coupling of any matter fields.
-That is, under these assumptions, the coupling of the matter fields in (21) preserves the Weyl symmetry regardless of whether $T^{(m)}$ vanishes or not.
-Our demonstration is based on the Weyl invariance of matter action in this case, as demonstrated in quiros_prd_2025 for fundamental fields and perfect fluids.
-Under the above assumptions, for fermions and gauge bosons as well as for perfect fluids, the matter action (5) is not only invariant but also form-invariant under WTs.
+III Conformal General Relativity with arbitrary matter fields
+The CGR theory deser-1970 whose Lagrangian density is given by the following expression (see action (18)):
+${\cal L}_{\text{tot}}=\frac{\sqrt{-g}}{2}\left[\frac{1}{6}\,\phi^{2}R+(\partial\phi)^{2}\right]+{\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu}),$ (24)
+is form-invariant under the following Weyl transformations:
+$g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu},\;\phi\rightarrow\hat{\phi}=\Omega^{-1}\phi,\;\chi_{i}\rightarrow\hat{\chi}_{i}=\Omega^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow\hat{m}=\Omega^{-1}m.$ (25)
+In this section, we will demonstrate that, if we assume that the masses of timelike fields are themselves point-dependent fields that transform as $m\rightarrow\hat{m}=\Omega^{-1}m$, while the energy density of a perfect fluid transforms as $\rho\rightarrow\hat{\rho}=\Omega^{-4}\rho$, it is not necessary to make any changes in (18)/(24) to get a consistent coupling of any matter fields $\xi_{i}$ to Weyl invariant gravity.
+That is, under these assumptions, the coupling of the matter fields in (24) preserves Weyl symmetry, regardless of whether $T^{(m)}$ vanishes.
+Our demonstration is based on the Weyl invariance of matter action, as demonstrated in quiros_prd_2025 for fundamental fields and perfect fluids.
+Under the above assumptions, for fermions, gauge bosons, and perfect fluids, the matter action (4) is not only invariant but also form-invariant under WTs.
 That is, under (25), the following equality takes place:
-$\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})=\sqrt{-\hat{g}}\,L_{m}(\hat{\chi}_{i},\hat{\partial}\hat{\chi}_{i},\hat{g}_{\mu\nu}).$ (27)
+${\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})={\cal L}_{m}(\hat{\chi}_{i},\hat{\partial}\hat{\chi}_{i},\hat{g}_{\mu\nu})\;\Leftrightarrow\;\sqrt{-g}\,L_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})=\sqrt{-\hat{g}}\,L_{m}(\hat{\chi}_{i},\hat{\partial}\hat{\chi}_{i},\hat{g}_{\mu\nu}).$ (26)
 From this equation it follows that if $g_{\mu\nu}$ is the physical metric with respect to the matter fields $\chi_{i}$; that is, the matter fields $\chi_{i}$ are minimally coupled to $g_{\mu\nu}$, the conformal metric $\hat{g}_{\mu\nu}$ is the physical metric with respect to the conformal matter fields $\hat{\chi}_{i}$ which are minimally coupled to $\hat{g}_{\mu\nu}$.
 In other words, if the metric $g_{\mu\nu}$ is the physical metric on some gauge, its conformal metric $\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu}$ is the physical metric on the conformal gauge.
-Equation (27) plays an important role because it shows that the total action (21) is Weyl invariant regardless of the type of matter coupled to gravity: fermions, gauge bosons, or perfect fluids.
-The transformation of masses $m\rightarrow\Omega^{-1}m$ and the consequent transformation of the energy density (and pressure) of a perfect fluid: $\rho\rightarrow\Omega^{-4}\rho$, is fundamental for the demonstration of form-invariance of the Lagrangian density of timelike fundamental matter fields quiros_prd_2025 .
-In this case, the Lagrangian density of matter fields is form-invariant under WTs, so the variational derivative: $\delta{\cal L}_{m}/\delta\phi$ obeys the Ward identity; that is, it is nonvanishing.
-This is the primary assumption underlying our approach to the Weyl symmetry of the CGR theory, which distinguishes our result from the well-known result that the matter SET trace must vanish if local scale invariance is a symmetry of the CGR theory.
-Let us apply the variational procedure under the assumption that the action of the matter degrees of freedom coupled to gravity: fermions, gauge bosons, and perfect fluids, is form-invariant under the following WTs:
-$g_{\mu\nu}\rightarrow\Omega^{2}g_{\mu\nu},\;\phi\rightarrow\Omega^{-1}\phi,\;\chi_{i}\rightarrow\Omega^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow\Omega^{-1}m.$ (28)
-Let us consider the Lagrangian density of matter fields: ${\cal L}_{m}={\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$.
-Since it is invariant under WTs (28), then, under infinitesimal Weyl transformations (18) plus $\delta\chi_{i}=w_{\chi_{i}}\theta\chi_{i}$ and $\delta m=-\theta m$, we have $\delta{\cal L}_{m}=\nabla_{\mu}V^{\mu}$, where $V^{\mu}$ is some vector or $\delta{\cal L}_{m}=0$.
+Equation (26) plays an important role because it shows that the total Lagrangian density (24) is Weyl-invariant regardless of the type of matter coupled to gravity: fermions, gauge bosons, or perfect fluids.
+Given that the Lagrangian density of matter fields ${\cal L}_{m}$ is form-invariant under WTs (25), the variational derivative: $\delta{\cal L}_{m}/\delta\phi$ obeys the Ward identity; that is, it is nonvanishing.
+This is the primary assumption underlying our approach to the Weyl symmetry of the CGR theory, which distinguishes our result from the well-known result deser-1970 that the matter SET trace must vanish if local scale invariance is a symmetry of the CGR theory.
+Let us apply the variational procedure under the assumption that the action of the matter degrees of freedom coupled to gravity: fermions, gauge bosons, and perfect fluids, is form-invariant under WTs (25).
+We consider the Lagrangian density of matter fields: ${\cal L}_{m}={\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$.
+Since it is invariant under WTs (25), then, under infinitesimal Weyl transformations (17) plus $\delta\chi_{i}=w_{\chi_{i}}\theta\chi_{i}$ and $\delta m=-\theta m$, we have $\delta{\cal L}_{m}=0$, or $\delta{\cal L}_{m}=\nabla_{\mu}V^{\mu}$, where $V^{\mu}$ is some vector.
 In any case, $\delta S_{m}=\int d^{4}x\,\delta{\cal L}_{m}=0$, so that under (17), $S_{m}\rightarrow S_{m}+\delta S_{m}=S_{m}$; that is, the matter action is invariant under infinitesimal Weyl transformations
@@
 $\frac{\delta{\cal L}_{m}}{\delta\chi_{i}}=0\;\Rightarrow\;\frac{\partial{\cal L}_{m}}{\partial\chi_{i}}-\nabla_{\mu}\left[\frac{\partial{\cal L}_{m}}{\partial(\partial_{\mu}\chi_{i})}\right]=0,$ (29)
-are satisfied. [^8] For classical gravitational problems, it is assumed that the matter fields are a priori given. From the last equality in (30), it follows the Ward identity:
+are satisfied. [^8] For most of classical gravitational problems, it is assumed that the matter fields are a priori given. From the last equality in (28), it follows the Ward identity:
 $g^{\mu\nu}\frac{\delta{\cal L}_{m}}{\delta g^{\mu\nu}}=-\frac{\phi}{2}\frac{\delta{\cal L}_{m}}{\delta\phi}.$ (30)
@@
 $\frac{\delta{\cal L}_{m}}{\delta\phi}=\frac{\sqrt{-g}}{\phi}\,T^{(m)}.$ (31)
-We have
+Let us now apply the variational principle to derive the equations of motion. We have
 $\delta S_{\text{tot}}=\delta S_{\text{grav}}+\delta S_{m}=\int d^{4}x$ $\left[\frac{\delta{\cal L}_{\text{grav}}}{\delta g^{\mu\nu}}\delta g^{\mu\nu}+\frac{\delta{\cal L}_{\text{grav}}}{\delta\phi}\delta\phi+\frac{\delta{\cal L}_{m}}{\delta g^{\mu\nu}}\delta g^{\mu\nu}+\frac{\delta{\cal L}_{m}}{\delta\phi}\delta\phi\right]=0,$
@@
 where in the last line of (32), in the factor multiplying $\delta\phi$ (expression within square brackets), we have taken into account the modified Ward identity (31).
-The variational principle leads to the following EOM:
+The variational principle leads to the following EOMs:
 ${\cal E}_{\mu\nu}=\frac{6}{\phi^{2}}T^{(m)}_{\mu\nu}$, (33)
 $\nabla^{2}\phi-\frac{\phi}{6}R=\frac{1}{\phi}T^{(m)}.$ (34)
-This is an outstanding result, since the trace of the matter SET is not required to vanish.
+This is an outstanding result because, while the EOM (33), (34), are form-invariant under Weyl transformations, the trace of the matter SET is not required to vanish.
 Note that the KG-EOM (34) is not an independent equation, as it coincides with the trace of Einstein-type EOM (33).
-That is, the scalar field $\phi$ is not a dynamical field since it does not obey an independent EOM.
-IV CGR: Physics beyond GR
-The CGR theory deser-1970 whose Lagrangian density is given by the following expression (see action (21)):
-${\cal L}_{\text{tot}}=\frac{\sqrt{-g}}{2}\left[\frac{1}{6}\,\phi^{2}R+(\partial\phi)^{2}\right]+{\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu}),$ (37)
-is form-invariant under the WTs (28). The independent EOM derived from (37), is the following Einstein-type equation
-${\cal E}_{\mu\nu}=\frac{6}{\phi^{2}}\,T^{(m)}_{\mu\nu}\;\Rightarrow\;G_{\mu\nu}$ $=\frac{6}{\phi^{2}}\left[\Theta^{(\phi)}_{\mu\nu}+T^{(m)}_{\mu\nu}\right$], (38)
+That is, the scalar field $\phi$ is not a dynamical field, as it does not satisfy an independent EOM.
+III.1 The “dark force”
+The independent EOM derived from (24), is the Einstein-type equation (33);
+${\cal E}_{\mu\nu}=\frac{6}{\phi^{2}}\,T^{(m)}_{\mu\nu}$, (35)
 where the tensor $\Theta^{(\phi)}_{\mu\nu}$ is defined in (13).
 There is no independent EOM that the gauge field $\phi$ must obey, since the KG-type equation (34), which is derived from (24) by varying with respect to $\phi$, coincides with the trace of (35).
-Another independent equation is provided by the continuity equation.
+The continuity equation provides another independent equation.
 If we take the divergence of (35), and use the second Bianchi identity $\nabla^{\lambda}G_{\lambda\mu}=0$, and the identity $(\nabla_{\lambda}\nabla_{\mu}-\nabla_{\mu}\nabla_{\lambda})\nabla^{\lambda}\phi^{2}=R_{\mu\lambda}\nabla^{\lambda}\phi^{2}$, we obtain
 $\nabla^{\lambda}(\phi^{2}{\cal E}_{\lambda\mu})=6\partial_{\mu}\phi\left(\nabla^{2}-\frac{1}{6}\,R\right)\phi=6\nabla^{\lambda}T^{(m)}_{\lambda\mu}$.
-Then, if we substitute the trace of the Einstein EOM (36) in the latter equation, we get the following inhomogeneous continuity equation:
+Then, if we substitute (34) in the latter equation, we get the following inhomogeneous continuity equation:
 $\nabla^{\lambda}T^{(m)}_{\lambda\mu}=\frac{\partial_{\mu}\phi}{\phi}\,T^{(m)}.$ (36)
 The inhomogeneous term on the right-hand side (RHS) of (36) is required for the above equation to be form-invariant under WTs (25).
-This term represents a fifth force, which can be related to the point-dependent property of the mass parameter in the present theory $m(x)=\kappa\,\phi(x)$ bekenstein_1980 ; hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 .
-Since timelike fields acquire masses through Yukawa interactions with the Higgs field $H$ after $SU(2)\times U(1)$ symmetry breaking, it is required to modify the standard model of particles (SMP) to get point-dependent masses.
-That is, the SMP must be Weyl invariant, and the $SU(2)\times U(1)$ symmetry breaking must not spoil Weyl symmetry.
-Therefore, the Higgs Lagrangian density must be appropriately modified bars-2014 ; mohamedi-2024 .
-In particular, the mass parameter $v$ in the Higgs potential, $V\propto(H^{\dagger}H-v^{2})^{2}$, must be replaced by a field: $v\rightarrow v_{0}\phi$, where $v_{0}$ is some dimensionless parameter bars-2014 .
-After the mentioned modification, the timelike fields acquire point-dependent masses through the Yukawa interaction.
-For example, for a fermion field, the Yukawa coupling is given by $\kappa_{f}\bar{\psi}\phi\psi=\bar{\psi}\,m_{f}\,\psi$ $\Rightarrow$ $m_{f}=\kappa_{f}\,\phi$.
-Since $\phi$ is the gravitational coupling in CGR theory, that is, $\phi\sim M_{\text{pl}}$, a mass scale issue could arise.
-However, we will never be able to measure the point-dependent mass, $m(x)=\kappa\,\phi(x)$, in local experiments, because measuring sticks and clocks are made of atoms (matter) that suffer the same universal point-dependence, so the mass unit also varies: $u_{m}(x)=u_{0}\,\phi(x)$.
-In experiments, we measure the ratio faraoni_prd_2007
-$\frac{m}{u_{m}}=\frac{\kappa}{u_{0}},$
-that does not depend on the spacetime point.
-Hence, even if we assume $\phi_{0}\sim M_{\text{pl}}$, the value $\phi_{0}$ is not what is measured in experiments.
-In other words, the gigantic mass $m=\kappa\,\phi_{0}$ is compensated for by an equally gigantic mass unit $u_{m}=u_{0}\,\phi_{0}$, so there is no mass scale problem associated with point-dependent masses.
-The point-dependence of masses (and the related fifth force) can be observed only in “nonlocal” experiments; that is, in experiments that imply emission and absorption of photons coming from distant spacetime points; for example, a photon emitted by some atom located at some point $(t,\vec{x})$ is absorbed by an identical atom at the origin $(0,\vec{0})$.
-Since the mass of electrons is point-dependent, so is the energy of the emitted/absorbed photon.
-On the other hand, photon paths are not sensitive to WTs (the Riemannian null-geodesics are already Weyl invariant).
-This mismatch makes it possible to reveal the point-dependence of the masses, but adequate experiments must be designed (see Section VIII for further discussion of this subject).
-IV.1 The fifth force
-An interesting property of the CGR theory is related to the inhomogeneous continuity equation (39): $\nabla^{\lambda}T^{(m)}_{\lambda\mu}=T^{(m)}\partial_{\mu}\phi/\phi$.
-While this equation is satisfied by matter fluxes of stress and energy, the equivalent EOM for point-like particles is the following inhomogeneous EOM for time-like fields:
+This term represents a fifth force, which can be related to the point-dependent property of the mass parameter in the present theory: $m(x)=\kappa\,\phi(x)$ bekenstein_1980 ; hobson-prd-2020 ; hobson-epjc-2022 ; casas_1992 .
+While (36) is satisfied by matter fluxes of stress and energy, the equivalent EOM for point-like particles is the following EOM for time-like fields:
 $\frac{d^{2}x^{\alpha}}{ds^{2}}+\left\{{}^{\alpha}_{\mu\nu}\right\}\frac{dx^{\mu}}{ds}\frac{dx^{\nu}}{ds}=\frac{\partial_{\mu}\phi}{\phi}\,h^{\mu\alpha},$ (37)
@@
 For that reason, we call it “dark force.” This opens up the possibility that this extra force could explain, in principle, the dark energy and dark matter components of the cosmic fluid.
-IV.2 Passive conformal transformations
-It has been demonstrated in woodard-1986 (see also hobson-epjc-2022 ) that there is no physics beyond GR in CGR theory.
-The argument goes like this.
-To assess the physical consequences of CGR (in woodard-1986 called “conformal scalar-metric theory” in contrast to GR, called “purely metric theory”), we must rely on Weyl-invariant objects and operators.
-In GR theory, every physical question can be reduced to the VEV of some coordinate-invariant functional of the metric ${\cal F}[g]$.
-Similarly, every physical question in CGR can be reduced to the VEV of a coordinate and conformally invariant functional of $\phi$ and $g_{\mu\nu}$: $\hat{{\cal F}}[\phi,g]$.
-In woodard-1986 , it is demonstrated that the VEV of ${\cal F}[g]$ for the theory whose action is $S[g]$, equals the VEV of $\hat{{\cal F}}[\phi,g]$ for the theory whose action is $\hat{S}[\phi,g]$.
-Hence, any conformal scalar-metric theory is gauge equivalent to the corresponding pure metric theory.
-Let us demonstrate the above result in an alternative way and show that the related point of view represents only one of two possible approaches to CTs (28): i) passive Weyl transformations (PWTs), and ii) active Weyl transformations (AWT’s) quiros_prd_2025 .
-Consider some abstract space of fields ${\cal M}_{f}$ where the metric $g_{\mu\nu}$, the gauge field $\phi$, the masses $m$, and any matter fields $\chi_{i}$ coupled to gravity, are “generalized coordinates.” Each point $P:(\phi,g_{\mu\nu},\chi_{i},m)$ in ${\cal M}_{f}$ represents a gravitational state.
-Passive CTs
+The assumption that $m=m(\phi)$ can have consequences for the standard model of particles (SMP).
+Actually, since timelike fields acquire masses through Yukawa interactions with the Higgs field $H$ after $SU(2)\times U(1)$ symmetry breaking, it is required to modify the SMP to get point-dependent masses.
+That is, the SMP must be Weyl-invariant, and the $SU(2)\times U(1)$ symmetry breaking must not spoil Weyl symmetry.
+Therefore, the Higgs Lagrangian density must be appropriately modified bars-2014 ; mohamedi-2024 .
+In particular, the mass parameter $v$ in the Higgs potential, $V\propto(H^{\dagger}H-v^{2})^{2}$, must be replaced by a field: $v\rightarrow v_{0}\phi$, where $v_{0}$ is some dimensionless parameter bars-2014 .
+After the mentioned modification, the timelike fields acquire point-dependent masses through the Yukawa interaction.
+For example, for a fermion field, the Yukawa coupling is given by $\kappa_{f}\bar{\psi}\phi\psi=\bar{\psi}\,m_{f}\,\psi$ $\Rightarrow$ $m_{f}=\kappa_{f}\,\phi$.
+Notice that, as long as $\phi$ is the gravitational coupling in CGR theory, that is, $\phi\sim M_{\text{pl}}$, a mass scale issue could arise.
+However, we will never be able to measure the $\phi$-dependent mass, $m(x)=\kappa\,\phi(x)$, in local experiments, because measuring sticks and clocks are made of atoms (matter) that share the same universal $\phi$-dependence, so the mass unit also varies: $u_{m}(x)=u_{0}\,\phi(x)$.
+In experiments, we measure the ratio faraoni_prd_2007
+$\frac{m}{u_{m}}=\frac{\kappa}{u_{0}},$
+that does not depend on the spacetime point.
+Hence, even if we assume $\phi_{0}\sim M_{\text{pl}}$, the value $\phi_{0}$ is not what is measured in experiments.
+In other words, the gigantic mass $m=\kappa\,\phi_{0}$ is compensated for by an equally gigantic mass unit $u_{m}=u_{0}\,\phi_{0}$, so there is no mass scale problem associated with point-dependent masses.
+The point-dependence of masses (and the related fifth force) can be observed only in “nonlocal” experiments.
+For example, in experiments that imply emission and absorption of photons coming from distant spacetime points, where a photon emitted by some atom located at some point $(t,\vec{x})$ is absorbed by an identical atom at the origin $(0,\vec{0})$.
+Since the mass of electrons is point-dependent, so is the energy of the emitted/absorbed photon.
+On the other hand, photon paths are not sensitive to WTs (the Riemannian null-geodesics are already Weyl-invariant).
+This mismatch makes it possible to reveal the point-dependence of the masses, but adequate experiments must be designed (see Section VIII for further discussion of this subject).
+IV CGR: Physics beyond GR
+It has been demonstrated in hobson-epjc-2022 ; woodard-1986 that there is no physics beyond GR in CGR theory.
+[^10] In woodard-1986 CGR is named “conformal scalar-metric theory” in contrast to GR, named “purely metric theory.” While in woodard-1986 , it is shown that any “conformal scalar-metric theory” is gauge equivalent to the corresponding “pure metric theory”, in hobson-epjc-2022 , it is argued that if one works in terms of quantities that can be physically measured (in particular, the metric to which matter fields are minimally coupled), conformal Schwarzschild-de Sitter (SdS) metric does not lead to any different result than SdS metric; that is, “in any conformal frame, the trajectories followed by ordinary matter particles are merely the timelike geodesics of the SdS metric.”
+Here we shall show that the above-cited result, although correct, is not universal.
+It represents only one of two possible approaches to WTs (25): i) passive Weyl transformations (PWTs), and ii) active Weyl transformations (AWT’s) quiros_prd_2025 .
+Let us consider some abstract space of fields ${\cal M}_{f}$ where the metric $g_{\mu\nu}$, the gauge field $\phi$, the masses $m$, and any matter fields $\chi_{i}$ coupled to gravity, are “generalized coordinates.” Each point $P:(\phi,g_{\mu\nu},\chi_{i},m)$ in ${\cal M}_{f}$ represents a “global gravitational state” (GGS).
+[^11] Here, by a global gravitational state we understand full knowledge of the fields at every spacetime point: $\phi=\phi(t,\vec{x})$, $g_{\mu\nu}=g_{\mu\nu}(t,\vec{x})$, $\chi_{i}=\chi_{i}(t,\vec{x})$ and $m=m(t,\vec{x})$.
+Passive Weyl transformations
 $g_{\mu\nu}\rightarrow g^{\prime}_{\mu\nu}=\Omega^{2}g_{\mu\nu},\;\phi\rightarrow\phi^{\prime}=\Omega^{-1}\phi,\;\chi_{i}\rightarrow\chi^{\prime}_{i}=\Omega^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow m^{\prime}=\Omega^{-1}m,$ (40)
 may be viewed as a “rotation” of the coordinate system $R(g_{\mu\nu},\phi,\chi_{i},m)$ $\rightarrow$ $R^{\prime}(g^{\prime}_{\mu\nu},\phi^{\prime},\chi^{\prime}_{i},m^{\prime})$ (see Figure 1 of quiros_prd_2025 where, for simplicity, the vacuum case is considered).
-The physically meaningful quantities are the invariants:
-$\mathfrak{g}_{\mu\nu}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{2}\,g_{\mu\nu},\;\Psi_{i}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{w_{\chi_{i}}}\chi_{i},\;\mathfrak{m}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-1}m.$ (44)
-Given that these are the physical (thus measurable) quantities according to PWTs, the fields $g_{\mu\nu}$, $\phi$, $\chi_{i}$, and $m$, which are transformed according to (43) we shall call “auxiliary fields.” The gravitational state is given by ${\cal S}:(\mathfrak{g}_{\mu\nu},\Psi_{i},\mathfrak{m})$.
-Hence, $R$ and $R^{\prime}$ are two different “coordinate representations” of the same point $P:(\mathfrak{g}_{\mu\nu},\Psi_{i},\mathfrak{m})\in{\cal M}_{f}$; that is, of the same gravitational state $\cal S$}.
+The physically meaningful quantities are the invariant metric $\mathfrak{g}_{\mu\nu}$, matter fields $\Psi_{i}$, and mass $\mathfrak{m}$:
+$\mathfrak{g}_{\mu\nu}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{2}\,g_{\mu\nu},\;\Psi_{i}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{w_{\chi_{i}}}\chi_{i},\;\mathfrak{m}=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-1}m,$ (41)
+respectively.
+We have that $\mathfrak{g}_{\mu\nu}$, $\Psi_{i}$, and $\mathfrak{m}$ are the physical (thus measurable) quantities according to PWTs, while the fields $g_{\mu\nu}$, $\phi$, $\chi_{i}$, and $m$, which are transformed according to (40) are just “auxiliary fields.” The global gravitational state is given by ${\cal S}:(\mathfrak{g}_{\mu\nu},\Psi_{i},\mathfrak{m})$.
+Hence, $R$ and $R^{\prime}$ are two different “coordinate representations” of the same point $P:(\mathfrak{g}_{\mu\nu},\Psi_{i},\mathfrak{m})\in{\cal M}_{f}$; that is, of the same GGS $\cal S$}.
 It has been demonstrated in quiros_prd_2025 that the Lagrangian density of matter fields (this includes both fundamental matter fields and perfect fluids) is form-invariant under WTs (40).
@@
 This equality relates the Lagrangian density of the auxiliary matter fields $\chi_{i}$ to that of the physical matter fields $\Psi_{i}$.
-In the Appendix A, for illustrative purposes, we show equality (45) for the Lagrangian density of the Dirac fermion and for that of perfect fluids.
-For other fundamental fields, e.g., gauge bosons, the demonstration is similar.
+In quiros-arxiv-2025 , equality (42) is shown for the Lagrangian density of the Dirac fermion and for that of perfect fluids.
+[^12] For perfect fluids, the demonstration of (42) is quite simple.
+The Lagrangian density for a perfect fluid with energy density $\rho$ can be written as ${\cal L}_{\text{fluid}}[\rho,g]=-\sqrt{-g}\,\rho$.
+If we define the Weyl invariant energy density $\mathfrak{r}:=(M_{\text{pl}}/\sqrt{6}\phi)^{4}\rho$, and take into account the definition of the Weyl invariant metric tensor (41), then, since $\sqrt{-\mathfrak{g}}=(\sqrt{6}\phi/M_{\text{pl}})^{4}\sqrt{-g}$, we get that $\sqrt{-\mathfrak{g}}\,\mathfrak{r}=\sqrt{g}\,\rho$; that is, ${\cal L}_{\text{fluid}}[\rho,g]={\cal L}_{\text{fluid}}[\mathfrak{r},\mathfrak{g}]$, which is (42) for the case of perfect fluids.
+For other fundamental fields, e.
+g., gauge bosons, the demonstration is similar.
 In view of equality (42), the equation (24) can be written as
@@
 ${\cal L}_{\text{tot}}=\sqrt{-\mathfrak{g}}\,\frac{M^{2}_{\text{pl}}}{2}\mathfrak{R}+{\cal L}_{m}(\Psi_{i},\mathfrak{D}\Psi_{i},\mathfrak{g}_{\mu\nu}),$ (44)
-where $\mathfrak{R}$ is the curvature scalar defined in terms of the Weyl invariant, thus physical, metric.
+where $\mathfrak{R}$ is the curvature scalar defined in terms of the Weyl-invariant, thus physical, metric.
 Written in this (covariant) way, the total Lagrangian density coincides with that of GR.
-The difference between (37) and (47) according to the PWTs is that the latter Lagrangian density is written in terms of physical quantities, while the former one is written in terms of auxiliary quantities without direct physical meaning.
+The difference between (24) and (44) according to the PWTs is that the latter Lagrangian density is written in terms of physical quantities, while the former one is written in terms of auxiliary fields without direct physical meaning.
 We notice that the auxiliary fields have been removed from the covariant formulation (44).
@@
 However, as we have already commented, this is only one of two possible understandings of WTs (25).
-IV.3 Active transformations
 […]
@@
 is already written covariantly in terms of physical fields $g_{\mu\nu}$, $\phi$, $\chi_{i}$, and $m$.
-In this case, Weyl invariants, such as those defined in (44), represent useful relationships between the physical fields in the two different gravitational states:
+In this case, Weyl invariants, such as those defined in (41), represent useful relationships between the physical fields in the two different global gravitational states:
 $\mathfrak{g}_{\mu\nu}$ $=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{2}\,g_{\mu\nu}=\left(\frac{\hat{\phi}}{\sqrt{6}M_{\text{pl}}}\right)^{2}\,\hat{g}_{\mu\nu}\;\Rightarrow\;g_{\mu\nu}=\left(\frac{\hat{\phi}}{\phi}\right)^{2}\,\hat{g}_{\mu\nu},$
@@
 For example, the GR gauge is achieved by choosing $\hat{\phi}=\sqrt{6}\,M_{\text{pl}}=$ const.
+[^13] We underline that once we make an specific gauge choice $\phi=\phi(x)$, the explicit Weyl invariance is lost.
+This is similar to the loss of explicit coordinate invariance after a specific choice of coordinates is made within the GR framework.
 Under this choice, the total Lagrangian density (24) is written as;
@@
 Here, we also consider the transformation $T^{(m)}_{\mu\nu}\rightarrow\hat{T}^{(m)}_{\mu\nu}=\Omega^{-2}\,T^{(m)}_{\mu\nu}$ of the matter SET; that is, $\hat{T}^{(m)}_{\mu\nu}=(\phi/\sqrt{6}M_{\text{pl}})^{-2}\,T^{(m)}_{\mu\nu}$.
-Since GR is a gauge of Weyl invariant theory (37), we refer to the latter theory as CGR.
+Since GR is a gauge of Weyl-invariant theory (24), we refer to the latter theory as CGR.
 Note that in an arbitrary gauge, where ${\cal L}_{m}={\cal L}_{m}(\chi_{i},\partial\chi_{i},g_{\mu\nu})$, the matter fields $\chi_{i}$ are minimally coupled to the metric $g_{\mu\nu}$.
-Therefore, it is the metric that defines the proper time intervals; that is, it is the physical metric in an arbitrary gauge.
+Therefore, it is the metric that defines the proper time intervals; that is, it is the physical metric in that gauge.
 In the same way, the matter fields $\hat{\chi}_{i}$ are minimally coupled to the metric $\hat{g}_{\mu\nu}$, so it is the physical metric in the GR gauge.
-The GR gauge $\mathfrak{G}_{\text{gr}}=\left\{\sqrt{6}M_{\text{pl}},\hat{g}_{\mu\nu}(x)\right\}$, where $\hat{g}_{\mu\nu}=\hat{g}_{\mu\nu}(x)$ is any solution of the GR-EOM (51), and any other arbitrary gauge $\mathfrak{G}=\{\phi(x),g_{\mu\nu}(x)\}$, where $g_{\mu\nu}=g_{\mu\nu}(x)$ is a solution of the Einstein-type EOM (38) once we have fixed the function $\phi=\phi(x)$, represent potential descriptions of the conformal invariant gravitational interactions of matter according to CGR theory.
-Both solutions are related by
+Let us take, for example, the GR gauge $\mathfrak{G}_{\text{gr}}=\left\{\sqrt{6}M_{\text{pl}},\hat{g}_{\mu\nu}(x),\hat{\chi}_{i},\hat{m}\right\}$, where $\hat{g}_{\mu\nu}=\hat{g}_{\mu\nu}(x)$ is any solution of the GR-EOM (48) and $\hat{m}$= const.
+is the GR constant mass parameter, and any other arbitrary gauge $\mathfrak{G}=\{\phi(x),g_{\mu\nu}(x),\chi_{i},m\}$, where $g_{\mu\nu}=g_{\mu\nu}(x)$ is a solution of the Einstein-type EOM (33) once we have fixed the function $\phi=\phi(x)$ (it is assumed that the matter fields are given).
+These gauges represent potential descriptions of the conformal invariant gravitational interactions of matter according to CGR theory, where the respective solutions are related by
 $g_{\mu\nu}=\left(\frac{\sqrt{6}M_{\text{pl}}}{\phi}\right)^{2}\hat{g}_{\mu\nu}.$ (49)
-Hence, we do not need to solve the Einstein-type EOM (38) after making a specific choice of $\phi=\phi(x)$.
+Hence, we do not need to solve the Einstein-type EOM (33) once we make a specific choice for $\phi=\phi(x)$.
 We only need to know some GR solution $\hat{g}_{\mu\nu}(x)$, then, if we fix the function $\phi=\phi(x)$, the corresponding conformal metric, according to (49), is necessarily a solution of (33).
-The usefulness of this result is illustrated, for example, in a series of papers bambi-2017-a ; bambi-2017-b ; bambi-2017-c ; bambi-2018-a ; bambi-2018-b , where conformal invariance is used in several contexts, for example, to look for the properties of some astrophysical black holes, to investigate the black hole evaporation, as well as the singularity issues.
+The usefulness of this result is illustrated, for example, in a series of papers bambi-2017-a ; bambi-2017-b ; bambi-2017-c ; bambi-2018-a ; bambi-2018-b , where Weyl invariance is used in several contexts, for example, to look for the properties of some astrophysical black holes, to investigate the black hole evaporation, as well as the singularity issues.
 Note that in these bibliographic references, each conformal solution is assumed to be physically meaningful.
@@
 Therefore, only the GR solution $\hat{g}_{\mu\nu}(x)$ matters.
-In this case, both the GR gauge $\mathfrak{G}_{\text{gr}}$ and any arbitrary gauge $\mathfrak{G}_{k}$;
-$\mathfrak{G}_{k}=\{\phi_{k}(x),g^{(k)}_{\mu\nu}(x)\},\;g^{(k)}_{\mu\nu}=\left(\frac{\sqrt{6}M_{\text{pl}}}{\phi_{k}}\right)^{2}\hat{g}_{\mu\nu},\;k=1,2,...,N,$
-in the conformal equivalence class $\mathfrak{C}_{\text{conf}}=\{\mathfrak{G}_{\text{gr}},\mathfrak{G}_{1},\mathfrak{G}_{2},...,\mathfrak{G}_{N}\}$ of (38), represent different gravitational states that account for different phenomenological consequences.
-The choice of one gauge or another can only be an experimental matter.
-This may be understood as a classical version of the many-worlds interpretation of quantum mechanics everett_rmp_1957 ; wheeler ; dewitt_phys_rev_1967 ; dewitt ; barvinsky_cqg_1990 ; omnes_rmp_1992 ; tegmark_1998 ; garriga_prd_2001 ; zurek_rmp_2003 ; tegmark_nature_2007 ; quiros_prd_2023 applied to the gravitational interactions of matter.
-V.1 Vacuum solution: Conformal Schwarzschild space
+In this case, the GR gauge $\mathfrak{G}_{\text{gr}}$ and any arbitrary gauge $\mathfrak{G}_{k}$;
+$\mathfrak{G}_{k}=\{\phi_{k}(x),g^{(k)}_{\mu\nu}(x),...\},\;g^{(k)}_{\mu\nu}=\left(\frac{\sqrt{6}M_{\text{pl}}}{\phi_{k}}\right)^{2}\hat{g}_{\mu\nu},\;k=1,2,...,N,$ (50)
+where $N\rightarrow\infty$ and the ellipsis within curled brackets stands for other fields coupled to gravity, belong in the conformal equivalence class
+$\mathfrak{C}_{\text{conf}}=\{\mathfrak{G}_{\text{gr}},\mathfrak{G}_{1},\mathfrak{G}_{2},...,\mathfrak{G}_{N}\},$ (51)
+of CGR theory given by the Lagrangian density (24) and the derived EOM (33).
+As a result, these gauges correspond to different global gravitational states that describe distinct phenomenological consequences.
+The choice between one gauge and another is ultimately an experimental matter.
+The arising picture may be understood as a classical version of the many-worlds interpretation of quantum mechanics everett_rmp_1957 ; wheeler ; dewitt_phys_rev_1967 ; dewitt ; barvinsky_cqg_1990 ; omnes_rmp_1992 ; tegmark_1998 ; garriga_prd_2001 ; zurek_rmp_2003 ; tegmark_nature_2007 ; quiros_prd_2023 applied to the gravitational interactions of matter.
+Besides, in the present case, the many worlds picture arises not in spacetime but in the configuration space.
+V.1 Cautionary note on gauge freedom
+The interpretation of gauge freedom that we have applied within the framework of CGR theory, as well as its physical meaning, has nothing in common with the usual interpretation of gauge freedom in standard gauge theories of fields.
+Consider, for example, the Lagrangian of a Maxwell field $A_{\mu}$ coupled to a Dirac fermion $\psi$:
+${\cal L}_{\text{em}}=-\frac{1}{4}\,F_{\mu\nu}F^{\mu\nu}+\bar{\psi}(i\cancel{D}-m)\psi,$ (52)
+where $F_{\mu\nu}\equiv\partial_{[\mu}A_{\nu]}/2$ is the Maxwell field strength, $\psi$ is the Dirac spinor of the fermion field ($\bar{\psi}$ is its adjoint), $\gamma^{\mu}$ are the Dirac gamma matrices, and $\cancel{D}\psi\equiv\gamma^{\mu}D_{\mu}\psi=\gamma^{\mu}(\partial_{\mu}\psi+iA_{\mu}\psi)$ is the gauge covariant derivative.
+It is assumed that the background space is Minkowski space.
+The Lagrangian (52) is form-invariant under gauge transformations:
+$A_{\mu}\rightarrow\hat{A}_{\mu}=A_{\mu}+\partial_{\mu}\lambda(x),\;\psi\rightarrow\hat{\psi}=e^{-i\lambda(x)}\psi,\;\bar{\psi}\rightarrow\hat{\bar{\psi}}=\bar{\psi}\,e^{i\lambda(x)},$
+where the free function $\lambda(x)$ is the gauge parameter.
+In this case, gauge invariance of (52) means that the derived EOM, as well as the measured quantities; that is, the physically meaningful quantities: the Lorentz scalar type $\bar{\psi}\psi$, vector type $\bar{\psi}\gamma^{\mu}\psi$, tensor type $\bar{\psi}\gamma^{\mu}\gamma^{\nu}\psi$ quantities, and the time and spatial components of the gauge invariant field strength $F_{\mu\nu}$; the electric $E_{i}=F_{0i}$ and magnetic $B_{i}=-\epsilon_{ijk}F^{jk}/2$ fields ($\epsilon_{ijk}$ is the Levi-Civita tensor), respectively, are not affected by the gauge transformations; that is, $\bar{\psi}\psi=\hat{\bar{\psi}}\hat{\psi}$, $\bar{\psi}\gamma^{\mu}\psi=\hat{\bar{\psi}}\gamma^{\mu}\hat{\psi}$, …, $E_{i}=\hat{E}_{i}$ and $B_{i}=\hat{B}_{i}$.
+This means that these transformations do not modify the state of the Maxwell field in interaction with the fermion.
+Note that, in the present case, the fields $A_{\mu}$, $\psi$, and $\bar{\psi}$, by themselves, do not represent measured quantities, so these are physically meaningless.
+Even if we try to associate different electromagnetic (EM) states with different sets of these fields: $\mathfrak{E}:(\psi,\bar{\psi},A_{\mu})$, $\hat{\mathfrak{E}}:(\hat{\psi},\hat{\bar{\psi}},\hat{A}_{\mu})$, etc., these states share the same measured quantities so that they are physically undistinguishable.
+To all practical purposes, $\mathfrak{E}$ and $\hat{\mathfrak{E}}$ represent the same physical EM state.
+This is in contrast with the active interpretation of Weyl transformations of the physically meaningful quantities; that is, the metric, the scalar field, and the other fields coupled to gravity: $g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu}$, $\phi\rightarrow\hat{\phi}=\Omega^{-2}\phi$, …, where $\{\phi(x),g_{\mu\nu}(x),$...\} and $\{\hat{\phi}(x),\hat{g}_{\mu\nu}(x),$...\} represent different global gravitational states, which are characterized by different measured quantities; that is, different curvature properties, proper time lapse, etc.
+In this document, we call gauge fixing and gauge freedom the specific choice of the free scalar $\phi(x)$, as explained above, because, historically, Weyl called conformal transformations (1) as “gauge transformations”.
+V.2 Vacuum solution: Conformal Schwarzschild space
 In a series of papers bambi-2017-a ; bambi-2017-b ; bambi-2017-c ; bambi-2018-a ; bambi-2018-b , conformal invariance is used to investigate the properties of astrophysical black holes and to examine black hole evaporation and singularity issues.
@@
 In these examples, the Schwarzschild vacuum solution $d\hat{s}^{2}$ given by (54) and its conformal singularity-free solutions; $\sqrt{6}M_{\text{pl}}/\phi(r)]^{2}\,d\hat{s}^{2}$, where $\phi(r)$ can be given by (56) or (57) (or possibly by many other $r$-functions), give suitable descriptions of vacuum CGR (55).
-The fact that gravitational states with a spacetime singularity are in the same equivalence class as singularity-free spacetimes is very interesting because this might shed some light on the singularity issue, as done, for example, in bambi-2017-a ; bambi-2017-b ; bambi-2017-c ; bambi-2018-a ; bambi-2018-b .
-If instead of the AWTs we adopted the passive approach to WTs, the results in these bibliographic references would not have any physical significance.
+The fact that global gravitational states with a spacetime singularity are in the same equivalence class as singularity-free spacetimes is very interesting because this might shed some light on the singularity issue, as done, for example, in bambi-2017-a ; bambi-2017-b ; bambi-2017-c ; bambi-2018-a ; bambi-2018-b .
+[^17] If instead of the AWTs we adopted the passive approach to WTs, the results in bambi-2017-a ; bambi-2017-b ; bambi-2017-c ; bambi-2018-a ; bambi-2018-b would not have any physical significance.
 In fact, according to passive WTs, the physically meaningful metric $\mathfrak{g}_{\mu\nu}=[\phi(r)/\sqrt{6}M_{\text{pl}}]^{2}\,g_{\mu\nu}=\hat{g}_{\mu\nu}$, coincides with the GR vacuum metric.
 That is, only the Schwarzschild black hole solution (53) matters.
-V.2 Cautionary note on gauge freedom
-The interpretation of gauge freedom (also called gauge fixing) we have applied within the framework of CGR theory, as well as its physical meaning, has nothing in common with the usual interpretation of gauge freedom in standard gauge theories of fields.
-Consider, for example, the Lagrangian of a Maxwell field $A_{\mu}$ coupled to a Dirac fermion $\psi$:
-${\cal L}_{\text{em}}=-\frac{1}{4}\,F_{\mu\nu}F^{\mu\nu}+\bar{\psi}(i\cancel{D}-m)\psi,$ (58)
-where $F_{\mu\nu}\equiv\partial_{[\mu}A_{\nu]}/2$ is the Maxwell field strength, $\psi$ is the Dirac spinor of the fermion field ($\bar{\psi}$ is its adjoint), $\gamma^{\mu}$ are the Dirac gamma matrices, and $\cancel{D}\psi\equiv\gamma^{\mu}D_{\mu}\psi=\gamma^{\mu}(\partial_{\mu}\psi+iA_{\mu}\psi)$ is the gauge covariant derivative.
-It is assumed that the background space is Minkowski space.
-The Lagrangian (58) is form-invariant under gauge transformations:
-$A_{\mu}\rightarrow\hat{A}_{\mu}=A_{\mu}+\partial_{\mu}\lambda(x),\;\psi\rightarrow\hat{\psi}=e^{-i\lambda(x)}\psi,\;\bar{\psi}\rightarrow\hat{\bar{\psi}}=\bar{\psi}\,e^{i\lambda(x)},$
-where the free function $\lambda(x)$ is the gauge parameter.
-In this case, gauge invariance of (58) means, in turn, that the measured quantities: the Lorentz scalar type $\bar{\psi}\psi$, vector type $\bar{\psi}\gamma^{\mu}\psi$, and tensor type $\bar{\psi}\gamma^{\mu}\gamma^{\nu}\psi$ quantities, as well as the time and spatial components of the gauge invariant field strength $F_{\mu\nu}$; the electric $E_{i}=F_{0i}$ and magnetic $B_{i}=-\epsilon_{ijk}F^{jk}/2$ fields ($\epsilon_{ijk}$ is the Levi-Civita tensor), respectively, are not affected by the gauge transformations, so $\bar{\psi}\psi=\hat{\bar{\psi}}\hat{\psi}$, $\bar{\psi}\gamma^{\mu}\psi=\hat{\bar{\psi}}\gamma^{\mu}\hat{\psi}$, …, $E_{i}=\hat{E}_{i}$ and $B_{i}=\hat{B}_{i}$.
-This means that these transformations do not modify the state of the Maxwell field in interaction with the fermion.
-This is in contrast with the interpretation of conformal transformations of the physically meaningful quantities; that is, the metric, the scalar field, and the other fields coupled to gravity, according to the active approach: $g_{\mu\nu}\rightarrow\hat{g}_{\mu\nu}=\Omega^{2}g_{\mu\nu}$, $\phi\rightarrow\hat{\phi}=\Omega^{-2}\phi$, …, where $\{\phi(x),g_{\mu\nu}(x),$...\} and $\{\hat{\phi}(x),\hat{g}_{\mu\nu}(x),$...\} represent different gravitational states, which are characterized by different curvature properties, proper time lapse, etc.
-In this document, we call gauge fixing and gauge freedom the specific choice of the free scalar $\phi(x)$, as explained above, because, historically, Weyl called conformal transformations (1) as “gauge transformations”.
+Yet, it is ultimately the experiment that decides which gauge is the one that better describes the gravitational phenomena.
+If it were, for example, the GR gauge, the one selected by experiment, then the spacetime singularity would still be there.
+Nevertheless, as we will show in Section IX, consideration of quantum-mechanical effects in the present framework may remove the classical spacetime singularity.
 VI Galactic rotation curves
-In harko-prd-2023 , the possibility that the observed behavior of test particles outside galaxies, which is usually explained with the help of dark matter, could be explained, alternatively, within the framework of so-called conformal invariant Weyl geometric quadratic gravity ghilen-2019 , was investigated.
-The static spherically symmetric solution found in harko-epjc-2022 to the equations of motion of the latter theory was one of three cases studied in harko-prd-2023 , showing a good fit of the galactic rotation curves.
-For the specific form of the Weyl vector with only a radial nonvanishing component, chosen in harko-prd-2023 ; harko-epjc-2022 , the resulting vacuum EOMs coincide exactly with (55) with the addition of a symmetry-preserving potential $V_{0}\phi^{4}/24$, for the specific choice $V_{0}=1$, once the static spherically symmetric metric (in harko-prd-2023 ; harko-epjc-2022 the BD field $\Phi=\phi^{2}$ was used):
+The possibility that the observed behavior of test particles outside galaxies, which is usually explained with the help of dark matter, could be explained, alternatively, within the framework of so-called conformal invariant Weyl geometric quadratic gravity ghilen-2019 , was investigated in harko-prd-2023 .
+The static spherically symmetric solution found in harko-epjc-2022 of the EOM of the latter theory was one of three cases studied in harko-prd-2023 , showing a good fit to the galactic rotation curves.
+For the specific form of the Weyl vector with only a radial nonvanishing component, chosen in harko-prd-2023 ; harko-epjc-2022 , the resulting vacuum EOMs coincide exactly with (55) with the addition of a Weyl symmetry-preserving potential $\phi^{4}/24$ (in harko-prd-2023 ; harko-epjc-2022 the BD field $\Phi=\phi^{2}$ was used).
+Then, there is no doubt that the solution found in these publications is a vacuum solution of Weyl-invariant CGR theory (24) with a cosmological constant.
+As we shall see, the solution involves a specific function $\phi=\phi(r)$, so that it is in fact a gauge of vacuum CGR theory (55) with a cosmological constant.
+Let us discuss the mentioned solution that provides a good explanation of the galaxies’ rotation curves.
+In harko-prd-2023 ; harko-epjc-2022 the static spherically symmetric metric
 $ds^{2}=-e^{\nu(r)}dt^{2}+e^{\lambda(r)}dr^{2}+r^{2}d\Omega^{2},$ (58)
-is substituted. The solution studied in harko-prd-2023 reads:
+was assumed. The solution studied in harko-prd-2023 reads:
 $e^{\nu}=e^{-\lambda}=\alpha-2+\frac{(\alpha-3)C_{2}}{3r}+(\alpha-1)\frac{r}{C_{2}}+C_{3}r^{2},$ (59)
@@
 In these equations, we have defined the following constants: $2m\equiv\sqrt{C_{1}/6}(1-C_{3}C^{2}_{2}-C_{1}/12)$ and $\Lambda_{0}\equiv C_{1}/4$.
-Therefore, according to active CTs, the GR Schwarzschild-de Sitter metric (63), which is not suitable for explaining the galactic rotation curves, and the conformal solution (62), which can provide a good background for explaining the rotation curves, belong to the same conformal equivalence class.
-Both are equally plausible solutions of vacuum EOMs (55) that represent different gravitational states according to AWTs.
-Only experiment (in the present case, the observational data on galactic rotation curves) can choose a specific gauge among the infinity of conformally related gauges in the conformal class.
+This is an expected result, since (59) is a vacuum solution of CGR with a cosmological constant.
+This means that GR $\Lambda$ is a gauge of this theory when we set $\phi(r)=\sqrt{6}M_{\text{pl}}$.
+Hence, any solution (61) that is conformal to the Schwarzschild-de Sitter vacuum solution (62) is also a solution of (55) plus a cosmological constant term.
+According to active CTs, the GR Schwarzschild-de Sitter metric (62), which is not suitable for explaining the galactic rotation curves, and the conformal solution (61), which can provide a good background for explaining the rotation curves, belong to the same conformal equivalence class $\mathfrak{C}_{\text{conf}}$ (51).
+Both are equally plausible solutions of vacuum EOM (55) (plus a $\Lambda$ term) that represent different global gravitational states according to AWTs.
+Only experiment (in the present case, the observational data on galactic rotation curves) can choose a specific gauge among the infinity of conformally related gauges in the conformal class $\mathfrak{C}_{\text{conf}}$ of vacuum CGR theory (plus a $\Lambda$ term).
 If we look at the above problem from the point of view of the passive approach to CTs, as done, for instance, in hobson-epjc-2022 , there are no physical consequences of Weyl symmetry.
@@
 This makes clear that the authors consider that only conformal invariant quantities, in particular the conformal invariant metric $\mathfrak{g}_{\mu\nu}\equiv(\phi/\sqrt{6}M_{\text{pl}})^{2}g_{\mu\nu}$, have observational meaning.
-However, since the Weyl invariant metric $\mathfrak{g}_{\mu\nu}=(\phi/\sqrt{6}M_{\text{pl}})^{2}g_{\mu\nu}=\hat{g}_{\mu\nu}$, only the GR-SdS geometry (63) has an observational meaning.
-Yet, we must recall that this analysis is correct only if one assumes from the start that the passive Weyl transformations represent the only correct approach to conformal symmetry.
+However, since the Weyl-invariant metric $\mathfrak{g}_{\mu\nu}=(\phi/\sqrt{6}M_{\text{pl}})^{2}g_{\mu\nu}=\hat{g}_{\mu\nu}$, only the GR-SdS geometry (62) has an observational meaning.
+However, we must recall that this analysis is correct only if one assumes from the start that the passive Weyl transformations represent the only correct approach to conformal symmetry.
 As we have shown, if we follow this approach, the action and derived EOM of CGR theory, when written in terms of the physical quantities, coincide with the GR action and derived EOM, so Weyl symmetry is lost or is, at least, a fictitious symmetry.
 Recall that in this case the “auxiliary” fields that suffer the Weyl transformations: $g_{\mu\nu}$, $\phi$, $\chi_{i}$, and $m$ are removed from the covariant formulation (44).
-VII Weyl invariant cosmology and the many-worlds perspective
-The CGR theory (37) is a much more versatile alternative to solving current gravitational puzzles than GR alone, since the latter belongs to the class of conformal equivalence of CGR.
-Suppose that we have an GR solution $\hat{g}_{\mu\nu}(x)$ and that a gravitational state is labeled “kk”;
-${\cal S}_{k}:\left(\phi_{k}(x),g^{(k)}_{\mu\nu}(x),...\right),\;\;k=1,2,...\,,N,$ (64)
-where $N\rightarrow\infty$ and the ellipsis stands for other fields coupled to gravity, is such that the fields $\phi_{k}(x)$ and $g^{(k)}_{\mu\nu}(x)$ satisfy the CGR-EOM (38) for any $k$ and any gauge choice $\phi_{k}=\phi_{k}(x)$.
-The different metric tensors $g^{(k)}_{\mu\nu}(x)$ are related to the GR solution through (52);
-$g^{(k)}_{\mu\nu}(x)=\left(\frac{\sqrt{6}M_{\text{pl}}}{\phi_{k}(x)}\right)^{2}\hat{g}_{\mu\nu}(x).$
-Hence, instead of a single GR gravitational state ${\cal S}_{\text{gr}}$, in CGR there is an infinite set of gravitational states: $\mathfrak{S}_{\text{cgr}}=\{{\cal S}_{\text{gr}},{\cal S}_{1},{\cal S}_{2},...,{\cal S}_{N}\}$, including the GR gauge itself, which are available to describe a given set of gravitational phenomena.
+VII Weyl-invariant cosmology and the many-worlds perspective
+The CGR theory (24) is a much more versatile alternative to solving current gravitational puzzles than GR alone, since the latter belongs to the class of conformal equivalence class $\mathfrak{C}_{\text{conf}}$ of CGR.
+Suppose that we have an arbitrary GR solution labeled “AA” (from now on the upper case latin indexes $A$, $B$, $C$, etc., run over all possible solutions of GR-EOM $A=1,2,3,...,N_{s}$, where $N_{s}$ is the number of different solutions), and that a corresponding global gravitational state labeled “kk” is represented as (compare with (50));
+$\mathfrak{G}(A|k):\{\phi_{k}(x),g^{(A|k)}_{\mu\nu}(x),...\},\;\;A=1,2,...,N_{s};\;\;k=1,2,...,N,$ (63)
+where $N\rightarrow\infty$, and the ellipsis in curly brackets stands for other fields coupled to gravity.
+Let us assume that the fields $\phi_{k}(x)$ and $g^{(A|k)}_{\mu\nu}(x)$ satisfy the CGR-EOM (35) for any $k$ and any gauge choice $\phi_{k}=\phi_{k}(x)$.
+The different metric tensors $g^{(A|k)}_{\mu\nu}(x)$ are related to the GR solution labeled “AA” through (49);
+$g^{(A|k)}_{\mu\nu}(x)=\left(\frac{\sqrt{6}M_{\text{pl}}}{\phi_{k}(x)}\right)^{2}\hat{g}^{(A)}_{\mu\nu}(x),$
+where $\hat{g}^{(1)}_{\mu\nu}(x)$, $\hat{g}^{(2)}_{\mu\nu}(x)$$\hat{g}^{(N_{s})}_{\mu\nu}(x)$, are different solutions of the GR-EOM.
+Hence, instead of a single GGS: $\mathfrak{G}_{\text{gr}}(A):\{\sqrt{6}M_{\text{pl}},\hat{g}^{(A)}_{\mu\nu}(x),$...\}, corresponding to each GR solution “AA”, in CGR there is an infinite set of global gravitational states: $\mathfrak{C}_{\text{conf}}(A)=\{\mathfrak{G}_{\text{gr}}(A),\mathfrak{G}(A|1),\mathfrak{G}(A|2),...,\mathfrak{G}(A|N-1)\}$ ($A=1,2,...,N_{s}$), including the GR solution itself, which are available to describe a given set of gravitational phenomena.
+Only experiments can pick out the GGS in $\mathfrak{C}_{\text{conf}}(A)$ that more closely describes given phenomena.
 Let us illustrate the versatility of CGR in the cosmological framework. Here, we assume a Friedmann-Robertson-Walker (FRW) line-element with flat spatial sections:
 $ds^{2}=-dt^{2}+a^{2}(t)\delta_{ij}dx^{i}dx^{j},$ (64)
-where $t$ is cosmic time, $a(t)$ is the scale factor, and the latin indices run through space: $i,j=1,2,3$.
+where $t$ is cosmic time, $a(t)$ is the scale factor, and the lower case latin indices run through space: $i,j=1,2,3$.
 We further assume that the background matter content is a perfect fluid with energy density $\rho$ and pressure $p$ that satisfy the following EOS: $p=(\gamma-1)\rho$, where $\gamma$ is the barotropic index of the fluid.
@@
 $a_{\text{gr}}(\tau)=\frac{\bar{C}}{\sqrt{6}M_{\text{pl}}}\,(\tau-\tau_{0})^{\frac{2}{3\gamma-2}}.$
-In CGR theory, instead of a single gravitational state for each value of the barotropic index $\gamma$, as in GR, we have an infinite class of gravitational states:
-$\mathfrak{S}_{\text{cgr}}=\left\{\left(\sqrt{6}M_{\text{pl}},a_{\text{gr}}(\tau)\right),\left(\phi_{1}(\tau),a_{1}(\tau)\right),...,(\phi_{N}(\tau),a_{N}(\tau))\right\}.$
+In CGR theory, instead of a single GGS for each value of the barotropic index $\gamma$, as in GR, we have an infinite class of global gravitational states:
+$\mathfrak{C}_{\text{conf}}=\left\{\left(\sqrt{6}M_{\text{pl}},a_{\text{gr}}(\tau)\right),\left(\phi_{1}(\tau),a_{1}(\tau)\right),...,(\phi_{N-1}(\tau),a_{N-1}(\tau))\right\}.$
 These are related by the following Weyl transformation:
-$a_{k}(\tau)=\left[\frac{\sqrt{6}M_{\text{pl}}}{\phi_{k}(\tau)}\right]a_{\text{gr}}(\tau),\;\;k=1,2,...,N.$
+$a_{k}(\tau)=\left[\frac{\sqrt{6}M_{\text{pl}}}{\phi_{k}(\tau)}\right]a_{\text{gr}}(\tau),\;\;k=0,1,2,...,N-1,$
+where the label $k=0$ is used exclusively to denote the GR gauge. That is, $\phi_{0}=\sqrt{6}M_{\text{pl}}$.
 The large (in principle infinite) variety of cosmic behaviors in $\mathfrak{S}_{\text{cgr}}$ is what makes the CGR theory and, therefore, Weyl symmetry attractive within the cosmological context and, at the same time, yields the most unappealing feature of this theory when it is analyzed under the active approach: the loss of predictability.
-The Weyl invariant CGR theory is not capable of defining a unique gravitational state, but a whole class of them, so that the definition of the state that correctly describes a given gravitational phenomenon is left to experiment.
+The Weyl-invariant CGR theory is not capable of defining a unique GGS, but a whole class of them, so that the definition of the state that correctly describes a given gravitational phenomenon is left to experiment.
 Nevertheless, this theory can be assigned some predictive power.
@@
 Hence, the above argument faraoni_prd_2007 is not universal and can be circumvented.
-In addition, even if a measuring unit, for example, the mass unit, is point-dependent, there are ways in which this effect can be experimentally revealed.
+VIII Observational consequences of CGR: The redshift measurements
+Even if measuring units, for example, the mass unit, are point-dependent, there are ways in which this effect can be experimentally revealed.
 For example, the point-dependent property of the masses of timelike fields, and so the fifth force appearing in (36), cannot be detected in local experiments because any measuring stick is made up of fields, so it suffers the same universal point-dependence as the mass being measured, and the effect cancels out.
-However, in experiments that involve the detection of photons emitted at a distant point in spacetime, namely redshift measurements, the effect can be measured (at least in principle).
-VIII Observational consequences of CGR: The redshift measurements
+However, in experiments that involve the detection of photons emitted at a distant point in spacetime, namely redshift measurements, at least in principle, the effect can be measured.
 Let us explain the peculiar contribution to the redshift of frequencies carried by the Weyl symmetry.
 In CGR theory, two phenomena contribute to the redshift of frequencies: 1) the gravitational redshift (also curvature redshift) and 2) the redshift due to variation of masses in spacetime (mass redshift).
-The former effect is due to the propagation of photons in a curved background, while the latter is due to the point-dependent property of masses: $m=m(x)=\mu\,\phi(x)$.
+The former effect is due to the propagation of photons in a curved background, while the latter is due to the point-dependent property of masses: $m=m(x)=\kappa\,\phi(x)$.
 This mass dependence affects the redshift because the frequency of a photon emitted/absorbed by some atom (atomic number $Z$) is given by the Rydberg formula:
@@
 where $m_{e}$ is the mass of the electron, $c$ is the speed of light, $h$ is the Planck constant (here, temporarily we assume that $c\neq h\neq 1$), $\alpha$ is the fine-structure constant ($\alpha\approx 1/137$), while $n_{l}$ and $n_{u}$ are the principal quantum numbers of the lower and upper energy levels, respectively.
-Suppose that two identical atoms are located at spacetime points $x$ and $x_{0}$, respectively.
+Consider two identical atoms, one of which emits a photon at some time $t$ in the past, and the other atom absorbs the photon at present time $t_{0}$.
 The relative shift of the frequency of the emitted/absorbed photons is given by
-$z_{m}=\frac{\gamma_{e/a}(x)-\gamma_{e/a}(x_{0})}{\gamma_{e/a}(x_{0})}=\frac{m_{e}(x)-m_{e}(x_{0})}{m_{e}(x_{0})}=\frac{\phi(x)}{\phi(x_{0})}-1.$ (72)
+$z_{m}=\frac{\gamma_{e/a}(t)-\gamma_{e/a}(t_{0})}{\gamma_{e/a}(t_{0})}=\frac{m_{e}(t)-m_{e}(t_{0})}{m_{e}(t_{0})}=\frac{\phi(t)}{\phi(t_{0})}-1.$ (71)
 Note that this emission/absorption frequency does not depend on the specific atom that emits/absorbs the photon, nor on the specific atomic transition $n_{u}\rightleftharpoons n_{l}$.
-Note also that in the GR gauge, since $\phi(x)=\phi(x_{0})=M_{\text{pl}}$, this type of redshift does not arise.
+Note also that in the GR gauge, since $\phi(t)=\phi(t_{0})=\sqrt{6}M_{\text{pl}}$, this type of redshift does not arise.
 Meanwhile, the curvature redshift is due to the dynamics of the photon frequency during its propagation in a curved background.
-The frequency of a photon emitted/absorbed at some point $x$ in spacetime can be found as a solution of the null-geodesics (42): $\gamma_{e/a}(x)=\gamma^{0}_{e/a}/\sqrt{-g_{00}(x)}$, where $\gamma^{0}_{e/a}$ is some constant frequency.
-The curvature redshift resulting from the propagation of a photon from point $x$ to $x_{0}$ and the overall redshift are given by:
-$z_{c}=\frac{\gamma_{e/a}(x)-\gamma_{e/a}(x_{0})}{\gamma_{e/a}(x_{0})}=\sqrt{\frac{g_{00}(x_{0})}{g_{00}(x)}}-1,$ (73)
+The frequency of a photon at any time $t$ while it propagates in an expanding universe can be found as a solution of the null-geodesics (39) (see equation (101) of appendix A): $\gamma(t)=\gamma_{0}/a(t)$, where $\gamma_{0}$ is some constant frequency.
+The curvature redshift resulting from the propagation of a photon in an expanding (homogeneous and isotropic) FRW background between times $t$ and $t_{0}$, and the overall redshift are given by:
+$z_{c}=\frac{\gamma(x)-\gamma(x_{0})}{\gamma(x_{0})}=\frac{a(t_{0})}{a(t)}-1\;\Rightarrow\;a(t)=\frac{a(t_{0})}{1+z_{c}},$ (72)
 and
-$z_{\text{tot}}=z_{c}+z_{m}=\sqrt{\frac{g_{00}(x_{0})}{g_{00}(x)}}+\frac{\phi(x)}{\phi(x_{0})}-2,$ (74)
-respectively. If we take into account the relationship (52) between the metric $g_{\mu\nu}(x)$ of an arbitrary gauge and the metric $\hat{g}_{\mu\nu}(x)$ in the GR gauge, we get,
-$z_{\text{tot}}=\frac{\phi(x)}{\phi(x_{0})}\sqrt{\frac{\hat{g}_{00}(x_{0})}{\hat{g}_{00}(x)}}+\frac{\phi(x)}{\phi(x_{0})}-2,$
-or if we realize that the GR redshift $z=\sqrt{\hat{g}_{00}(x_{0})/\hat{g}_{00}(x)}-1$, then
-$z_{\text{tot}}=\frac{\phi(x)}{\phi(x_{0})}\,z+2\left(\frac{\phi(x)}{\phi(x_{0})}-1\right).$ (75)
-In a cosmological context, if we replace $t\rightarrow x$, where $t$ is cosmic time, and set $x_{0}=0$ (present time), under the assumption of very small $\phi$-variations, we can write
-$\frac{\phi(t)}{\phi(0)}=1+\frac{\dot{\phi}(0)}{\phi(0)}\Delta t=1+\epsilon,$ (76)
-where a dot means derivative with respect to cosmic time, we assume that $\phi(0)=\sqrt{6}M_{\text{pl}}$, and that $\epsilon>0$ is a small positive number, so that
-$z_{\text{tot}}=(1+\epsilon)z+2\epsilon.$ (77)
-This simplified expression shows that, in CGR cosmology, the measured redshift $z_{\text{tot}}$ is greater than the redshift according to GR.
-Although (77) is an approximate equation, for redshifts $\sim 1$ and a bit larger, this result can have obvious implications for the explanation of the late-time dynamics of our Universe.
-Since the fifth force that arises in (39) and in (40) is due to the point-dependent property of the mass: $m(x)=\kappa\,\phi(x)$, accurate measurements of the second term in the RHS of (77) would entail indirect measurements of the extra force $f^{\alpha}=h^{\alpha\lambda}\partial_{\lambda}\phi/\phi$.
-To illustrate the modification of the frequency redshift achieved by the assumption of CGR theory under the active approach to Weyl transformations, let us consider the Pound-Rebka and Pound-Snider experiments pound-1960 ; pound-1964 ; pound-1965 .
+$z_{\text{tot}}=z_{c}+z_{m}=\frac{a(t_{0})}{a(t)}+\frac{\phi(t)}{\phi(t_{0})}-2,$ (73)
+respectively. If we take into account the relationship (46); that is,
+$a(t)=\frac{\sqrt{6}M_{\text{pl}}}{\phi(t)}\,\hat{a}(t),$
+where $\hat{a}$ is the scale factor in the GR gauge, we get that
+$z_{\text{tot}}=\frac{\phi(t)}{\phi(t_{0})}\frac{\hat{a}(t_{0})}{\hat{a}(t)}+\frac{\phi(t)}{\phi(t_{0})}-2,$
+or if we realize that the GR redshift $z={\hat{a}(t_{0})/\hat{a}(t)}-1$, then
+$z_{\text{tot}}=\frac{\phi(t)}{\phi(t_{0})}\,z+2\left(\frac{\phi(t)}{\phi(t_{0})}-1\right).$ (74)
+In a cosmological context, we replace $t\rightarrow z$, and set $t_{0}=0$ (present time):
+$z_{\text{tot}}=\frac{\phi(z)}{\phi(0)}\,z+2\left(\frac{\phi(z)}{\phi(0)}-1\right).$ (75)
+Let us for simplicity and definiteness choose the linear $z$ gauge
+$\frac{\phi(z)}{\phi(0)}=1+\frac{\phi^{\prime}(0)}{\phi(0)}z=1+\mu z,$ (76)
+where a prime means derivative with respect to $z$ and $\mu$ is a dimensionless constant. We assume that $\phi(0)=\sqrt{6}M_{\text{pl}}$. Hence, in this “linear gauge”
+$z_{\text{tot}}=(1+2\mu+\mu z)z.$ (77)
+The expression (77) shows that in this gauge of CGR cosmology, the measured redshift $z_{\text{tot}}$ is greater than the redshift according to the GR gauge.
+This result can have obvious implications for the explanation of the late-time dynamics of our Universe.
+Since the fifth force that arises in (36) and in (37) is due to the point-dependent property of the mass: $m(x)=\kappa\,\phi(x)$, accurate measurements of the second term in the RHS of (75); $2\mu z$, would entail indirect measurements of the extra force $f^{\alpha}=h^{\alpha\lambda}\partial_{\lambda}\phi/\phi$.
+VIII.1 SNIa redshift measurements: Weyl symmetry vs dark energy
+[…]
+Measuring distances to distant stars is a challenging task.
+Several quantities, such as the apparent magnitude $m$, the absolute magnitude $M$ (logarithmic measures of flux and luminosity, respectively), and the luminosity distance $d_{L}$ are involved ccp-padman ; dodelson-book .
+The latter is related to the energy flux $\Phi$ measured by the observer at $z=0$, which comes from a distant source with actual luminosity $L$; $d_{L}=\sqrt{L/4\pi\Phi}$.
+It is, in general, a model-dependent quantity.
+It can be related to theoretical quantities as it follows dodelson-book :
+$d_{L}=\frac{d_{C}}{a}=(1+z)d_{C},$ (78)
+where the comoving distance $d_{C}$ is given by
+$d_{C}=\int_{t}^{0}\frac{dt^{\prime}}{a(t^{\prime})}=\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})}.$ (79)
+[…]
+$H(z)=H_{0}\sqrt{\Omega_{m,0}(z+1)^{3}+\Omega_{\Lambda,0}},$ (80)
+where the dimensionless energy density of matter and of the cosmological constant are given by:
+$\Omega_{m,0}\equiv\frac{\rho_{0}}{3H^{2}_{0}},\;\Omega_{\Lambda,0}\equiv\frac{\Lambda}{3H^{2}_{0}},$
+respectively. In the above equations, $H_{0}$ and $\rho_{m,0}$ are the present values of the Hubble parameter and of the matter density.
+For simplicity, we shall investigate the modification to the redshift in the linear gauge (76) where the relationship (77) between the overall (measured) redshift in this gauge and the GR redshift $z$ takes place.
+Note that in the expression for the comoving distance $d_{C}$, the expression under the integral: $dt/a(t)$ is a Weyl form-invariant quantity, so it has the same functional dependence in the linear gauge and in the GR gauge ($\Lambda=0$, $\Omega_{m,0}=1$).
+That is, $\int_{t}^{0}dt^{\prime}/a(t^{\prime})=\int_{0}^{z}dz^{\prime}/(z^{\prime}+1)^{3/2}$.
+However, a subtlety must be considered: while the integration limit in GR gauge is $z$, in the “linear gauge” it must be $z_{\text{tot}}$, which is the measured redshift in that gauge.
+This means that in GR gauge we have:
+$H_{0}d_{C}=\int_{0}^{z}\frac{dz^{\prime}}{(z^{\prime}+1)^{3/2}}=2\left(\frac{\sqrt{z+1}-1}{\sqrt{z+1}}\right),$ (81)
+while in the linear gauge (77) we get that
+$H_{0}d^{\text{lin}}_{C}=\int_{0}^{z_{\text{tot}}}\frac{dz^{\prime}}{(z^{\prime}+1)^{3/2}}=2\left(\frac{\sqrt{z_{\text{tot}}+1}-1}{\sqrt{z_{\text{tot}}+1}}\right)=2\left(1-\frac{1}{\sqrt{1+(1+2\mu+\mu z)z}}\right),$ (82)
+where $z_{\text{tot}}$ is given by (77).
+Another subtlety to be considered is that, according to (72) and (78), in an arbitrary gauge $d_{L}=d_{C}/a=d_{C}(1+z_{c})$.
+That is, only the curvature redshift matters, since it is the one related to the scale factor (72) (see EQ.
+(102) in appendix A).
+Hence, in the linear gauge (76), the expression for the luminosity distance is given by:
+$H_{0}d^{\text{lin}}_{L}=(1+z_{c})H_{0}d^{\text{lin}}_{C}=(1+\mu z)(1+z)H_{0}d^{\text{lin}}_{C},$
+where $H_{0}d^{\text{lin}}_{C}$ is given by (82).
+According to Eq.
+(31) of reference ccp-padman , the redshift-dependent apparent magnitude $m$ of distant SNIa is given by [^20] Since we consider the dimensionless combination $H_{0}d_{L}$ rather than $d_{L}$, this means that in the equation to determine the apparent magnitude a term $5\log_{10}h$ is to be taken away, where $H_{0}=100h\,\text{km}\,\text{s}^{-1}\,\text{Mpc}^{-1}$.
+$m=M-5\log_{10}h+42.38+5\log_{10}\left(H_{0}d_{L}\right).$
+Here we take M=19.09M=-19.09 and $h=0.72$ so that,
+$m=5\log_{10}(H_{0}d_{L})+24,$ (83)
+which is the master equation to determine the theoretical values of the apparent magnitude.
+[…]
+Now we are in a position to explain how the accelerated expansion is naturally avoided in the present conformal invariant theory.
+As an illustration, we shall consider one of the first data sets on high redshift supernovae measurements perlmutter-1999 , which provided early evidence for accelerated expansion of the universe.
+We shall consider three cases: 1) the GR model, where
+$H_{0}d_{L}=(1+z)H_{0}d_{C},$ (84)
+[…]
+$H_{0}d^{\Lambda}_{L}=(1+z)H_{0}d^{\Lambda}_{C},\;H_{0}d^{\Lambda}_{C}=\int_{0}^{z}\frac{dz^{\prime}}{\sqrt{\Omega_{m,0}(z^{\prime}+1)^{3}+\Omega_{\Lambda,0}}},$ (85)
+and 3) the Weyl-invariant CGR model:
+$H_{0}d^{\text{lin}}_{L}=(1+\mu z)(1+z)H_{0}d^{\text{lin}}_{C},$ (86)
+where $H_{0}d^{\text{lin}}_{C}$ is given by (82).
+We compute the theoretical values of the apparent magnitude $m$ (83) for the three cases and compare the results.
+These are checked against redshift data.
+[…]
+[…]
+VIII.2 Pound-Rebka-Snider experiments
+To further illustrate the modification of the frequency redshift in terrestrial experiments under the assumption of CGR theory and the active approach to Weyl transformations, let us consider the Pound-Rebka and Pound-Snider experiments pound-1960 ; pound-1964 ; pound-1965 .
 A target $T_{h}$ made of some atom is placed at some height $h$ on top of a tower on the Earth’s surface (radial coordinate $r=R_{\oplus}+h$, where $R_{\oplus}$ is the mean radius of the Earth) and an identical target $T_{0}$ (made of the same atom) is placed at the bottom of the tower (radial coordinate $r_{0}=R_{\oplus}$.) Atoms in $T_{h}$ are in some excited state, so these can emit photons to return to the ground state.
@@
 [^21] We omit experimental details such as, for instance, that the target at the bottom of the tower must be mounted in a centrifuge with controlled rotation speed, or that the temperature of the emitter had to be controlled to high precision, or that the role of the emitter and of the absorber was continuously exchanged, etc.
-If we expand (75) $\phi(r)=\phi(r_{0})+\phi^{\prime}(r_{0})\Delta r$, where $\Delta r=r-r_{0}=h$, then we get:
+If we expand (74) $\phi(r)=\phi(r_{0})+\phi^{\prime}(r_{0})\Delta r$, where $\Delta r=r-r_{0}=h$, then
 $z_{tot}=(1+\alpha)z+2\alpha,$ (87)
-where we have introduced the following notation:
-$\alpha\equiv\frac{\phi^{\prime}(r_{0})}{2\phi(r_{0})}\,h.$ (79)
+where
+$\alpha\equiv\frac{\phi^{\prime}(r_{0})}{\phi(r_{0})}\,h\ll 1.$ (88)
 Let us evaluate (88) for different choices of the gauge scalar $\phi=\phi(r)$.
@@
 In this case,
-$\alpha=\frac{2m^{4}}{R^{4}_{\oplus}}\left(\frac{h}{R_{\oplus}}\right)=1.66\times 10^{-42},$ (80)
+$\alpha\approx\frac{2m^{4}}{R^{4}_{\oplus}}\left(\frac{h}{R_{\oplus}}\right)=1.66\times 10^{-42},$ (89)
 where we used the following values of the parameters: the gravitational constant $G=6.67\times 10^{-11}\,\text{m}^{3}/\text{s}^{2}\text{kg}$, the mass of Earth $M_{\oplus}=5.97\times 10^{24}\,\text{kg}$, mean radius of Earth $R_{\oplus}=6.37\times 10^{6}\,\text{m}$, the speed of light in vacuum $c=2.99\times 10^{8}\,\text{m}/\text{s}$, and the distance between the source and the absorber established in the original experiment by Pound and Rebka pound-1960 : $h=22.55\,\text{m}$.
@@
 For comparison, the value of the redshift measured in pound-1960 was $z=2.5\times 10^{-15}$.
-Another illustrative example can be given by the arbitrary gauge choice $\phi(r)=\sqrt{6}M_{\text{pl}}(1-2m/r)^{-1/4}$. In this case,
-$\alpha=\frac{m\,h}{2R^{2}_{\oplus}}\approx 1.23\times 10^{-15},$
+Another gauge choice that is suitable for removing the spacetime singularity, according to bambi-2017-a , is given by (57): $\phi(r)=\sqrt{6}M_{\text{pl}}/(1+m^{2}/r^{2})$.
+In this case,
+$\alpha\approx 2\left(\frac{m}{R_{\oplus}}\right)^{2}\frac{h}{R_{\oplus}}=3.46\times 10^{-24},$
+so that experimental check of this gauge is also beyond the current experimental capabilities.
+Another illustrative example can be given by the arbitrary gauge choice $\phi(r)=\sqrt{6}M_{\text{pl}}/(1+m/r)^{1/4}$. In this case,
+$\alpha\approx\frac{m\,h}{4R^{2}_{\oplus}}=0.6\times 10^{-15},$
 which is of the order of the redshift measured in the Pound-Rebka experiment.
 Since no such additional contribution has been obtained in pound-1960 , this means that this gauge choice is experimentally ruled out.
-IX Discussion and concluding remarks
-Here we have demonstrated that the very well-known result within the HEP community that local-scale invariance in flat spacetime requires the vanishing of the trace of the matter stress-energy tensor is not universal in curved spacetime.
+IX Quantum-mechanical removal of spacetime singularities
+Let us discuss the possibility that quantum-mechanical considerations within the CGR framework may resolve the spacetime-singularity issue; that is, that quantum theory may remove the spacetime singularities that plague classical gravitational theory penrose-1 ; penrose-2 ; penrose-3 ; hawking-1 ; berger ; bosma .
+According to the previous discussion in sections V and VII, a global gravitational state $\mathfrak{G}(A|k)=\{\phi_{k}(x),g^{(A|k)}_{\mu\nu}$,...\} can be thought as a fourth-dimensional sheet where the whole history of all fields is drawn: the metric tensor $g^{(A|k)}_{\mu\nu}=g^{(A|k)}_{\mu\nu}(t,\vec{x})$, the scalar field $\phi_{k}=\phi_{k}(t,\vec{x})$, and each matter field coupled to gravity $\chi^{(A)}_{i}=\chi^{(A)}_{i}(t,\vec{x})$ (for simplicity we assume that the matter fields are fixed, that is, these are the same in any gauge).
+Each sheet is labeled by the pair $(A|k)$, i.e., by the index corresponding to the given GR solution $A=1,2,...,N_{s}$, and by the index corresponding to the gauge choice k=1,2,,Nk=1,2,...,N ($N\rightarrow\infty$).
+Let us call it a “gauge sheet”.
+[^22] The gauge sheet is just a vague illustration of the actual meaning of a gauge or what we call here a global gravitational state.
+A clearer illustration must be consistent with the well-posed initial-value formulation of CGR theory.
+We assume that such a gauge sheet, for example, the one associated to an arbitrary solution “AA” and a gauge choice $\phi_{k}=\phi_{k}(x)$, is a globally hyperbolic spacetime $({\cal M},g^{(A|k)}_{\mu\nu})_{\phi_{k}}$ admitting foliation by Cauchy hypersurfaces $\sum_{t}$, parametrized by a global time function $t$.
+We may view this globally hyperbolic spacetime as representing the time development of a Riemannian metric on a fixed three-dimensional manifold wald-book .
+A given classical gravitational phenomenon is well-described by a single gauge sheet within the infinite collection of $N\to\infty$ gauge sheets, but quantum gravity must take into account all of them, since all of the possible gauges of a given solution contribute to the probability amplitude; [^23] Although the probability amplitude of some solution, in particular (90), is expressible as hawking-2 ; hawking-2 ; hawking-3 : $\varphi=\int{\cal D}\phi\,{\cal D}g\,\exp{iS_{\text{cgr}}[\phi,g]},$ where ${\cal D}\phi$ and ${\cal D}g$, are some measures in the space of all scalar fields $\phi=\phi(x)$ and metrics $g_{\mu\nu}=g_{\mu\nu}(x)$, respectively, and $S_{\text{cgr}}[\phi,g]$ is the classical action of vacuum CGR theory (91), for our qualitative discussion it will be more adequate to work with the sum (90).
+$\varphi(A)=\lim_{N\to\infty}\sum_{k=1}^{N}\varphi(A|k),\;\;\varphi(A|k)=\frac{\varphi_{0}}{N}\exp iS_{\text{cgr}}[\phi_{k},g_{(A|k)}],$ (90)
+where $\varphi(A)$ is the probability amplitude of the solution “AA”, while $\varphi(A|k)$ is the probability amplitude of the gauge “kk” of solution “AA”, $\varphi_{0}/N$ is some normalization constant, and the sum is over all gauges in the conformal class of the GR solution “AA”: $\mathfrak{C}_{\text{conf}}(A)=\{\mathfrak{G}_{\text{gr}}(A),\mathfrak{G}(A|1),\mathfrak{G}(A|2),...,\mathfrak{G}(A|N-1)\}$.
+For simplicity of the analysis, we consider the vacuum CGR theory.
+Hence,
+$S_{\text{cgr}}[\phi,g]=\frac{1}{12}\int_{{\cal V}_{4}}d^{4}x\sqrt{-g}\left[\phi^{2}R+6(\partial\phi)^{2}\right]+\frac{1}{6}\int_{\partial{\cal V}_{4}}d^{3}x\sqrt{-h}\phi^{2}K,$ (91)
+where ${\cal V}_{4}$ is some integration volume, $h_{\mu\nu}=g_{\mu\nu}\pm n_{\mu}n_{\nu}$ is the metric induced on the boundary of the integration volume $\partial{\cal V}_{4}$, which is orthogonal to the unit vector field $n_{\mu}$, and $K=h^{\mu\nu}K_{\mu\nu}$ ($K_{\mu\nu}=h^{\lambda}_{\;\mu}h^{\sigma}_{\;\nu}\nabla_{\lambda}n_{\sigma}$ is the extrinsic curvature tensor of the boundary).
+The boundary term is introduced to compensate for the second derivatives of the gravitational action in the variational procedure hawking-2 .
+This term does not modify the conformal form-invariance of the action (91) (see, for example, appendix A of wands-rev ).
+Due to form-invariance of the classical action $S_{\text{cgr}}[\phi,g]$ under Weyl rescalings (45), each probability amplitude $\varphi(A|k)$ in (90) contributes not only a same amplitude $\varphi_{0}/N$, but also a same phase $S_{\text{cgr}}$, so that $\varphi(A|1)=\varphi(A|2)=\varphi(A|*)=...=\varphi(A|N-1)$, where $\mathfrak{G}(A|*)$ can be any of the gauges in the conformal equivalence class $\mathfrak{C}_{\text{conf}}(A)$.
+Therefore,
+$\varphi(A)=\varphi_{0}\exp iS_{\text{cgr}}[\phi_{*},g_{(A|*)}].$ (92)
+To explain how the spacetime singularity is removed in quantum-mechanical computations, let us consider a specific solution.
+Here, for illustration, we consider the GR vacuum Schwarzschild solution (53):
+$d\hat{s}^{2}=\hat{g}^{(\cancel{0})}_{\mu\nu}dx^{\mu}dx^{\nu}=-\left(1-\frac{2m}{r}\right)dt^{2}+\frac{dr^{2}}{\left(1-\frac{2m}{r}\right)}+r^{2}d\Omega^{2}.$
+Any conformal Schwarzschild metric (54)
+$ds^{2}=g^{(\cancel{0}|*)}_{\mu\nu}dx^{\mu}dx^{\nu}=\left[\frac{\sqrt{6}M_{\text{pl}}}{\phi_{*}(r)}\right]^{2}\left[-\left(1-\frac{2m}{r}\right)dt^{2}+\frac{dr^{2}}{\left(1-\frac{2m}{r}\right)}+r^{2}d\Omega^{2}\right],$ (93)
+is also a solution of the vacuum CGR-EOM (55).
+For example, in bambi-2017-a , the authors showed that the gauge choice $\phi_{*}(r)=\sqrt{6}M_{\text{pl}}/\sqrt{1+(L/r)^{4}},$ where $L$ is some length scale, leads to the Schwarzschild conformal metric being complete, so the resulting geometry represents a nonsingular black hole.
+According to (92), the probability amplitude of the Schwarzschild solution may be given by: $\varphi(\cancel{0})=\varphi_{0}\exp iS_{\text{cgr}}[\phi_{*},g_{(\cancel{0}|*)}],$ where $\phi_{*}$ corresponds to the gauge choice made in bambi-2017-a (see above) and $g^{(\cancel{0}|*)}_{\mu\nu}$ is the conformal Schwarzschild metric (93), which is complete, so there will be no singularity issue.
+Due to conformal form-invariance of the classical CGR action, leading to the same phase contribution to the probability amplitude from any gauge, we can choose any gauge, say $\mathfrak{G}(A|*)$ $\Rightarrow\,S_{\text{cgr}}[\phi_{*},g_{(\cancel{0}|*)}],$ where the singularity is removed.
+This is the way in which the spacetime singularities are avoided in quantum-mechanical computations.
+A final comment on this issue: Within the path integral formulation of gravity, not only those fields that are solutions of the gravitational EOM, like, for example, $\phi_{*}$ and $g^{(\cancel{0},*)}_{\mu\nu},$ matter.
+Any field configurations that do not satisfy the gravitational EOM are to be considered as well.
+Notice, however, that the classical action $S_{\text{cgr}}[\phi,g]$ itself is conformal form-invariant regardless of whether the fields satisfy equations of motion or not.
+That is, any probability amplitudes contribute the same amplitude and phase.
+Hence, we are free to choose one of the singularity-free configurations and substitute it in (92), even if the fields do not satisfy any equation of motion.
+X Concluding remarks
+In this document, we have demonstrated that the very well-known result within the HEP community that local-scale invariance in flat spacetime requires vanishing trace of the matter stress-energy tensor is not universal in curved spacetime.
 The requirement $T^{(m)}=0$ holds wherein the masses of timelike fields are constant parameters that are not affected by the Weyl recalings.
@@
 Our result that under the WTs (45), vanishing of the trace of the stress-energy tensor of matter is not required for Weyl invariance to be a symmetry of theory (24), may be of importance for a revised perspective of the Weyl anomaly capper-duff-1974 ; duff-1977 ; rusos-1984 ; affleck-1986 ; duff-1994 ; odintsov-2000 ; cai-2010 .
-Unfortunately, the investigation of this issue goes clearly beyond the aims of the present document.
+Unfortunately, the investigation of this issue clearly goes beyond the scope of the present document.
+[…]
 The results discussed in this document are valid only for the Weyl-invariant CGR theory (24).
@@
 In this case, the demonstration in Section III is applicable again, so the traceless condition does not arise.
-Interestingly, the above Weyl invariant framework may provide a suitable classical description of the gravitational interactions of matter, which meets Dicke’s principle that the (classical) laws of physics must be invariant under the transformations of units; the name used by Dicke for the Weyl transformations (28).
+Interestingly, the above Weyl-invariant framework may provide a suitable classical description of the gravitational interactions of matter, which meets Dicke’s principle that the (classical) laws of physics must be invariant under the transformations of units; the name used by Dicke for the Weyl transformations (25).
 Frequent arguments against the Weyl invariance of physical laws must be adequately identified as arguments against local scale invariance.
@@
 $g_{\mu\nu}\rightarrow\Omega^{2}g_{\mu\nu},\;\phi\rightarrow\Omega^{-1}\phi,\;\chi_{i}\rightarrow\Omega^{w_{\chi_{i}}}\chi_{i},\;m\rightarrow\Omega^{-1}m.$
-Acknowledgments The author thanks Felipe F Faria for pointing out several bibliographic references to us and for the useful exchange of ideas and critical comments.
-Funding: This work has been supported by FORDECYT-PRONACES-CONACYT under grant CF-MG-2558591.
+Acknowledgments The author thanks Felipe F.
+Faria for pointing out several bibliographic references and for the useful exchange of ideas and critical comments.
+FORDECYT-PRONACES-CONACYT has supported this work under grant CF-MG-2558591.
 Declaration of generative AI and AI-assisted technologies in the manuscript preparation process.
-During the preparation of this work, the author used the WRITEFULL tool on overleaf in order to check spelling and grammar.
+During the preparation of this work, the author used the WRITEFULL tool on Overleaf to check spelling and grammar.
 After using this tool, the author reviewed and edited the content as needed and takes full responsibility for the content of the published article.
-Appendix A Lagrangian density of matter in terms of Weyl invariant quantities
-Here, for illustration, we demonstrate equality (45) for perfect fluids and also for the Dirac fermion.
-Let us start with the Lagrangian density of a perfect fluid.
-In this case, the demonstration of equality (45) is a very simple exercise.
-Let us define the Weyl invariant energy density of the fluid:
-$\mathfrak{r}:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-4}\rho.$ (86)
-Hence, trivially,
-${\cal L}_{\text{fluid}}:=-\sqrt{-g}\,\rho\equiv-\sqrt{-\mathfrak{g}}\,\mathfrak{r},$ (87)
-where the left-hand Lagrangian density is written in terms of the fields that suffer the WTs, while the right-hand Lagrangian is written in terms of Weyl invariant quantities.
-A.1 The Dirac Lagrangian density
-The Lagrangian density of a Dirac fermion $\psi$ ($\bar{\psi}$ is the adjoint spinor of the fermion) in curved space is given by:
-[…]
-where $\cancel{\cal D}:=\gamma^{\mu}{\cal D}_{\mu}$ and the following covariant derivative has been defined:
-${\cal D}_{\mu}\psi=\left(D_{\mu}-\frac{1}{2}\sigma_{ab}e^{b\nu}\nabla_{\mu}e^{a}_{\nu}\right)\psi.$ (89)
-In these equations $a,b,c=0,1,2,3$ are the flat spacetime indices, while $\gamma^{a}$ are the Dirac gamma matrices, $e^{a}_{\mu}$ are the tetrad fields such that $g_{\mu\nu}=\eta_{ab}e^{a}_{\mu}e^{b}_{\nu}$ ($\eta_{ab}$ is the Minkowski metric).
-Besides $\gamma^{\mu}=e^{\mu}_{c}\gamma^{c}$, and
-$\sigma_{ab}=\frac{1}{2}\left[\gamma_{a},\gamma_{b}\right]=\frac{1}{4}\left(\gamma_{a}\gamma_{b}-\gamma_{b}\gamma_{a}\right),$
-are the generators of the Lorentz group in the spin representation.
-Above, we have used the standard definition of the $SU(2)\times U(1)$ gauge derivative, $D_{\mu}\psi$ (see, for example, Equation (32) of quiros_prd_2025 ).
-Let us define the following Weyl invariant quantities:
-$\mathfrak{e}^{a}_{\mu}:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)e^{a}_{\mu},\;\mathfrak{g}_{\mu\nu}:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{2}g_{\mu\nu},\;\Psi:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-\frac{3}{2}}\psi,\;\mathfrak{m}:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-1}m.$ (90)
-We can also compute the Levi-Cività (LC) connection (Christoffel symbols) of the Weyl invariant metric $\mathfrak{g}_{\mu\nu}$:
-$\mathfrak{C}^{\alpha}_{\mu\nu}=\frac{1}{2}\mathfrak{g}^{\alpha\lambda}\left(\mathfrak{g}_{\mu\lambda,\nu}+\mathfrak{g}_{\nu\lambda,\mu}-\mathfrak{g}_{\mu\nu,\lambda}\right),$ (91)
-and the related covariant derivative operator $\mathfrak{V}_{\mu}$. We have
-$e^{b\nu}\nabla_{\mu}e^{a}_{\nu}=\mathfrak{e}^{b\nu}\mathfrak{V}_{\mu}\mathfrak{e}^{a}_{\nu}+\left(\mathfrak{e}^{a}_{\mu}\mathfrak{e}^{b\lambda}-\mathfrak{e}^{b}_{\mu}\mathfrak{e}^{a\lambda}\right)\frac{\partial_{\lambda}\phi}{\phi},$
-so
-$\sigma_{ab}e^{b\nu}\nabla_{\mu}e^{a}_{\nu}=\sigma_{ab}\left(\mathfrak{e}^{b\nu}\mathfrak{V}_{\mu}\mathfrak{e}^{a}_{\nu}+2\mathfrak{e}^{a}_{\mu}\mathfrak{e}^{b\lambda}\frac{\partial_{\lambda}\phi}{\phi}\right).$
-Let us next define the Weyl invariant gamma matrix in a curved background;
-$\Gamma^{\mu}:=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{-1}\gamma^{\mu}=\mathfrak{e}^{\mu}_{a}\gamma^{a}=\mathfrak{e}^{\mu a}\gamma_{a}.$ (92)
-It is straightforward to find
-$\cancel{D}\psi=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{\frac{5}{2}}\left[\Gamma^{\mu}\left(D_{\mu}-\frac{1}{2}\sigma_{ab}\mathfrak{e}^{b\nu}\mathfrak{V}_{\mu}\mathfrak{e}^{a}_{\nu}\right)\Psi+\left(\frac{3}{2}\mathfrak{e}^{b\lambda}\gamma_{b}-\sigma_{ab}\gamma^{a}\mathfrak{e}^{b\lambda}\right)\frac{\partial_{\lambda}\phi}{\phi}\Psi\right],$
-that is,
-$\cancel{D}\psi=\left(\frac{\phi}{\sqrt{6}M_{\text{pl}}}\right)^{\frac{5}{2}}\cancel{\mathfrak{D}}\Psi,$ (93)
-where $\cancel{\mathfrak{D}}\Psi=\Gamma^{\mu}\left(D_{\mu}-\frac{1}{2}\sigma_{ab}\mathfrak{e}^{b\nu}\mathfrak{V}_{\mu}\mathfrak{e}^{a}_{\nu}\right)\Psi$ is the Weyl invariant gauge covariant slash derivative and we have used the identity $\sigma_{ab}\gamma^{a}=3\gamma_{b}/2$.
-Hence,
-$\sqrt{-g}\,\bar{\psi}\cancel{D}\psi=\sqrt{-\mathfrak{g}}\,\bar{\Psi}\cancel{\mathfrak{D}}\Psi.$
-The right-hand side (RHS) of this equality is written in terms of Weyl invariant quantities.
-This serves as an additional demonstration that the left-hand side (LHS) is form-invariant under WTs.
-It is not difficult to demonstrate that,
-$\sqrt{-g}\,m\bar{\psi}\psi=\sqrt{-\mathfrak{g}}\,\mathfrak{m}\bar{\Psi}\Psi.$
-Hence, we have demonstrated that the Lagrangian density of the Dirac spinor fulfills the following equality:
-[…]
-where the LHS is written in terms of the fields that suffer the Weyl transformations, while the RHS is written in terms of Weyl invariant quantities.
+Appendix A Curvature reshift as a solution of the null-geodesic equation
+If we write the null geodesic equation (39) in FRW spacetime (64) we have that,
+$\frac{d\omega}{d\xi}+a\dot{a}\delta_{ij}k^{i}k^{j}=0,$ (99)
+where $\xi$ is the affine parameter along the photon’s path and $k^{\mu}=dx^{\mu}/d\xi$ is the four-wave vector, with components: $k^{0}=dt/d\xi=\omega$ (photon’s cyclic frequency) and $k^{i}=dx^{i}/d\xi$.
+The following expression is satisfied by the wave vector:
+$g_{\mu\nu}k^{\mu}k^{\nu}=0\;\Rightarrow\;-\omega^{2}+a^{2}(t)\delta_{ij}k^{i}k^{j}=0.$
+If substitute this equation back into (99) we get that,
+$\frac{d\omega}{d\xi}+\frac{\dot{a}}{a}\omega^{2}=0,$
+or, if recall that $\omega=2\pi\gamma$ ($\gamma$ is the photon’s frequency) and that $\omega=dt/d\xi$ $\Rightarrow\dot{a}\omega=da/d\xi$, then the null geodesic equation can be written in the following way:
+$\frac{d\gamma}{d\xi}+\frac{1}{a}\frac{da}{d\xi}\gamma=0.$ (100)
+Straightforward integration of this equation leads to the textbook expression for the cosmological redshift:
+$\gamma(t)=\frac{\gamma_{0}}{a(t)},\;\gamma_{0}\equiv e^{C},$ (101)
+where $C$ is an integration constant. Since this redshift effect is due exclusively to the propagation of photons in a curved background (expanding FRW spacetime), we call it curvature redshift:
+$z_{c}=\frac{a(t_{0})}{a(t)}-1\;\Rightarrow\;a(t)=\frac{a(t_{0})}{1+z_{c}}.$ (102)